### v1

# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

Abstract.
Compared with standard learning, adversarially robust learning is widely recognized to demand significantly more training examples. Recent works propose the use of self-supervised adversarial training (SSAT) with external or synthetically generated unlabeled data to enhance model robustness. However, SSAT requires a substantial amount of extra unlabeled data, significantly increasing memory usage and model training times. To address these challenges, we propose novel methods to strategically select a small subset of unlabeled data essential for SSAT and robustness improvement. Our selection prioritizes data points near the model’s decision boundary based on latent clustering-based techniques, efficiently identifying a critical subset of unlabeled data with a higher concentration of boundary-adjacent points. While focusing on near-boundary data, our methods are designed to maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting. Our experiments on image benchmarks show that integrating our selection strategies into self-supervised adversarial training can largely reduce memory and computational requirements while achieving high model robustness. In particular, our latent clustering-based selection method with k-means is the most effective, achieving nearly identical test-time robust accuracies with $5$ to $10$ times less external or generated unlabeled data when applied to image benchmarks. Additionally, we validate the generalizability of our approach across various application scenarios, including a real-world medical dataset for COVID-19 chest X-ray classification. Our Implementations are available as open-source code at this url.
1. Introduction
Over the past decade, it has been repeatedly confirmed that deep neural networks (DNNs) are highly vulnerable to adversarial perturbations (Szegedy et al., 2013). This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness (Goodfellow et al., 2014; Papernot et al., 2016; Buckman et al., 2018; Biggio and Roli, 2018). Among them, adversarial training is regarded as one of the most effective methods (Madry et al., 2017; Wang et al., 2019; Zhang et al., 2019). However, as stated by Schmidt et al. (Schmidt et al., 2018), learning a model that is resilient to adversarial perturbations requires a significantly larger dataset than standard learning. Recent studies have explored self-supervised techniques to expand the training set size of adversarial training algorithms by leveraging unlabeled external (Carmon et al., 2019) or generated data (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023). Despite alleviating the sample complexity barrier and producing models with improved robust accuracies, these methods typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store those data and a much longer training time for algorithms like adversarial training to converge.
Witnessing the challenges of additional memory and computational requirements, we investigate whether the significantly large amount of utilized additional data is inevitable for achieving state-of-the-art adversarial robustness. The ultimate goal of our work is to maximize the model robustness achieved by SSAT algorithms by using as few additional unlabeled data points as possible. Inspired by Zhang et al. (Zhang et al., 2020), which highlights the unequal importance of training examples, we argue that with limited model capacity, self-supervised adversarial learning should also focus on optimizing critical data samples near the model’s decision boundary. Consequently, we propose multiple data selection strategies, including a simple prediction confidence-based selection (PCS) scheme that prioritizes unlabeled data points in which the model’s prediction is most uncertain and more advanced latent clustering-based selection (LCS) schemes. Specifically, we propose two LCS approaches depending on the corresponding clustering technique it builds on: latent clustering-based selection with Gaussian mixture models (LCS-GMM) and latent clustering-based selection with k-means (LCS-KM). For each unlabeled data point, we compute the difference between the highest and second-highest posterior probabilities in the latent feature space for LCS-GMM, while we select vulnerable points by calculating the distance difference between the two nearest cluster centroids for LCS-KM. The set of unlabeled data with the smallest probability or distance difference is considered more critical and further incorporated into self-supervised adversarial training. Moreover, our selection strategies balance the ratio of boundary critical points and the remaining non-boundary points to avoid undesirable overfitting to the boundary data distribution.
By applying these targeted data reduction approaches, we streamline the self-supervised adversarial training process, significantly reducing the required additional unlabeled data while maintaining comparable model robustness against adversarial perturbations. Our work offers valuable insights into how to acquire additional data and employ SSAT in resource-constrained application domains. Below, we further summarize the main contributions of our work.
Contributions. We motivate and formalize the problem task of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section 3). To realize such a goal, we propose various data selection schemes to identify the most critical unlabeled data points, including a straightforward approach based on prediction confidence and two advanced methods based on clustering techniques of unlabeled data in the latent embedding space (Section 4). The proposed methods optimize the performance of SSAT by refining the model’s decision boundary in the input regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points. By focusing on critical unlabeled data points, our methods largely reduce the memory consumption and time complexities of self-supervised adversarial training algorithms. We conduct comprehensive experiments on two image benchmarks and a real-world medical dataset, demonstrating that our data selection schemes significantly reduce the memory consumption and the total running time of self-supervised adversarial training algorithms under various experimental settings (Section 5). Additionally, we explain the insights of the potential reasoning behind the superior efficacy of our LCS-KM method to other data selection schemes (Section 6.1) and provide general guidelines on how to employ our LCS methods for maximized robustness performance (Section 6.2). Finally, we examine the robustness of our methods by varying the adversarial training algorithms, perturbation size and metric (Section 6.3), as well as studying the role of the intermediate model under the regime with insufficient labeled data (Section 6.4), all confirming the usefulness and generalizability of the proposed data selection schemes for improving the efficiency of SSAT algorithms. We believe our work can help promote the development of more scalable, robust learning algorithms, which would be particularly beneficial for resource-constrained environments.
2. Related Work
In this section, we review the most relevant literature on adversarial examples, (self-supervised) adversarial training, and its variants. We also discuss the most representative techniques for improving the efficiency of standard deep learning, including dataset distillation and coreset selection.
2.1. Adversarially Robust Learning
Adversarial Examples. Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions (Szegedy et al., 2013). The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications. Earlier-proposed attacks like fast gradient sign method (FGSM) (Goodfellow et al., 2014) and projected gradient descent (PGD) (Madry et al., 2017) introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks (Carlini and Wagner, 2017), which search for the minimal perturbation that causes misclassification. Recent attacks like Auto-PGD (Croce and Hein, 2020a) automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack (Croce et al., 2020) is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
Adversarial Training. On the defense side, adversarial training and its variants are the most popular learning methods for adversarial robustness. For instance, Madry et al. proposed to train DNNs on adversarially perturbed inputs using PGD attacks (Madry et al., 2017), while Zhang et al. introduced TRADES (Zhang et al., 2019), which balances the trade-off between adversarial robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications. In particular, Rice et al. illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training (Rice et al., 2020), where early stopping is shown to be effective in alleviating such an issue. Techniques, such as Free AT (Shafahi et al., 2019) and Fast AT (Wong et al., 2020), offer promising solutions to reduce the computation for adversarial training. Nevertheless, they suffer from decreased robustness performance.
Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations (Zhang et al., 2020). As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data. In a similar vein, Hua et al. advocated for a more targeted approach (Hua et al., 2021), recommending that PGD training be applied primarily to examples situated near the decision boundaries. This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks. Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
Self-Supervised Adversarial Training. Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research (Chapelle et al., 2009). In the adversarial context, Schmidt et al. (Schmidt et al., 2018) highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning. They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly. Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training (Carmon et al., 2019; Alayrac et al., 2019; Najafi et al., 2019; Zhai et al., 2019). In particular, Carmon et al. proposed robust self-training methods to address the sample complexity issue (Carmon et al., 2019), which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework incorporating labeled and unlabeled data (Najafi et al., 2019), while providing formal guarantees on adversarial robustness.
In addition, leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023). For instance, Gowal et al. employed various unconditional generative models (Gowal et al., 2021), including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further. While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
2.2. Data-Efficient Deep Learning
Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information. Two prominent approaches are dataset distillation and coreset selection.
Dataset Distillation. Dataset distillation compresses large datasets into compact, synthetic datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points. For instance, Wang et al. proposed a foundational approach to dataset distillation (Wang et al., 2018), proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching (Zhao et al., 2020), where distilled data points are optimized to match the gradients of the original dataset. This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data. Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data (Sucholutsky and Schonlau, 2021), facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. On the other hand, coreset selection (Sener and Savarese, 2017; Kim and Shin, 2022) aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset. These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts (Kaushal et al., 2019), while Xia et al. introduced the concept of “moderate coreset” (Xia et al., 2022), focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function (Mirzasoleiman et al., 2020).
In the adversarial context, Dolatabadi et al. proposed adversarial coreset selection (Dolatabadi et al., 2023), offering a task-specific solution to enhance the computational efficiency of adversarial training. In particular, the method constructs a compact data subset by minimizing the gradient approximation error. The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals. While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges. These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time. In such settings, repeatedly selecting data during training may become computationally infeasible. In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
3. Improving Data Efficiency for SSAT
In this section, we first introduce the necessary notations and definitions for readers to understand self-supervised adversarial training (Section 3.1) and then motivate and formulate the problem task of enhancing the data efficiency of SSAT algorithms (Section 3.3).
3.1. Preliminaries on SSAT
We work with self-supervised adversarial training algorithms, initially introduced to the field by Carmon et al. (Carmon et al., 2019). Let $\mathcal{X}\subseteq\mathbb{R}^{n}$ be a $n$-dimensional input space, $\mathcal{Y}$ be the output space of class labels, and $D_{l}$ be the underlying labeled distribution supported on $\mathcal{X}\times\mathcal{Y}$ that we aim to learn an adversarially robust classifier. Let $\mathcal{S}_{l}$ be the training set with examples i.i.d. drawn from $D_{l}$ and $\mathcal{S}_{u}$ be a set of inputs sampled from some unlabeled distribution $D_{u}$ supported on $\mathcal{X}$. In the adversarial machine learning literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and $|\mathcal{S}_{u}|$ is set to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are easier to acquire than well-annotated labeled data. Throughout the paper, we use ||$\cdot$|| ⋅ | to denote the cardinality of the corresponding set. For instance, when $\mathcal{S}_{l}$ corresponds to the $50$K labeled CIFAR-10 training images, Carmon et al. (Carmon et al., 2019) selected $500$K “most CIFAR-10-like but non-identical” images from the whole $80$M Tiny ImageNet data as the unlabeled dataset $\mathcal{S}_{u}$ to be used by SSAT.
To be more specific, SSAT first standardly trains an intermediate model $f_{\hat{\theta}}$ on $\mathcal{S}_{l}$, known as the pseudo labeling function, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$. More rigorously, the intermediate model’s weight parameters $\hat{\theta}$ are learned by minimizing the following objective function:
(1) $\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}(\theta,\bm{x},y),$
where $\mathcal{L}(\cdot)$ denotes the standard loss, such as cross-entropy, that measures the discrepancy between model prediction and the class label. Finally, SSAT trains a model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled dataset $\mathcal{S}_{l}$ and unlabeled dataset $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
(2) $\min_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)+\frac{1}{\gamma|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\mathcal{L}_{\mathrm{adv}}\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$
where $\gamma\geq 0$ controls the contributions between labeled and unlabeled data. In prior literature (Gowal et al., 2021), a typical approach is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\alpha$) but fix the total batch size to optimize for the best SSAT performance. For simplicity, we term $\gamma$ as the per-batch ratio hyperparameter in this paper. When $\frac{1}{\gamma}$ is set as 00, Equation 2 reduces to the training objective of vanilla adversarial training (AT) (Madry et al., 2017). In Equation 2, $\mathcal{L}_{\mathrm{adv}}(\cdot)$ stands for the adversarial loss function: for any $(\bm{x},y)\in\mathcal{S}_{l}$ and $\epsilon>0$,
(3) $\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}(\theta,\bm{x}+\bm{\delta},y),$
where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the ball centered at $\bm{0}$ with radius $\epsilon$ measured in some distance metric, such as $\ell_{p}$-norm. Aligned with prior literature on adversarial training (Madry et al., 2017), we adopt multi-step PGD to obtain an approximated solution to the inner maximization problem specified by Equation 3 during self-supervised adversarial training.
3.2. Limitations of Existing SSAT Methods
Although SSAT methods alleviate the sample complexity barriers of adversarially robust learning (Schmidt et al., 2018) and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. This trend is clearly documented in the leaderboard of RobustBench (Croce et al., 2020). For example, state-of-the-art SSAT methods (Carmon et al., 2019; Gowal et al., 2021; Wang et al., 2023) either select $500$K external Tiny Imagenet data samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original $50$K CIFAR-10 training examples. For example, the method proposed by Gowal et al. (Gowal et al., 2021) achieves around $65\%$ robust accuracy on CIFAR-10 using a WideResNet-70-16 model architecture against $\ell_{\infty}$ perturbations with $\epsilon=8/255$, increasing the robustness performance of vanilla adversarial training (Madry et al., 2017) by a large margin, but relies an extra unlabeled set of $100$M DDPM-generated data. The huge unlabeled dataset required for SSAT significantly increases memory consumption for holding the whole training dataset, which we argue is inefficient and likely to be prohibited for resource-constrained application scenarios. These observations motivate us to explore whether such a considerable amount of unlabeled data can be reduced while preserving the high model robustness attained by SSAT algorithms.
Moreover, we note that SSAT requires a much longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of training epochs compared with vanilla AT (see Figure LABEL:fig:convergence for supporting evidence). Intuitively, extracting useful and robust features through adversarial training from such a large and potentially diverse dataset is more difficult. As we will illustrate in our experiments, the slower convergence can be attributed to the additional large unlabeled datasets (with pseudo labels) that often exhibit higher variance than the original labeled data samples. Once we reduce the size of the unlabeled data involved with specifically designed techniques, faster convergence is expected for SSAT. Since vanilla adversarial training has already been criticized in prior literature (Shafahi et al., 2019; Wong et al., 2020) for its high computational costs for running multi-step PGD to solve the inner maximization problem in Equation 3, state-of-the-art SSAT algorithms incur even higher computational costs (see Table 1 for runtime comparisons), due to the slower convergence rate and larger model architecture they typically adopt. Therefore, it is essential to design efficient SSAT methods, especially for resource-constrained applications, that can better utilize the large amount of extra unlabeled data.
3.3. Problem Formulation
Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for training models with good robustness performance. Inspired by the idea of coreset selection for efficient deep learning (Kaushal et al., 2019; Xia et al., 2022; Mirzasoleiman et al., 2020), we propose to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that self-supervised adversarial training based on $\mathcal{S}_{l}$ and the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using full unlabeled dataset $\mathcal{S}_{u}$. More formally, we aim to solve the following constrained optimization problem:
(4) $\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big{(}\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big{)},\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ stands for the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation 2, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the robustness of the model with parameters $\theta$ against $\epsilon$ perturbations:
(5) $\mathrm{AdvRob}_{\epsilon}(\theta)=1-\mathbb{E}_{(\bm{x},y)\sim D_{l}}\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$
where $\mathcal{L}_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in the following sections, we design different data selection schemes with respect to the extra unlabeled dataset for SSAT that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
4. Proposed Data Selection Schemes
So far, we have introduced the problem task of reducing the unlabeled dataset size to improve the efficiency of SSAT algorithms while maintaining the robustness enhancement and illustrated its importance. Motivated by the significant role of boundary-adjacent data points in optimizing model performance (Section 4.1), in this section, we design three efficient data selection schemes to address the optimization problem in Equation 4 (Sections 4.2 and 4.3).
4.1. Prioritize Boundary Unlabeled Data
Achieving high model robustness with restricted data resources remains a challenging task in machine learning. It involves striking a balance among data significance, optimizing the use of labeled and unlabeled data, and tackling constraints imposed by model capacity. Inspired by the prior literature (Zhang et al., 2020; Hua et al., 2021) that emphasizes the imbalanced data importance for vanilla adversarial training (see Section 2.1 for detailed discussions of these works), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose identifying a small set of vulnerable yet valuable unlabeled data points in $\mathcal{S}_{u}$, which are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield more robustness enhancement to adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy for locating difficult-to-classify data points. Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points. SSAT algorithms can then leverage this information to ensure these boundary-adjacent points are correctly classified, even under adversarial conditions. By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies (see Figure 6 in Appendix A for an illustration of the overall training pipeline).
While a straightforward extension of existing approaches is to select unlabeled data based on how much their predictions change under perturbations found by PGD attacks, this method is computationally expensive, undermining the efficiency gains we seek. Specifically, these methods entail iterative optimization processes to pinpoint boundary points, making them less efficient, especially for large-scale datasets. Additionally, while these methods provide valuable insights, their complexity often limits their interpretability. Therefore, we seek alternatives to effectively identify boundary data points without incurring high computational overhead.
Algorithm 1 Prediction Confidence-based Selection
1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
4:  Initialize $\mathcal{A}_{u}$ as empty set
5:  $\mathrm{Conf}(\bm{x})\leftarrow$ get confidence of $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
7:  $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with lowest $\mathrm{Conf}(\bm{x})$
8:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
10:  Output: selected subset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
4.2. Prediction Confidence-Based Selection
To identify the critical set of vulnerable data points near the model’s decision boundary while accounting for computational efficiency, scalability, and interpretability, we first propose a straightforward approach, termed as Prediction Confidence-based Selection (PCS), which utilizes the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. The pseudocode for such a selection scheme is depicted in Algorithm 1. Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section 6.2.
The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data using model confidence scores. Note that for self-supervised adversarial training, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset. Thus, an efficient data selection scheme is desirable to avoid high computational overhead. Nevertheless, we discovered in our experiments that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT. We hypothesize that PCS overlooks the geometric relationships and distributional properties, which are crucial for characterizing boundary-adjacent points, particularly for complex datasets. In addition, DNNs have been shown to be overconfident in their predictions (Guo et al., 2017), suggesting prediction confidence score might be an inherently biased indicator.
4.3. Latent Clustering-Based Selection
To overcome the above issues, we propose latent clustering-based selection (LCS) strategies, which identify data points near the model’s decision boundary in the latent space using different clustering techniques. Our approach begins by generating latent embeddings for all unlabeled data $\bm{z}=h_{\hat{\theta}}(\bm{x})$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$. The penultimate layer captures more abstract, high-level features and better represents the underlying data structure, which helps identify points near decision boundaries. It avoids the biases of overconfident predictions from the final layer, offering more reliable clustering. Here, the goal is to identify boundary-adjacent data points inferred by examining distances to cluster centroids. Points equidistant from multiple centroids are expected to be closer to decision boundaries in the latent embedding space.
In particular, we explore two classical clustering techniques for their unique benefits in identifying boundary points in the LCS framework. Algorithm 2 presents the pseudocode of the two proposed LCS methods. In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to capture local geometric structures around decision boundaries effectively. On the other hand, LCS with Gaussian mixture models (LCS-GMM) fits the latent representations to Gaussian mixture models, using posterior probabilities across multiple fitted Gaussians to identify points that are likely near decision boundaries. Compared to PCS, both LCS-KM and LCS-GMM leverage latent space clustering to provide a more accurate characterization of boundary vulnerabilities.
Algorithm 2 Latent Clustering-based Selection
1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
4:  Initialize $\mathcal{A}_{u}$ as empty set
5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in D_{u}$
6:  if using LCS-KM then
7:     $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}\leftarrow$ k-means clustering on $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
8:     $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in D_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
9:     $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta d$
10:  else if using LCS-GMM then
11:     Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
12:     Compute posterior probabilities $\bm{p}(\bm{z})$ for each $\bm{z}$
13:     $\Delta p\leftarrow$ compute $\Delta p=|p_{1}-p_{2}|$ for each $\bm{z}$, where $p_{1}$ and $p_{2}$ are the top two highest probabilities
14:     $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta\gamma$
15:  end if
16:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
18:  Output: selected dataset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
LCS-KM. In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the $N$ unlabeled data points into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster. For each latent embedding $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$. Data points are selected based on the minimal difference in distance between each latent embedding $\bm{z}$ to the corresponding two closest cluster centroids $\Delta d=|d_{1}-d_{2}|$, where $d_{1}$ and $d_{2}$ are the Euclidean distances to the closest and second closest centroids, respectively. The set of unlabeled inputs with the smallest $\Delta d$ values are selected. Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset followed by the same balancing step using the ratio parameter $\beta$. As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during self-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
LCS-GMM. In this variant of LCS, we again start by computing latent embeddings for unlabeled data $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$. Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models. A GMM assumes that the data is generated from a mixture of $k$ Gaussian distributions, each with its own mean $\bm{\mu}_{j}$ and covariance matrix $\mathbf{\Sigma}_{j}$ for the $j$-th component. Mathematically, each data point $\bm{z}$ has a probability of belonging to the $j$-th Gaussian component given by the posterior probability:
$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$
where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ represents the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$. We denote $\bm{p}(\bm{z})=[p_{1}(\bm{z}),p_{2}(\bm{z}),\ldots,p_{k}(\bm{z})]$ as the probability vector. To identify data points near the decision boundary, we focus on those points for which the posterior probabilities $\bm{p}(\bm{z})$ are similar across multiple Gaussian components, indicating that they are near the boundary between different clusters. Specifically, for each data point $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively. Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries. Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset. As will be illustrated in Section 5, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$). More detailed discussions of the difference between these data selection schemes are provided in Section 6.1, where we visualize the selected unlabeled data in a two-dimensional latent space.
5. Experiments
In this section, we comprehensively evaluate the performance of our proposed data selection schemes, as described in Algorithms 1 and 2, using two widely recognized image benchmarks: SVHN (Netzer et al., 2011) and CIFAR-10 (Alex, 2009) (Section 5.2). Following existing literature, we consider $\ell_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10. We also highlight the computational advantages of our methods, particularly their ability to reduce the convergence time of SSAT algorithms (Section 5.3). Furthermore, we assess the generalizability of our techniques by applying them to a real-world medical dataset, demonstrating the applicability of our methods beyond standard image benchmarks (Section 5.4).
Table 1. Comparison results of self-supervised adversarial training performance under different unlabeled data selection schemes and varying ratios on two image benchmarks. For SVHN, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ and train models on Wide ResNet 16-8. For CIFAR-10, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$ and train models on Wide ResNet 28-10.
Dataset Ratio $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
[…]
SVHN $0\%$ No Sel. $94.4$ $75.6$ $68.3$ $2.62$ $94.4$ $75.6$ $68.3$ $2.62$
$100\%$ No Sel. $97.1$ $86.0$ $75.3$ $8.38$ $97.4$ $86.3$ $75.2$ $17.41$
$1\%$ Random $94.2$ $78.4$ $70.3$ $3.28$ $95.1$ $80.3$ $70.5$ $3.28$
PCS $95.4$ $81.8$ $71.6$ $3.75$ $95.4$ $81.5$ $71.5$ $4.08$
LCS-GMM $95.1$ $81.8$ $72.1$ $3.97$ $95.6$ $81.2$ $71.8$ $4.27$
LCS-KM $95.2$ $82.7$ $72.9$ $3.90$ $96.4$ $82.4$ $72.6$ $4.18$
$10\%$ Random $95.3$ $82.0$ $73.0$ $3.28$ $95.6$ $83.6$ $72.4$ $3.28$
PCS $96.1$ $82.8$ $74.2$ $3.85$ $95.3$ $84.0$ $73.5$ $4.08$
LCS-GMM $96.2$ $83.0$ $74.3$ $3.97$ $95.5$ $84.1$ $74.1$ $4.27$
LCS-KM $96.1$ $86.3$ $75.2$ $3.90$ $96.6$ $86.6$ $74.8$ $4.18$
$20\%$ Random $96.2$ $82.3$ $73.2$ $3.28$ $96.3$ $84.6$ $74.0$ $3.28$
PCS $96.3$ $83.9$ $74.7$ $3.75$ $96.6$ $85.4$ $74.5$ $4.08$
LCS-GMM $96.4$ $84.3$ $75.0$ $3.97$ $95.8$ $85.5$ $74.7$ $4.27$
LCS-KM $96.2$ $86.6$ $75.1$ $3.90$ $96.7$ $87.2$ $75.3$ $4.18$
CIFAR-10 $0\%$ No Sel. $84.9$ $55.4$ $49.2$ $10.69$ $84.9$ $55.4$ $49.2$ $10.69$
$100\%$ No Sel. $89.7$ $62.5$ $58.6$ $28.50$ $85.7$ $61.8$ $58.4$ $57.13$
$1\%$ Random $83.2$ $54.5$ $49.8$ $14.25$ $83.3$ $54.3$ $49.6$ $14.25$
PCS $84.9$ $55.4$ $51.5$ $14.73$ $84.3$ $55.2$ $50.7$ $15.16$
LCS-GMM $84.2$ $55.1$ $52.1$ $14.89$ $85.8$ $54.6$ $51.8$ $15.36$
LCS-KM $85.6$ $56.4$ $52.9$ $14.78$ $85.9$ $56.1$ $52.5$ $15.27$
$10\%$ Random $85.2$ $56.0$ $52.4$ $14.25$ $85.0$ $56.5$ $52.6$ $14.25$
PCS $85.3$ $56.9$ $53.7$ $14.73$ $85.4$ $56.9$ $54.1$ $15.16$
LCS-GMM $85.9$ $57.1$ $54.2$ $14.89$ $85.8$ $57.1$ $54.5$ $15.36$
LCS-KM $87.2$ $58.2$ $55.3$ $14.78$ $86.1$ $58.0$ $55.8$ $15.27$
$20\%$ Random $87.1$ $57.5$ $54.2$ $14.25$ $85.4$ $57.2$ $54.2$ $14.25$
PCS $87.0$ $57.9$ $54.5$ $14.73$ $85.6$ $58.0$ $55.0$ $15.16$
LCS-GMM $87.0$ $58.2$ $55.5$ $14.89$ $85.9$ $58.6$ $55.8$ $15.36$
LCS-KM $88.7$ $60.7$ $57.8$ $14.78$ $85.5$ $60.2$ $57.2$ $15.27$
5.1. Experimental Settings
First, we introduce the necessary details to understand our experimental results. All the remaining details for reproducing our experiments are provided in Appendix B.
Dataset. For the image benchmark experiments, the proposed selection schemes are initially applied to external unlabeled datasets, following the implementation protocols outlined by Carmon et al. (2019). For the SVHN experiments, models are trained on $73$K labeled digit images from the original SVHN dataset, supplemented by $531$K additional unlabeled SVHN images. Specifically, the CIFAR-10 dataset comprises $50$K labeled training images and $10$K labeled test images, with models trained using SSAT augmented by an external unlabeled dataset $\mathcal{S}_{u}$ containing $500$K images sampled from the $80$M Tiny Images (80M-TI) dataset. To assess the generalizability of our selection schemes, we further evaluate SSAT algorithms using $1$M synthetically generated images created by the Denoising Diffusion Probabilistic Model (DDPM) for both CIFAR-10 and SVHN, following the protocols described by Gowal et al. (2021).
Configuration. We primarily conduct experiments on SSAT using WideResNet architectures. We train models using TRADES (Zhang et al., 2019) on both labeled and unlabeled data coupled with pseudo labels generated by the intermediate model. In Section 6.3, we evaluate the sensitivity of our methods for varying $\ell_{\infty}$ perturbation size, PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations. To standardize training, an epoch is defined as processing $50$K data points, regardless of the total dataset size. This is achieved by calculating the required number of batches to cover $50$K data points based on the batch size. This approach, adopted in prior works (Carmon et al., 2019; Gowal et al., 2021), ensures that training time per epoch remains consistent across datasets of varying sizes.
SSAT. For experiments with external unlabeled data, SSAT models are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs. To evaluate the performance of each data selection scheme, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the “best” model. This approach aligns with the early stopping practices commonly used in adversarial ML literature (Zhang et al., 2019; Rice et al., 2020). In experiments with external data and selection ratios $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN. Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs. Total training time is reported as the duration required to achieve the best model performance.
Intermediate Model. As a foundational step, the proposed selection schemes require training an intermediate model that utilizes the same architecture as the final model. This intermediate model is trained using standard supervised learning for $100$ epochs, with training on the WRN-28-10 architecture taking approximately $54$ minutes and $43$ seconds. Notably, this time is excluded from the reported training durations presented in the results table. The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for the implementation of any data selection strategy. As such, this training step is integral to the overall workflow, irrespective of whether a specific selection method is employed.
Evaluation Metric. To evaluate model robustness, we primarily focus on using multi-step PGD attacks (PGD) (Madry et al., 2017) due to their wide adoption in existing adversarial ML literature. We also report the robust accuracy using AutoAttack (AA) (Croce and Hein, 2020b) in our main experiments on image benchmarks for more rigorous evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
5.2. Main Evaluation on Image Benchmarks
SVHN. Our SVHN evaluation results, summarized in Table 1, demonstrate the efficacy of SSAT algorithms with unlabeled data using various selection schemes and $\alpha\in\{0\%,1\%,10\%,20\%,100\%\}$ trained on a WRN-16-8 architecture. Specifically, employing the LCS-KM selection strategy with $10\%$ or $20\%$ of the additional data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which closely approximate the $86.0\%$ robustness achieved using the entire unlabeled dataset. In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM. The robust accuracies evaluated using AutoAttack (i.e., the “AA” column of Table 1) demonstrate similar patterns compared to that using multi-step PGD Attacks (i.e., the “PGD” column of Table 1). Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, reinforcing the robustness and adaptability of the proposed data selection framework. Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases. We provide detailed analyses and discussions of these ratio hyperparameter configurations in Section 6.2.
We also note that the use of ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance improvements stem from the proposed data selection strategy rather than label accuracy (see Table 4 in Appendix D.2 for details). Additionally, experiments with $1$M data points generated using DDPM for SSAT (Gowal et al., 2021) highlight that selecting only $10\%$$20\%$ of data through LCS-KM achieves comparable performance to utilizing the full unlabeled dataset, with LCS-KM consistently demonstrating superior results. Importantly, training SSAT with selected data requires only $75$ epochs to optimize robust accuracy, whereas training with the full dataset demands up to $400$ epochs. This significant reduction in computational cost further underscores the practical advantages of the proposed selection methodology, which will be elaborated in Section 5.3.
CIFAR-10. For CIFAR-10 experiments, we adopt the WRN-28-10 architecture to train models using self-supervised adversarial training algorithms. To assess the generalizability of our observations, we also conduct ablation studies on a ResNet-18 architecture, which are detailed in Table 5 in Appendix D.4). The impact of varying proportions, $\alpha$, of pseudo-labeled unlabeled data was analyzed, revealing that selecting $20\%$ of the external data using LCS-KM achieved a robust accuracy of $60.7\%$, comparable to the $61.2\%$ obtained with the entire dataset. In contrast, the random selection of the same proportion yielded only $57.8\%$, underscoring the advantages of strategic data selection techniques such as LCS-KM. Building on the SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data. Using $1$ million images generated by DDPM (Gowal et al., 2021), we discover that utilizing just $20\%$ of the generated data via LCS-KM resulted in performance metrics nearly identical to those achieved with the full dataset. These trends are consistent across experiments with $\ell_{2}$ perturbations (see Figure LABEL:fig:two_graphs_ab for evidence).
Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations. In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection. When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training. By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance. These results validate the efficacy of our latent clustering-based selection approach in enhancing self-supervised adversarial training (SSAT). The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in adversarial training frameworks. Table 3 in Appendix C summarizes the comparisons, showcasing that LCS-KM improves model robustness while significantly improving the efficiency of SSAT algorithms with full unlabeled data.
5.3. Computational Benefits
As suggested by the runtime comparisons depicted in Table 1, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall computation costs of SSAT while maintaining competitive, robust accuracy. This section further visualizes the learning curves of SSAT and discusses the computational benefits associated with our proposed methods.
Figure LABEL:fig:convergence depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing $50$K data points, with $10$ steps for PGD training and $20$ for PGD testing. When no additional data is utilized, as shown in Figure LABEL:fig:0per, the highest robust test accuracy is achieved after approximately $75$ epochs. As illustrated in Figures LABEL:fig:10percrandom and LABEL:fig:10perc, incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs. In contrast, as depicted in Figure LABEL:fig:100perc, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs. These findings suggest that including extra unlabeled data in SSAT increases computational demands, as evidenced by the longer training convergence time. Strategic selection reduces redundancy in the dataset, allowing the model to achieve optimal performance with fewer training epochs.
Specifically, we observe that the size of the selected unlabeled dataset plays a critical role in determining the time required for the model to achieve peak accuracy. Larger unlabeled datasets usually extend the convergence times of SSAT, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. However, continued training beyond this point often leads to robust overfitting. Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations. Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time. These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
5.4. Application to Real-World Medical Data
We extend our study to a real-world healthcare application with COVID-19-related datasets. Specifically, the COVIDGR dataset (Tabik et al., 2020) is utilized as the labeled dataset, and the CoronaHack dataset (Govi, 2020) serves as the source of unlabeled data. This extension allowed us to validate the efficacy of the proposed methods in a domain where data scarcity and model robustness are critical considerations.
Setting. The experimental setup consists of multiple stages, leveraging both datasets in conjunction with ResNet-18 as the base model. The COVIDGR dataset contains $852$ chest X-ray images evenly distributed across two classes: $426$ COVID-19 positive cases (confirmed via RT-PCR testing) and $426$ negative cases. This dataset was split into an $80\%$ training set and a $20\%$ test set for evaluation purposes. The CoronaHack dataset, which comprises $5,910$ chest X-ray images, is divided into an $89\%$ training set and an $11\%$ test set. This dataset has an imbalanced class distribution, with $73\%$ of the images classified as COVID-19 positive and $27\%$ as negative (non-COVID). For our experiments, we strategically select a subset of the training data from the CoronaHack dataset as unlabeled data to assess the effectiveness of our adversarial robustness techniques. We applied the LCS-KM selection strategy to identify and incorporate a minimal yet representative subset of unlabeled data from the CoronaHack dataset into the training process. This approach enabled us to maintain robust model performance while significantly reducing the volume of external data required. The models are trained on SSAT with varying unlabeled data selection schemes using a ResNet-18 architecture. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and vanilla PGD-based adversarial training.
Result Analysis. The results of our experiments confirm the effectiveness of strategic data selection in enhancing robust accuracy and optimizing training efficiency. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieved a robust accuracy of $56\%$, compared to $53\%$ when the same proportion was selected randomly, as shown in Figures LABEL:fig:10perrandom and LABEL:fig:10per. Additionally, smaller, strategically selected datasets led to faster convergence, underscoring the benefits of early stopping. Specifically, the best accuracy of $56\%$ was reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM in Figure LABEL:fig:10per. Figure LABEL:fig:20per shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset (i.e., $\alpha=100\%$) required $60$ epochs to achieve the highest robust accuracy of $58\%$ depicted in Figure LABEL:fig:100per. These findings demonstrate that smaller but strategically selected subsets of data can achieve robust accuracy comparable to that of the full dataset while significantly reducing computational time and memory requirements. The success of our approach in this real-world application highlights its practical relevance, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
6. Additional Analyses
In this section, we conduct additional analyses and experiments to assess the generalizability of our proposed method. We start by visualizing and comparing the latent-space behavior of PCS, LCS-KM, and LCS-GMM to evaluate their effectiveness in selecting critical data points (Section 6.1). Next, we explore the influence of key ratio hyperparameters in our data selection schemes on the robustness performance achieved by SSAT (Section 6.2). Additionally, we evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section 6.3). Finally, we examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section 6.4). Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table 1.
6.1. Comparative Visualizations
To facilitate a comprehensive understanding of the proposed data selection schemes, we perform a comparative visualization using the TinyImages dataset in the latent space. The latent representations, extracted from the penultimate layer of a pre-trained WRN-28-10 model on CIFAR-10, encode high-dimensional feature embeddings that capture the data’s underlying structure. To render these high-dimensional features interpretable, we employ a two-step dimensionality reduction process that balances the retention of global structure and the visualization of local relationships.
Setting. The first step in this two-step pipeline is Principal Component Analysis (PCA). By reducing the dimensionality of the latent features to $50$ components, PCA retains the majority of the variance within the data. This step ensures that the global structure of the feature space is preserved, maintaining the relative distances and relationships that reflect the overall organization of the data (Gewers et al., 2021). PCA also acts as a preprocessing step, making the subsequent visualization process more efficient and effective. In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) to project the $50$-dimensional PCA-reduced features into a two-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE is designed to maintain local relationships among data points (Xia et al., 2021). This combination of PCA and t-SNE allows us to visualize the latent feature space in a manner that is both interpretable and faithful to the original high-dimensional structure.
After deciding on the two-dimensional latent space, we visualize the selected unlabeled data using our proposed schemes: PCS, LCS-KM, and LCS-GMM. To ensure methodological rigor and consistency, we strictly adhere to the procedures outlined in Algorithms 1 and 2 for implementing the data selection techniques. The ratio hyperparameters $\alpha$ and $\beta$ are set to $0.1$ and $1$, respectively. Importantly, these parameters influence the specific distribution of selected points but do not alter the overall interpretability of the visualizations, ensuring that the comparative analysis remains robust by their precise values.
PCS. Figure LABEL:fig:clustering_techniques presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy. Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT across various experimental settings. In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes. Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers or ambiguous samples that offer limited benefit for refining the decision boundary. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points, as shown in Figure LABEL:fig:PCS_selection. This amplifies PCS’s shortcomings compared to the more structured sampling of LCS-KM or the probabilistic clustering offered by LCS-GMM. Thus, PCS often dedicates training efforts to less informative or highly noisy examples, resulting in weaker robustness improvements.
LCS. LCS-GMM relies on fitting Gaussian distributions to the latent representations. Although this probabilistic approach can effectively capture regions where classes overlap, it inherently presupposes that local data clusters follow (approximately) Gaussian distributions, an assumption that may fail in real-world settings characterized by irregular or noisy data. As a result, LCS-GMM’s contours can become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Steinley and Brusco, 2011). A closer look at the selected boundary samples in Figure LABEL:fig:gmm also reveals that the region these points occupy is broader and less tightly focused on the decision boundary, making it harder to capture truly critical examples for training. Due to the concentration, selected data points overlap with each other in the two-dimensional visualization space. In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM, further underscoring the latter’s advantage in both coverage and precision.
Finally, our best-performing LCS-KM method emphasizes leveraging the latent-space clustering structure to identify critical decision-boundary regions in a more balanced manner. By partitioning the data representations into multiple k-means clusters, LCS-KM systematically locates the points of lowest classification confidence within each cluster, illustrated in Figure LABEL:fig:kmeans, thereby ensuring that the algorithm focuses on diverse boundary-adjacent regions. This cluster-based approach yields a fine-grained analysis of local uncertainties and avoids the risk of overcommitting to isolated, potentially noisy areas in the latent space.
6.2. Hyperparameter Sensitivity
The efficacy of our selection strategies critically relies on several key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$. We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally. Due to page limits, experimental results with varying values of $\gamma$ are deferred to Appendix D.1.
Ratio of Boundary Points $\bm{\beta}$. Concentrating exclusively on points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away. The proportion of these selections is determined by the value of $\beta$. We conduct experiments with different $\beta$ values to assess their impact on robust accuracy. Figure LABEL:fig:betacifar illustrates how robust accuracy varies with $\beta$ on the CIFAR-10 dataset using $20\%$ generated data from DDPM, which is approximately four times the size of the original dataset. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure LABEL:fig:betasvhn illustrates how robust accuracy varies with $\beta$ on the SVHN dataset using $10\%$ extra data from the same dataset. Here, due to the smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. Consequently, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, a smaller $\beta$, i.e., fewer points near the boundary, may yield better robust accuracy. Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.
Ratio of Selected Unlabeled Data $\bm{\alpha}$. Figures LABEL:fig:cofarper and LABEL:fig:svhnper illustrate how the performance varies with selection ratio $\alpha$ for CIFAR-10 and SVHN, respectively. We analyze how the quantity of extra data influences the effectiveness of our data-selection method, which is designed to prioritize data points located near decision boundaries. When the external data supply is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of external or generated data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
Table 2. Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
[…]
$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
pre-trained CLIP $86.8$ $83.9$ $40.5$
$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
pre-trained CLIP $86.8$ $85.3$ $43.2$
$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
pre-trained CLIP $86.8$ $85.6$ $47.7$
100% fully-supervised $90.5$ $86.2$ $58.2$
pre-trained CLIP $86.8$ $86.5$ $57.4$
6.3. Further Exploration of Generalizability
All experiments in this section are conducted on CIFAR-10. We compare four different settings: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data (i.e., $\alpha=0\%$), and using the entire dataset (i.e., $\alpha=100\%$).
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table 1. The results are depicted in Figures LABEL:fig:4_255 and LABEL:fig:16_255, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data. The trends are shown to be consistent across varying perturbation sizes.
Adversarial Training. We explore the impact of replacing TRADES with vanilla AT as the training algorithm. Similarly, we use external data from TinyImages as the unlabeled dataset. The results for PGD-based AT are shown in Figure LABEL:fig:-pgd. Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure LABEL:fig:pgd-5. Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are shown in LABEL:fig:externall2 and LABEL:fig:generatedl2, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in self-supervised adversarial training algorithms.
6.4. Low Labeled Data Regime
We further study how the choice of the intermediate model and the amount of labeled data impact the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with 10% of data selected from a pool of $500$K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low. Additionally, we include a pre-trained CLIP model (Radford et al., 2021) as an alternative intermediate model for comparison. The results of these experiments are presented in Table 2. The intermediate model’s quality significantly impacts the final model’s robust accuracy. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations underscore the importance of having a strong intermediate model for robust performance. Additionally, we find that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-label regimes. Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix D.3.
6.5. Further Discussion
Our study introduces strategic data selection methods, including Latent Clustering-Based Selection with k-means (LCS-KM), Latent Clustering-Based Selection with Gaussian Mixture Models (LCS-GMM), and prediction confidence-based selection, to improve the efficiency of self-supervised adversarial training. These approaches tackle key challenges in adversarially robust learning, particularly the significant memory and computational overhead of existing methods. By prioritizing boundary-adjacent data, these processes aim to enhance model robustness. However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models. While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset. By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization. Nevertheless, deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization. Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
7. Conclusion
In this work, we illustrated the importance of latent clustering-based data selection in improving the efficiency of SSAT while maintaining robust accuracy. Emphasizing boundary data points during selection significantly enhances the model’s overall robustness. For SVHN, we also showed improvement in robust accuracy by adding ground-truth labels to a small set of data. This suggests manually annotating a few unlabeled data points close to the boundary could be a potential future direction to further enhance model robustness with considerations of memory and computational efficiency. Our work opens new avenues for further research, including exploring advanced data selection schemes, making the selection process more interpretable, and investigating the trade-offs between model performance, data selection, and computational costs.
Appendix A Overall Training Pipeline
Figure 6 illustrates the overall pipeline, based on the robust self-training framework proposed by Carmon et al. (Carmon et al., 2019). We first train our intermediate model using the available labeled data and use this intermediate model to select a subsection of our unlabelled or generated data. The final robust training is performed on all the labeled data and a strategically selected subset of the unlabelled or generated data.
Figure 6. Overview of the training pipeline. An intermediate model is first trained with labeled data and then used to select data from the unlabeled set. Finally, we adopt SSAT to learn a robust model on both the labeled and the reduced unlabeled datasets.
Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately $73$K images and an extra training set of around $531$K images. Initially, the model is trained on the $73$K labeled images. To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model (Gowal et al., 2021).
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture (Zagoruyko and Komodakis, 2016) in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. (Zhang et al., 2019), with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. (Carmon et al., 2019) except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with $1$M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al. (Carmon et al., 2019) for better comparison. We use step size $0.005$, number of attack steps $K=100$. We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in $80$M-TI do not correspond to CIFAR-10 image categories. Carmon et al. (Carmon et al., 2019) used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional $50$K images from 80M-TI using the model’s predicted scores to create a $500$K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model with $50$K labeled data and use this model to select data from the $500$K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments utilizing $1$M synthetical data generated by the DDPM model (Gowal et al., 2021).
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) (Zagoruyko and Komodakis, 2016) architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section 6.3 when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. (Carmon et al., 2019) except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the $500$K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process. Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. (Carmon et al., 2019) for fair comparisons. We use step size $\alpha=0.01$, number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
Medical Application. Below, we detail the setup employed for our experiments on the medical datasets outlined in Section 5.4. For this experiment, we trained models using a ResNet-18 (He et al., 2016) architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained with SSAT using the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Table 3. Comparisons of SSAT methods with different selection schemes based on average efficiency score and robust accuracy. Here, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the “No Selection” method, we use $\alpha=100\%$, taking the average of the external and generated data results for each setting.
[…]
SVHN CIFAR-10
Random $1.00$ $83.1\pm 1.04$ $56.8\pm 0.59$
PCS $0.89$ $84.0\pm 0.92$ $57.4\pm 0.53$
LCS-GMM $0.87$ $84.2\pm 0.89$ $57.8\pm 0.67$
LCS-KM $0.88$ $86.6\pm 0.33$ $59.3\pm 1.15$
No Selection $0.33$ $86.2\pm 0.15$ $62.2\pm 0.35$
Appendix C Detailed Comparisons of Different Data Selection Schemes
This section presents a comprehensive comparison of all the selection methods, including the case where no selection is applied. It highlights the tradeoff between performance and efficiency, offering a detailed analysis of the conditions under which each method is most suitable. Table 3 summarizes the performance of SSAT with different data selection strategies in terms of efficiency and robust accuracy. Each method strikes a balance between computational cost and accuracy improvements. Random selection serves as a baseline with minimal computation but lacks precision, while our strategic selection methods prioritize boundary points, which achieve progressively improved performance but with moderate increases in computational demands. Using $100\%$ of the data without selection yields the highest robust accuracy but at a significant computational cost. More specifically, we calculate a normalized efficiency score for each method defined as $\text{Eff Score}_{ij}=\frac{T_{\text{min},j}}{T_{ij}}$, where $T_{\text{min},j}$ is the minimum training time for dataset $j$, and $T_{ij}$ is the actual training time for method $i$. The overall efficiency score for method $i$ is the average across all datasets computed as $\frac{1}{n}\sum_{j}\text{Eff Score}_{ij}$, where $n$ denotes the number of datasets and sources considered. Effectiveness is measured by the average robust accuracy across CIFAR-10 and SVHN. In particular, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the no-selection method, we use $\alpha=100\%$. In all cases, we take the average of the results from both the external and generated data. These metrics are introduced to provide a fair comparison between methods.
Appendix D Additional Experiments
D.1. Other Hyperparameter Analysis Details and Results
In Section 6.2, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. In these experiments, we varied one hyperparameter at a time while keeping the others constant. Figure LABEL:fig:betacifar illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the $1$M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure LABEL:fig:betasvhn shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$. In Figure LABEL:fig:ratiocifar, we present how robust accuracy for the CIFAR-10 dataset is affected by varying the ratio of per-batch ratio $\gamma$ when $20\%$ extra data from the $1$M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure LABEL:fig:ratiosvhn displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$. In addition, Figures LABEL:fig:cofarper and LABEL:fig:svhnper explore how robust accuracy varies with the amount of data used. For CIFAR-10 in Figure LABEL:fig:cofarper, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$. For the SVHN dataset in Figure LABEL:fig:svhnper, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures LABEL:fig:cofarper and LABEL:fig:svhnper, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
Table 4. Comparison results of self-supervised adversarial training algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned with the ground-truth SVHN labels provided in the dataset.
[…]
$1\%$ LCS-KM $95.6$ $82.9$ $75$
$10\%$ Random $96.3$ $82.7$ $75$
$10\%$ LCS-KM $96.7$ $86.3$ $75$
$20\%$ LCS-KM $96.4$ $86.5$ $75$
$100\%$ No Selection $97.5$ $86.4$ $200$
D.2. SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table 4 shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
D.3. Pre-Trained CLIP as Intermediate Model
In all our previous experiments, we began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. This intermediate model was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with generated or external data that already comes with labels, training an intermediate model specifically for selection undermines the selection process’s intended benefits, primarily to reduce computational and time complexity. In addition, in application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings (Radford et al., 2021). CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution allows us to select a relevant subset without having to train an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure LABEL:fig:comparison_clip shows our result when using CLIP instead of the previous intermediate model for data selection.
D.4. Resnet-18 Model Architecture
Our main experiments focus on the WidResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on $50$K labeled samples from CIFAR-10 and a chosen subset of data from $500$K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table 5, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
Table 5. Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of $50$K labeled CIFAR-10 images and $500$K unlabeled images drawn from Tiny-ImageNet.
[…]
$1\%$ Random $79.8$ $50.6$
PCS $79.4$ $51.1$
LCS-GMM $78.7$ $51.6$
LCS-KM $78.1$ $51.8$
$10\%$ Random $82.3$ $52.0$
PCS $82.1$ $52.6$
LCS-GMM $82.0$ $53.2$
LCS-KM $83.3$ $53.6$
$20\%$ Random $82.4$ $53.1$
PCS $82.1$ $53.3$
LCS-GMM $82.1$ $53.7$
LCS-KM $83.1$ $54.0$
$100\%$ No Selection $83.2$ $54.7$

### v2

@@
-Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection
-Abstract.
-Compared with standard learning, adversarially robust learning is widely recognized to demand significantly more training examples.
-Recent works propose the use of self-supervised adversarial training (SSAT) with external or synthetically generated unlabeled data to enhance model robustness.
-However, SSAT requires a substantial amount of extra unlabeled data, significantly increasing memory usage and model training times.
-To address these challenges, we propose novel methods to strategically select a small subset of unlabeled data essential for SSAT and robustness improvement.
-Our selection prioritizes data points near the model’s decision boundary based on latent clustering-based techniques, efficiently identifying a critical subset of unlabeled data with a higher concentration of boundary-adjacent points.
-While focusing on near-boundary data, our methods are designed to maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting.
-Our experiments on image benchmarks show that integrating our selection strategies into self-supervised adversarial training can largely reduce memory and computational requirements while achieving high model robustness.
-In particular, our latent clustering-based selection method with k-means is the most effective, achieving nearly identical test-time robust accuracies with $5$ to $10$ times less external or generated unlabeled data when applied to image benchmarks.
-Additionally, we validate the generalizability of our approach across various application scenarios, including a real-world medical dataset for COVID-19 chest X-ray classification.
-Our Implementations are available as open-source code at this url.
-1. Introduction
-Over the past decade, it has been repeatedly confirmed that deep neural networks (DNNs) are highly vulnerable to adversarial perturbations (Szegedy et al., 2013).
-This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness (Goodfellow et al., 2014; Papernot et al., 2016; Buckman et al., 2018; Biggio and Roli, 2018).
-Among them, adversarial training is regarded as one of the most effective methods (Madry et al., 2017; Wang et al., 2019; Zhang et al., 2019).
-However, as stated by Schmidt et al.
-(Schmidt et al., 2018), learning a model that is resilient to adversarial perturbations requires a significantly larger dataset than standard learning.
-Recent studies have explored self-supervised techniques to expand the training set size of adversarial training algorithms by leveraging unlabeled external (Carmon et al., 2019) or generated data (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023).
-Despite alleviating the sample complexity barrier and producing models with improved robust accuracies, these methods typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store those data and a much longer training time for algorithms like adversarial training to converge.
-Witnessing the challenges of additional memory and computational requirements, we investigate whether the significantly large amount of utilized additional data is inevitable for achieving state-of-the-art adversarial robustness.
-The ultimate goal of our work is to maximize the model robustness achieved by SSAT algorithms by using as few additional unlabeled data points as possible.
+Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction
+Abstract
+Achieving high model robustness under adversarial settings is widely recognized as demanding considerable training samples.
+Recent works propose semi-supervised adversarial training (SSAT) methods with external unlabeled or synthetically generated data, which are the current state-of-the-art.
+However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage.
+In this paper, we propose unlabeled data reduction strategies to improve the efficiency of SSAT.
+Specifically, we design novel latent clustering-based techniques to select or generate a small critical subset of data samples near the model’s decision boundary.
+While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting.
+Comprehensive experiments on benchmark datasets demonstrate that our methods can significantly reduce SSAT’s data requirement and computation costs while preserving its strong robustness advantages.
+In particular, our latent-space selection scheme based on k-means clustering and our guided DDPM fine-tuning approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data and approximately $4\times$ less total runtime.
+I Introduction
+Deep neural networks (DNNs) are highly vulnerable to adversarially perturbed inputs [1, 2].
+This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness [2, 3, 4, 5].
+Among them, adversarial-training-based algorithms are most popular [7, 8, 9].
+As illustrated by Schmidt et al.
+[10], adversarially robust learning is inherently challenging, requiring a much larger sample complexity than standard learning.
+To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [11, 12] or generated synthetic data [13, 14, 15].
+Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store the extra data and a much longer training time for algorithms like adversarial training to converge (see Appendix A for detailed discussions on related work).
+Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy.
+The ultimate goal is to maximize model robustness while utilizing as few additional data samples as possible.
 Inspired by Zhang et al.
-(Zhang et al., 2020), which highlights the unequal importance of training examples, we argue that with limited model capacity, self-supervised adversarial learning should also focus on optimizing critical data samples near the model’s decision boundary.
-Consequently, we propose multiple data selection strategies, including a simple prediction confidence-based selection (PCS) scheme that prioritizes unlabeled data points in which the model’s prediction is most uncertain and more advanced latent clustering-based selection (LCS) schemes.
-Specifically, we propose two LCS approaches depending on the corresponding clustering technique it builds on: latent clustering-based selection with Gaussian mixture models (LCS-GMM) and latent clustering-based selection with k-means (LCS-KM).
-For each unlabeled data point, we compute the difference between the highest and second-highest posterior probabilities in the latent feature space for LCS-GMM, while we select vulnerable points by calculating the distance difference between the two nearest cluster centroids for LCS-KM.
-The set of unlabeled data with the smallest probability or distance difference is considered more critical and further incorporated into self-supervised adversarial training.
-Moreover, our selection strategies balance the ratio of boundary critical points and the remaining non-boundary points to avoid undesirable overfitting to the boundary data distribution.
-By applying these targeted data reduction approaches, we streamline the self-supervised adversarial training process, significantly reducing the required additional unlabeled data while maintaining comparable model robustness against adversarial perturbations.
-Our work offers valuable insights into how to acquire additional data and employ SSAT in resource-constrained application domains.
-Below, we further summarize the main contributions of our work.
+[16], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary.
+Specifically, we introduce three data selection strategies, including PCS, a simple approach that prioritizes unlabeled data points with low prediction confidence, and two advanced latent clustering-based selection schemes: LCS-KM and LCS-GMM, where we identify the most uncertain data in the model’s latent embedding space based on the corresponding clustering technique.
+The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset.
+In addition, we propose novel DDPM fine-tuning approaches with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models.
+Our data reduction schemes balance the ratio of boundary critical points and the remaining non-boundary points to avoid potential overfitting to the boundary distribution.
+By applying the aforementioned data reduction schemes, the high unlabeled data and computation requirements of SSAT can be significantly alleviated while largely preserving its robustness performance.
 Contributions.
-We motivate and formalize the problem task of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section 3).
-To realize such a goal, we propose various data selection schemes to identify the most critical unlabeled data points, including a straightforward approach based on prediction confidence and two advanced methods based on clustering techniques of unlabeled data in the latent embedding space (Section 4).
-The proposed methods optimize the performance of SSAT by refining the model’s decision boundary in the input regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points.
-By focusing on critical unlabeled data points, our methods largely reduce the memory consumption and time complexities of self-supervised adversarial training algorithms.
-We conduct comprehensive experiments on two image benchmarks and a real-world medical dataset, demonstrating that our data selection schemes significantly reduce the memory consumption and the total running time of self-supervised adversarial training algorithms under various experimental settings (Section 5).
-Additionally, we explain the insights of the potential reasoning behind the superior efficacy of our LCS-KM method to other data selection schemes (Section 6.1) and provide general guidelines on how to employ our LCS methods for maximized robustness performance (Section 6.2).
-Finally, we examine the robustness of our methods by varying the adversarial training algorithms, perturbation size and metric (Section 6.3), as well as studying the role of the intermediate model under the regime with insufficient labeled data (Section 6.4), all confirming the usefulness and generalizability of the proposed data selection schemes for improving the efficiency of SSAT algorithms.
-We believe our work can help promote the development of more scalable, robust learning algorithms, which would be particularly beneficial for resource-constrained environments.
-2. Related Work
-In this section, we review the most relevant literature on adversarial examples, (self-supervised) adversarial training, and its variants.
-We also discuss the most representative techniques for improving the efficiency of standard deep learning, including dataset distillation and coreset selection.
-2.1. Adversarially Robust Learning
-Adversarial Examples.
-Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions (Szegedy et al., 2013).
-The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications.
-Earlier-proposed attacks like fast gradient sign method (FGSM) (Goodfellow et al., 2014) and projected gradient descent (PGD) (Madry et al., 2017) introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks (Carlini and Wagner, 2017), which search for the minimal perturbation that causes misclassification.
-Recent attacks like Auto-PGD (Croce and Hein, 2020a) automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack (Croce et al., 2020) is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
-Adversarial Training.
-On the defense side, adversarial training and its variants are the most popular learning methods for adversarial robustness.
-For instance, Madry et al.
-proposed to train DNNs on adversarially perturbed inputs using PGD attacks (Madry et al., 2017), while Zhang et al.
-introduced TRADES (Zhang et al., 2019), which balances the trade-off between adversarial robustness and standard accuracy.
-Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications.
-In particular, Rice et al.
-illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training (Rice et al., 2020), where early stopping is shown to be effective in alleviating such an issue.
-Techniques, such as Free AT (Shafahi et al., 2019) and Fast AT (Wong et al., 2020), offer promising solutions to reduce the computation for adversarial training.
-Nevertheless, they suffer from decreased robustness performance.
-Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently.
-For example, Zhang et al.
-argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations (Zhang et al., 2020).
-As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data.
-In a similar vein, Hua et al.
-advocated for a more targeted approach (Hua et al., 2021), recommending that PGD training be applied primarily to examples situated near the decision boundaries.
-This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks.
-Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
-Self-Supervised Adversarial Training.
-Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research (Chapelle et al., 2009).
-In the adversarial context, Schmidt et al.
-(Schmidt et al., 2018) highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning.
-They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly.
-Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training (Carmon et al., 2019; Alayrac et al., 2019; Najafi et al., 2019; Zhai et al., 2019).
-In particular, Carmon et al.
-proposed robust self-training methods to address the sample complexity issue (Carmon et al., 2019), which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data.
-The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model.
-Concurrently, Najafi et al.
-proposed an optimization framework incorporating labeled and unlabeled data (Najafi et al., 2019), while providing formal guarantees on adversarial robustness.
-In addition, leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023).
-For instance, Gowal et al.
-employed various unconditional generative models (Gowal et al., 2021), including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further.
-While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement.
-In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
-2.2. Data-Efficient Deep Learning
-Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information.
-Two prominent approaches are dataset distillation and coreset selection.
-Dataset Distillation.
-Dataset distillation compresses large datasets into compact, synthetic datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points.
-For instance, Wang et al.
-proposed a foundational approach to dataset distillation (Wang et al., 2018), proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance.
-Later, Zhao et al.
-introduced the concept of gradient matching (Zhao et al., 2020), where distilled data points are optimized to match the gradients of the original dataset.
-This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data.
-Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data (Sucholutsky and Schonlau, 2021), facilitating the training of complex models with minimal data while maintaining strong performance.
-Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands.
-Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
-Coreset Selection.
-On the other hand, coreset selection (Sener and Savarese, 2017; Kim and Shin, 2022) aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset.
-These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning.
-For example, Kaushal et al.
-utilized diverse models to select training data subsets to reduce labeling efforts (Kaushal et al., 2019), while Xia et al.
-introduced the concept of “moderate coreset” (Xia et al., 2022), focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions.
-Similarly, Mirzasoleiman et al.
-proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function (Mirzasoleiman et al., 2020).
-In the adversarial context, Dolatabadi et al.
-proposed adversarial coreset selection (Dolatabadi et al., 2023), offering a task-specific solution to enhance the computational efficiency of adversarial training.
-In particular, the method constructs a compact data subset by minimizing the gradient approximation error.
-The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals.
-While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges.
-These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time.
-In such settings, repeatedly selecting data during training may become computationally infeasible.
-In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
-3. Improving Data Efficiency for SSAT
-In this section, we first introduce the necessary notations and definitions for readers to understand self-supervised adversarial training (Section 3.1) and then motivate and formulate the problem task of enhancing the data efficiency of SSAT algorithms (Section 3.3).
-3.1. Preliminaries on SSAT
-We work with self-supervised adversarial training algorithms, initially introduced to the field by Carmon et al.
-(Carmon et al., 2019).
-Let $\mathcal{X}\subseteq\mathbb{R}^{n}$ be a $n$-dimensional input space, $\mathcal{Y}$ be the output space of class labels, and $D_{l}$ be the underlying labeled distribution supported on $\mathcal{X}\times\mathcal{Y}$ that we aim to learn an adversarially robust classifier.
-Let $\mathcal{S}_{l}$ be the training set with examples i.i.d.
-drawn from $D_{l}$ and $\mathcal{S}_{u}$ be a set of inputs sampled from some unlabeled distribution $D_{u}$ supported on $\mathcal{X}$.
-In the adversarial machine learning literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and $|\mathcal{S}_{u}|$ is set to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are easier to acquire than well-annotated labeled data.
-Throughout the paper, we use ||$\cdot$|| ⋅ | to denote the cardinality of the corresponding set.
-For instance, when $\mathcal{S}_{l}$ corresponds to the $50$K labeled CIFAR-10 training images, Carmon et al.
-(Carmon et al., 2019) selected $500$K “most CIFAR-10-like but non-identical” images from the whole $80$M Tiny ImageNet data as the unlabeled dataset $\mathcal{S}_{u}$ to be used by SSAT.
-To be more specific, SSAT first standardly trains an intermediate model $f_{\hat{\theta}}$ on $\mathcal{S}_{l}$, known as the pseudo labeling function, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$.
-More rigorously, the intermediate model’s weight parameters $\hat{\theta}$ are learned by minimizing the following objective function:
-(1) $\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}(\theta,\bm{x},y),$
-where $\mathcal{L}(\cdot)$ denotes the standard loss, such as cross-entropy, that measures the discrepancy between model prediction and the class label.
-Finally, SSAT trains a model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled dataset $\mathcal{S}_{l}$ and unlabeled dataset $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
-(2) $\min_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)+\frac{1}{\gamma|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\mathcal{L}_{\mathrm{adv}}\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$
-where $\gamma\geq 0$ controls the contributions between labeled and unlabeled data.
-In prior literature (Gowal et al., 2021), a typical approach is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\alpha$) but fix the total batch size to optimize for the best SSAT performance.
-For simplicity, we term $\gamma$ as the per-batch ratio hyperparameter in this paper.
-When $\frac{1}{\gamma}$ is set as 00, Equation 2 reduces to the training objective of vanilla adversarial training (AT) (Madry et al., 2017).
-In Equation 2, $\mathcal{L}_{\mathrm{adv}}(\cdot)$ stands for the adversarial loss function: for any $(\bm{x},y)\in\mathcal{S}_{l}$ and $\epsilon>0$,
-(3) $\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}(\theta,\bm{x}+\bm{\delta},y),$
-where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the ball centered at $\bm{0}$ with radius $\epsilon$ measured in some distance metric, such as $\ell_{p}$-norm.
-Aligned with prior literature on adversarial training (Madry et al., 2017), we adopt multi-step PGD to obtain an approximated solution to the inner maximization problem specified by Equation 3 during self-supervised adversarial training.
-3.2. Limitations of Existing SSAT Methods
-Although SSAT methods alleviate the sample complexity barriers of adversarially robust learning (Schmidt et al., 2018) and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement.
-This trend is clearly documented in the leaderboard of RobustBench (Croce et al., 2020).
-For example, state-of-the-art SSAT methods (Carmon et al., 2019; Gowal et al., 2021; Wang et al., 2023) either select $500$K external Tiny Imagenet data samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original $50$K CIFAR-10 training examples.
-For example, the method proposed by Gowal et al.
-(Gowal et al., 2021) achieves around $65\%$ robust accuracy on CIFAR-10 using a WideResNet-70-16 model architecture against $\ell_{\infty}$ perturbations with $\epsilon=8/255$, increasing the robustness performance of vanilla adversarial training (Madry et al., 2017) by a large margin, but relies an extra unlabeled set of $100$M DDPM-generated data.
-The huge unlabeled dataset required for SSAT significantly increases memory consumption for holding the whole training dataset, which we argue is inefficient and likely to be prohibited for resource-constrained application scenarios.
-These observations motivate us to explore whether such a considerable amount of unlabeled data can be reduced while preserving the high model robustness attained by SSAT algorithms.
-Moreover, we note that SSAT requires a much longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of training epochs compared with vanilla AT (see Figure LABEL:fig:convergence for supporting evidence).
-Intuitively, extracting useful and robust features through adversarial training from such a large and potentially diverse dataset is more difficult.
-As we will illustrate in our experiments, the slower convergence can be attributed to the additional large unlabeled datasets (with pseudo labels) that often exhibit higher variance than the original labeled data samples.
-Once we reduce the size of the unlabeled data involved with specifically designed techniques, faster convergence is expected for SSAT.
-Since vanilla adversarial training has already been criticized in prior literature (Shafahi et al., 2019; Wong et al., 2020) for its high computational costs for running multi-step PGD to solve the inner maximization problem in Equation 3, state-of-the-art SSAT algorithms incur even higher computational costs (see Table 1 for runtime comparisons), due to the slower convergence rate and larger model architecture they typically adopt.
-Therefore, it is essential to design efficient SSAT methods, especially for resource-constrained applications, that can better utilize the large amount of extra unlabeled data.
-3.3. Problem Formulation
-Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for training models with good robustness performance.
-Inspired by the idea of coreset selection for efficient deep learning (Kaushal et al., 2019; Xia et al., 2022; Mirzasoleiman et al., 2020), we propose to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that self-supervised adversarial training based on $\mathcal{S}_{l}$ and the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using full unlabeled dataset $\mathcal{S}_{u}$.
-More formally, we aim to solve the following constrained optimization problem:
-(4) $\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big{(}\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big{)},\text{ s.t.
-}|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$
-where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ stands for the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation 2, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint.
-$\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the robustness of the model with parameters $\theta$ against $\epsilon$ perturbations:
-(5) $\mathrm{AdvRob}_{\epsilon}(\theta)=1-\mathbb{E}_{(\bm{x},y)\sim D_{l}}\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$
-where $\mathcal{L}_{0/1}$ denotes the 0-1 loss function.
+We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section II).
+To realize this goal, we propose to optimize the performance of SSAT by refining the model’s decision boundary in the regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points (Section III-A).
+Specifically, we propose various data selection schemes to identify boundary-adjacent points (Section III-B), including a simple method based on prediction confidence and two advanced selection techniques by clustering the unlabeled or generated data in the latent embedding space.
+Additionally, we introduce three guided DDPM fine-tuning approaches (Section III-C), aiming to generate the subset of desirable data directly, thus avoiding the computation overhead of pre-generating the whole extra dataset.
+Comprehensive experiments on image benchmarks demonstrate that our data selection schemes, especially for the best-performing LCS-KM approach, significantly improve the data and computational efficiency of baseline SSAT methods, while achieving comparable robust accuracies (Section IV-A).
+For instance, utilizing $10\%$ extra unlabeled data identified by LCS-KM in SSAT can achieve similar robustness performance on SVHN compared to using the entire dataset, while reducing the total runtime to approximately $25\%$.
+By directly generating the exact amount of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can reduce the computation overhead while maintaining robustness (Section IV-C).
+Moreover, we demonstrate the generalizability of our data reduction schemes to a real-world COVID-related medical application (Section V-A) and visualize the distributions of selected data points to further illustrate why our LCS-KM method always performs the best (Section V-B).
+Finally, we provide general guidelines on how to choose the key hyperparameters for maximizing the SSAT performance (Section V-C), and conclude with more discussions on related literature, limitations of our study, and future research directions (Section VI).
+Our work reveals the importance of prioritizing boundary-adjacent points for improving SSAT’s efficiency and highlights the superiority of latent-based k-means clustering, potentially assisting the development of more efficient and scalable robust learning algorithms for real-world applications.
+II Inefficiency of Existing SSAT Methods
+II-A Semi-Supervised Adversarial Training
+We work with adversarially robust classification tasks and focus on semi-supervised adversarial training (SSAT), which was initially introduced to the field by Carmon et al.
+[11].
+Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be the input space and $\mathcal{Y}$ be the output space of class labels.
+Assume $D_{l}$ is the underlying labeled data distribution over $\mathcal{X}\times\mathcal{Y}$ that we aim to learn adversarially robust classifiers.
+Let $\mathcal{S}_{l}$ be the training dataset with examples independently and identically drawn from $D_{l}$, and $\mathcal{S}_{u}$ be a collection of inputs sampled from an unlabeled data distribution $D_{u}$ defined over $\mathcal{X}$.
+In adversarial ML literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and the set size of $\mathcal{S}_{u}$, denoted as $|\mathcal{S}_{u}|$, to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are much easier to acquire than well-annotated labeled data.
+For instance, when $\mathcal{S}_{l}$ corresponds to the 5050K labeled CIFAR-10 training images, $\mathcal{S}_{u}$ is the 500500K “most CIFAR-10-like but non-identical” images selected from the whole 8080M Tiny ImageNet dataset [11].
+Specifically, SSAT usually first trains an intermediate model $f_{\hat{\theta}}$ on the labeled training set $\mathcal{S}_{l}$, known as the pseudo labeler, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$.
+The weight parameters of the intermediate model are optimized by minimizing the following objective function:
+$\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell(\theta,\bm{x},y),$ (1)
+where $\ell(\cdot)$ stands for the individual standard loss (e.g., cross-entropy), capturing the discrepancy between model predictions and the ground-truth labels.
+Finally, SSAT adversarially trains a classification model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled data $\mathcal{S}_{l}$ and unlabeled data $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
+$\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},$ $\gamma)=\operatorname*{argmin}_{\theta}\quad\frac{\gamma}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell_{\mathrm{adv}}(\theta,\bm{x},y)$
+$+\frac{1-\gamma}{|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\ell_{\mathrm{adv}}\!\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$ (2)
+where $\gamma\geq 0$ is a hyperparameter that controls the trade-off between labeled and unlabeled data.
+A typical approach used in prior work is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\gamma$) but fix the batch size to optimize for the best SSAT performance [13].
+For simplicity, we term $\gamma$ as the per-batch ratio.
+When $\gamma$ is set as $1$, Equation II-A reduces to the optimization objective of vanilla adversarial training (AT) [7].
+In Equation II-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss.
+Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
+$\ell_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\ell(\theta,\bm{x}+\bm{\delta},y),$ (3)
+where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the perturbation ball centered at $\bm{0}$ with radius $\epsilon$ in some distance metric such as $\ell_{p}$-norm.
+Aligned with prior literature [7], we adopt multi-step PGD to approximately solve for the optimization problem in Equation 3.
+II-B Limitations of Prior Work
+Data Inefficiency.
+Although incorporating unlabeled data alleviates the sample complexity barriers of adversarially robust learning [10] and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement.
+Such a trend is clearly documented in the leaderboard of RobustBench [17].
+For CIFAR-10 image classification tasks, state-of-the-art SSAT algorithms either select 500500K external Tiny Imagenet samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original 5050K CIFAR-10 training images.
+Specifically, the method proposed by Gowal et al.
+[13] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [7] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data.
+Training SSAT algorithms on the huge unlabeled dataset demands significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and carbon footprint.
+In addition, handling a larger dataset in ML often means increased GPU processing load and memory usage.
+High Computation Cost.
+Apart from the data inefficiency issue, SSAT also requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of total training epochs compared with vanilla AT (see Figure 1 for supporting evidence).
+Intuitively, learning meaningful and robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more difficult.
+As our experiments illustrate, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples.
+Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster.
+We note that vanilla AT has already been criticized for its high computation [18, 19], due to the iterative PGD steps involved in dealing with the inner maximization problem (Equation 3).
+SSAT methods incur even higher computations, since they converge more slowly and usually adopt larger model architectures (see Table I for concrete runtime comparisons).
+Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
+II-C Problem Formulation
+Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for achieving high model robustness.
+To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
+Strategic Selection.
+Inspired by the idea of coreset selection for data-efficient deep learning [20, 21, 22], the first approach aims to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that SSAT with the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using the full unlabeled dataset $\mathcal{S}_{u}$.
+Formally, we aim to solve the following constrained optimization problem:
+$\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big),\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (4)
+where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation II-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint.
+Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
+$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$ (5)
+where $\ell_{0/1}$ denotes the 0-1 loss function.
 Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly.
-As we will illustrate in the following sections, we design different data selection schemes with respect to the extra unlabeled dataset for SSAT that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
-4. Proposed Data Selection Schemes
-So far, we have introduced the problem task of reducing the unlabeled dataset size to improve the efficiency of SSAT algorithms while maintaining the robustness enhancement and illustrated its importance.
-Motivated by the significant role of boundary-adjacent data points in optimizing model performance (Section 4.1), in this section, we design three efficient data selection schemes to address the optimization problem in Equation 4 (Sections 4.2 and 4.3).
-4.1. Prioritize Boundary Unlabeled Data
-Achieving high model robustness with restricted data resources remains a challenging task in machine learning.
-It involves striking a balance among data significance, optimizing the use of labeled and unlabeled data, and tackling constraints imposed by model capacity.
-Inspired by the prior literature (Zhang et al., 2020; Hua et al., 2021) that emphasizes the imbalanced data importance for vanilla adversarial training (see Section 2.1 for detailed discussions of these works), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT.
-In particular, we propose identifying a small set of vulnerable yet valuable unlabeled data points in $\mathcal{S}_{u}$, which are close to the model’s decision boundary.
+As we will illustrate in Section III-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
+Guided Diffusion.
+The second method proposes to leverage existing pretrained diffusion models but fine-tune them to directly generate a small set of extra data with desirable properties, ready to be incorporated into the SSAT pipeline.
+Similar to our first strategic selection approach, we aim to achieve comparable robustness to existing SSAT while maximizing the efficiency for the utilization of the additional data:
+$\max_{\mathcal{G}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)\big),\text{ s.t. }|\mathcal{G}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (6)
+where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model.
+As will be detailed in Section III-C, we adopt a pretrained DDPM [23] as the backbone while strategically fine-tuning it with novel training losses for guided data generations.
+Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
+III Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
+So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement.
+In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section III-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections III-B and III-C).
+III-A Prioritize Boundary-Adjacent Unlabeled Data
+Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity.
+Inspired by prior literature [16, 24] that emphasizes the unbalanced data importance for vanilla adversarial training (see Appendix A for more discussions), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT.
+In particular, we propose to identify or generate a small set of vulnerable yet valuable unlabeled data points, which are close to the model’s decision boundary.
 Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify.
 Thus, improving their classification can yield more robustness enhancement to adversarial inputs.
-We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy for locating difficult-to-classify data points.
+We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy function for locating difficult-to-classify data points.
 Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points.
-SSAT algorithms can then leverage this information to ensure these boundary-adjacent points are correctly classified, even under adversarial conditions.
-By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies (see Figure 6 in Appendix A for an illustration of the overall training pipeline).
-While a straightforward extension of existing approaches is to select unlabeled data based on how much their predictions change under perturbations found by PGD attacks, this method is computationally expensive, undermining the efficiency gains we seek.
-Specifically, these methods entail iterative optimization processes to pinpoint boundary points, making them less efficient, especially for large-scale datasets.
-Additionally, while these methods provide valuable insights, their complexity often limits their interpretability.
-Therefore, we seek alternatives to effectively identify boundary data points without incurring high computational overhead.
+By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.
+While a plausible approach is to select a set of unlabeled data based on how much their label predictions change under adversarial perturbations returned by PGD attacks, it is computationally expensive, undermining the efficiency gains we aim for.
+In particular, such a PGD-based method entails iterative optimization processes to pinpoint boundary points, making them impractical for large-scale unlabeled datasets.
+Therefore, we seek better alternatives to effectively identify boundary data points without incurring high computational overhead.
 Algorithm 1 Prediction Confidence-Based Selection
-1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
+1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
 2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
@@
 8:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
-9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
+9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
 10:  Output: selected subset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
-4.2. Prediction Confidence-Based Selection
-To identify the critical set of vulnerable data points near the model’s decision boundary while accounting for computational efficiency, scalability, and interpretability, we first propose a straightforward approach, termed as Prediction Confidence-based Selection (PCS), which utilizes the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point.
+III-B Strategic Selection via Latent Clustering
+PCS.
+To identify a small, critical subset of boundary-adjacent data points while accounting for computational efficiency and scalability, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point.
 The pseudocode for such a selection scheme is depicted in Algorithm 1.
 Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized.
-The underlying assumption is that data points with low confidence scores are more likely to lie near the decision boundary, making them ideal candidates for our selection.
+The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection.
 Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes.
-The reason for involving such a trade-off parameter will be further discussed in Section 6.2.
-The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data using model confidence scores.
-Note that for self-supervised adversarial training, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset.
-Thus, an efficient data selection scheme is desirable to avoid high computational overhead.
-Nevertheless, we discovered in our experiments that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT.
-We hypothesize that PCS overlooks the geometric relationships and distributional properties, which are crucial for characterizing boundary-adjacent points, particularly for complex datasets.
-In addition, DNNs have been shown to be overconfident in their predictions (Guo et al., 2017), suggesting prediction confidence score might be an inherently biased indicator.
-4.3. Latent Clustering-Based Selection
-To overcome the above issues, we propose latent clustering-based selection (LCS) strategies, which identify data points near the model’s decision boundary in the latent space using different clustering techniques.
-Our approach begins by generating latent embeddings for all unlabeled data $\bm{z}=h_{\hat{\theta}}(\bm{x})$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$.
+The reason for involving such a trade-off parameter will be further discussed in Section V-C.
+The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data.
+Note that for SSAT, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset.
+Thus, an efficient selection scheme is desirable to avoid high computational overhead.
+Nevertheless, our experiments showed that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT.
+We hypothesize that PCS overlooks the geometric relationships and distributional properties, which can be crucial for characterizing boundary-adjacent points.
+In addition, DNNs have been shown to be overconfident in their predictions [25], suggesting that prediction confidence scores returned by neural network models might be inherently biased.
+LCS.
+To overcome the aforementioned issues, we introduce latent clustering-based selection (LCS) schemes to pinpoint desirable data in the model’s latent space using different clustering techniques.
+Our approach begins by generating latent embeddings for all $\bm{z}=h_{\hat{\theta}}(\bm{x})$ with $\bm{x}\in\mathcal{S}_{u}$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$.
 The penultimate layer captures more abstract, high-level features and better represents the underlying data structure, which helps identify points near decision boundaries.
@@
 Algorithm 2 Latent Clustering-Based Selection
-1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
+1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
 2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
@@
 4:  Initialize $\mathcal{A}_{u}$ as empty set
-5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in D_{u}$
+5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
 6:  if using LCS-KM then
 7:   $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}\leftarrow$ k-means clustering on $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
-8:     $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in D_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
+8:   $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in\mathcal{S}_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
 9:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta d$
@@
 11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
-12:     Compute posterior probabilities $\bm{p}(\bm{z})$ for each $\bm{z}$
+12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
 13:   $\Delta p\leftarrow$ compute $\Delta p=|p_{1}-p_{2}|$ for each $\bm{z}$, where $p_{1}$ and $p_{2}$ are the top two highest probabilities
@@
 16:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
-17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
+17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
 18:  Output: selected dataset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
 LCS-KM.
-In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the $N$ unlabeled data points into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster.
-For each latent embedding $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$.
+In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the unlabeled dataset into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster.
+For each $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$.
 Data points are selected based on the minimal difference in distance between each latent embedding $\bm{z}$ to the corresponding two closest cluster centroids $\Delta d=|d_{1}-d_{2}|$, where $d_{1}$ and $d_{2}$ are the Euclidean distances to the closest and second closest centroids, respectively.
-The set of unlabeled inputs with the smallest $\Delta d$ values are selected.
-Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset followed by the same balancing step using the ratio parameter $\beta$.
-As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during self-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
+Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset, followed by the same balancing step using the ratio parameter $\beta$.
+As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during semi-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
 LCS-GMM.
 In this variant of LCS, we again start by computing latent embeddings for unlabeled data $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$.
-Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models.
+Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models (GMMs).
 A GMM assumes that the data is generated from a mixture of $k$ Gaussian distributions, each with its own mean $\bm{\mu}_{j}$ and covariance matrix $\mathbf{\Sigma}_{j}$ for the $j$-th component.
 Mathematically, each data point $\bm{z}$ has a probability of belonging to the $j$-th Gaussian component given by the posterior probability:
-$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$
-where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ represents the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$.
-We denote $\bm{p}(\bm{z})=[p_{1}(\bm{z}),p_{2}(\bm{z}),\ldots,p_{k}(\bm{z})]$ as the probability vector.
-To identify data points near the decision boundary, we focus on those points for which the posterior probabilities $\bm{p}(\bm{z})$ are similar across multiple Gaussian components, indicating that they are near the boundary between different clusters.
-Specifically, for each data point $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively.
+$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$ (7)
+where $\pi_{j}$ is the mixing coefficient for the $j$-th Gaussian component, and $\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ denotes the Gaussian distribution with mean $\bm{\mu}_{j}$ and covariance $\Sigma_{j}$.
+We focus on data points with similar top posterior probabilities, suggesting that they are near the boundary between Gaussian distributions.
+Specifically, for each $\bm{z}$, we calculate the difference between the highest and the second-highest posterior probabilities: $\Delta p=|p_{1}(\bm{z})-p_{2}(\bm{z})|$, where $p_{1}(\bm{z})$ and $p_{2}(\bm{z})$ are the highest and second-highest posterior probabilities, respectively.
 Data points with the smallest $\Delta\gamma$ values are selected, as they are more likely to be near decision boundaries.
 Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset.
-As will be illustrated in Section 5, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$).
-More detailed discussions of the difference between these data selection schemes are provided in Section 6.1, where we visualize the selected unlabeled data in a two-dimensional latent space.
-5. Experiments
-In this section, we comprehensively evaluate the performance of our proposed data selection schemes, as described in Algorithms 1 and 2, using two widely recognized image benchmarks: SVHN (Netzer et al., 2011) and CIFAR-10 (Alex, 2009) (Section 5.2).
-Following existing literature, we consider $\ell_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10.
-We also highlight the computational advantages of our methods, particularly their ability to reduce the convergence time of SSAT algorithms (Section 5.3).
-Furthermore, we assess the generalizability of our techniques by applying them to a real-world medical dataset, demonstrating the applicability of our methods beyond standard image benchmarks (Section 5.4).
-Table 1.
-Comparison results of self-supervised adversarial training performance under different unlabeled data selection schemes and varying ratios on two image benchmarks.
-For SVHN, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ and train models on Wide ResNet 16-8.
-For CIFAR-10, we consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$ and train models on Wide ResNet 28-10.
+As will be illustrated in Section IV, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$).
+More detailed discussions of the difference between these data selection schemes are provided in Section V-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
+Algorithm 3 SSAT with Guided DDPM Fine-Tuning
+1:  Input: labeled $\mathcal{S}_{l}$; pre-trained DDPM $\theta_{\mathrm{pre}}$; fine-tuning epochs $S$; trade-off parameter $\lambda$; other parameters $\epsilon$, $\alpha|\mathcal{S}_{u}|$, $\beta$, $\gamma$, $k$, $\eta$
+2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
+3:  Initialize $\mathcal{G}_{u}$ as empty set
+4:  $\mathcal{G}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points generated by pre-trained $\theta_{\mathrm{pre}}$
+5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
+6:  for $s=1,2,\ldots,S$ do
+7:   Sample batches of size $b$: $\bm{x}_{0}\sim\mathcal{S}_{l}$, $t\sim\mathcal{U}(T)$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I})$
+8:   $\bm{x}_{t}\leftarrow\sqrt{\bar{\alpha}_{t_{i}}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$
+9:   $L_{\mathrm{DDPM}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|_{2}^{2}$
+10:   if using PCG then
+11:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{PC}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+12:   else if using LCG-KM then
+13:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{KM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+14:   else if using LCG-GMM then
+15:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{GMM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+16:   end if
+17:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\big(L_{\mathrm{DDPM}}+\lambda\cdot L_{\mathrm{reg}}\big)$
+18:  end for
+19:  $\mathcal{G}_{u}\leftarrow$ add $\beta\cdot\alpha|\mathcal{S}_{u}|$ points generated by fine-tuned $\theta$
+20:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)$
+21:  Output: generated dataset $\mathcal{G}_{u}$, final model $\theta_{\mathrm{final}}$
+III-C Guided Diffusion via Latent Clustering
+While our selection-based strategy can significantly reduce the unlabeled data set size and lower the overall training time, it has a critical drawback–when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement.
+As shown in Table II, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU.
+Note that the best current SSAT method documented in the RobustBench leaderboard [17] utilizes a few hundred million synthetically generated images, suggesting the total generation time can be a significant computational overhead.
+Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved.
+In this section, we propose an alternative generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively.
+In particular, we fine-tune a pretrained DDPM using novel guidance losses for more efficient unlabeled data generation based on the insights from previous sections.
+In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the computation overhead.
+DDPM.
+Diffusion model resembles a milestone in the history of generative modeling, which is the current state-of-the-art for high-quality image generation and synthesis.
+A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [23].
+By gradually removing Gaussian noise through the reverse process, DDPM can thus generate real-looking images from pure noise.
+Let $q_{0}$ be a real image distribution over $\mathbb{R}^{d}$ and $\mathcal{N}(\bm{0},\mathbf{I})$ be the noise distribution.
+The forward process defines how to map images $\bm{x}_{0}\sim q_{0}$ to Gaussian noises $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ through conditional sampling: for any $t\in[T]$, let
+$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I}),$ (8)
+where $\beta_{1},\beta_{2},\ldots,\beta_{T}$ are scheduling parameters.
+According to the properties of Gaussian distributions, Equation 8 implies that we can directly sample $\bm{x}_{t}$ for any $t\in[T]$ based on:
+$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (9)
+where $\bar{\alpha}_{t}=\Pi_{s=1}^{t}(1-\beta_{s})$. Based on Equation 9, Ho et al. [23] proposed to train DDPM by minimizing a regression loss:
+$L_{\mathrm{DDPM}}(\theta)=\mathbb{E}_{\bm{x}_{0}\sim q_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I}),t\in\mathcal{U}(T)}\big\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big\|_{2}^{2},$ (10)
+where $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ denotes the noisy image corresponding to $\bm{x}_{0}$ at $t$-th forward step, $\bm{\epsilon}_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ stands for a neural net with weights $\theta$ for noise prediction from $\bm{x}_{t}$, and $\mathcal{U}(T)$ denotes the uniform distribution over the set $\{1,2,\ldots,T\}$.
+Once the model $\bm{\epsilon}_{\theta}$ is trained, we can then generate synthetic images from Gaussian noise via the following reverse (or denoising) process: for $t=T,\ldots,1$,
+$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\big(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\big)+\sqrt{\beta_{t}}\bm{\delta},$ (11)
+where $\alpha_{t}=1-\beta_{t}$, and $\bm{\delta}\sim\mathcal{N}(\bm{0},\mathbf{I})$.
+For simplicity, denote by $g_{\theta}(\bm{x}_{t},t)$ the right-hand side of Equation 11, and $g_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ can be viewed as one-step DDPM denoising.
+Define DDPM Guidance Loss.
+Now that we have explained DDPM’s forward and backward processes, we can introduce the detailed design of our fine-tuning schemes.
+Borrowing the insights from previous sections, our method aims to fine-tune a DDPM model to directly generate the required subset of boundary-adjacent unlabeled data points for more efficient SSAT.
+Similar to our selection-based methods, we consider three designs: prediction confidence-guided generation (PCG), latent clustering-guided generation with k-means (LCG-KM), and latent clustering-guided generation with Gaussian mixture models (LCG-GMM).
+Specifically, we define a guidance loss $\ell_{\mathrm{guide}}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ corresponding to each design, which will serve as a fine-tuning regularization term to bias the diffusion model to generate more boundary-adjacent samples.
+PCG. The first design utilizes the prediction confidence of the intermediate model $f_{\hat{\theta}}$ and defines the guidance loss as
+[…]
+A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary.
+As we’ve explained in Section III-B, the prediction confidence of a neural network classifier can be an unreliable estimator.
+Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches.
+Based on different clustering techniques, we again introduce two possible variants.
+LCG-KM.
+For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters.
+Note that this step is different from LCS-KM, because we now do not pre-generate the unlabeled dataset.
+Let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters.
+We define the guidance loss as:
+[…]
+where $d_{1}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k]}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ denotes the Euclidean distance between the latent embeddings of $\bm{x}$ and the closest centroid, $d_{2}(\bm{x};h_{\hat{\theta}})=\min_{j\in[k],j\neq j_{1}}\|h_{\hat{\theta}}(\bm{x})-\bm{\mu}_{j}\|_{2}$ is the second smallest distance, and $j_{1}$ stands for the index of the closest centroid.
+By definition, we know $\ell_{\mathrm{KM}}(\bm{x})$ is always non-negative, and a smaller value of $\ell_{\mathrm{KM}}(\bm{x})$ indicates that the input $\bm{x}$ is more likely to be a boundary-adjacent point.
+LCG-GMM.
+Similarly, we can define an informative guidance loss based on the coefficients of the Gaussian mixture models.
+Let $\{\bm{\mu}_{j},\mathbf{\Sigma}_{j})$ be the mean and covariance parameters returned by GMM with respect to the latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$.
+Let $p_{j}(\bm{z})$ be the $j$-th GMM posterior probability defined by Equation 7.
+Then, the guidance loss is defined as:
+[…]
+[…]
+Guided DDPM Fine-Tuning.
+To utilize our designed guidance losses in DDPM fine-tuning, we propose to first denoise $\bm{x}_{t}$ via a single DDPM reverse step $g_{\theta}(\bm{x}_{t},t)$, where $\bm{x}_{t}$ is a noisy image at $t$-th time step defined by the forward process, and then apply the respective guidance loss to $g_{\theta}(\bm{x}_{t},t)$.
+For any sample $\bm{x}_{0}\sim q_{0}$, time $t\sim\mathcal{U}(T)$ and noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we define the following loss function $L_{\mathrm{reg}}(\theta)$ that will be used as an additional regularization term to fine-tune DDPM:
+$L_{\mathrm{reg}}(\theta)=\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\big[\ell_{\mathrm{guide}}\big(g_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\big)\big],$ (15)
+where $\ell_{\mathrm{guide}}$ corresponds to Equation 12 when we use prediction confidence, and refers to Equations 13 and 14 when using LCG-KM and LCG-GMM, respectively.
+Putting all the pieces together, the fine-tuning loss for direct generation of boundary-adjacent data is defined as:
+$L_{\mathrm{tot}}(\theta)=L_{\mathrm{DDPM}}(\theta)+\lambda\cdot L_{\mathrm{reg}}(\theta),$ (16)
+where $\lambda\geq 0$ balances the trade-off between high-quality image generation captured by the original DDPM loss and boundary-seeking data generation driven by our guidance loss.
+By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD).
+Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to bias the diffusion model to generate the desirable subset of synthetic images.
+IV Experiments
+In this section, we test the performance of our data reduction strategies proposed in Section III compared with baseline SSAT algorithms.
+In particular, we first conduct experiments to evaluate our selection-based methods in terms of robustness and efficiency advantages when integrated into the SSAT pipeline (Section IV-A).
+Then, we highlight the computational advantages of our approaches, particularly their capabilities of accelerating the convergence and reducing the total runtime of SSAT (Section IV-B).
+Finally, we study the effectiveness of our DDPM fine-tuning schemes in further improving the efficiency of SSAT algorithms under a generative setup (Section IV-C).
+To begin with, we lay out the main experimental setup.
+All our experiments are conducted using a single A100 GPU.
+All the remaining details, including all the hyperparameters, for reproducing our experiments are provided in Appendix B.
+Dataset.
+We test two image benchmarks in our evaluation, SVHN [26] and CIFAR-10 [27], both widely used in prior adversarial ML literature.
+We consider two scenarios in terms of the source of the utilized unlabeled dataset $\mathcal{S}_{u}$ for testing our selection-based strategies: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model.
+For external scenarios, we follow the implementation protocols outlined by Carmon et al.
+[11].
+To be more specific, SVHN models are trained on 7373K labeled digit images from the original dataset, supplemented by 531531K extra unlabeled SVHN images.
+For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset.
+For generative setup, we adopt 11M synthetic images generated by DDPM for both datasets, following the protocols documented by Gowal et al.
+[13].
+Configuration.
+Following prior literature, we mainly consider $L_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10.
+We adopt WideResNet (WRN) architectures and train models using TRADES [9] on both labeled and unlabeled data, coupled with pseudo labels returned by the intermediate model.
+In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations.
+To standardize training, an epoch is defined as processing 5050K data points, regardless of the total dataset size.
+Prior works also adopted this approach [11, 13], which ensures that training time per epoch remains consistent across datasets of varying sizes.
+SSAT.
+For experiments with external unlabeled data, SSAT algorithms are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs.
+To compare the performance of our methods in a fair way, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the best model.
+This approach aligns with the early stopping practices commonly used in adversarial ML literature [9, 28].
+In experiments with external data and ratio parameter $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN.
+Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs.
+Total training time is reported as the duration required to achieve the best model performance.
+Intermediate Model.
+Our data selection and guided diffusion schemes require training an intermediate model $f_{\hat{\theta}}$, which we adopt the same architecture as the final model.
+This intermediate model is trained using standard supervised learning for $100$ epochs on a WRN-28-10 architecture, which takes around $55$ minutes.
+Notably, this time is excluded from the reported total runtime.
+The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our proposed strategies.
+As such, this training step is integral to the overall workflow, irrespective of whether a specific selection or fine-tuning method is employed.
+Evaluation Metric.
+To test the robustness of the final model returned by SSAT, we mainly use multi-step PGD attacks (PGD) [7], due to their broad adoption in adversarial ML.
+We also report robust accuracies against AutoAttack (AA) [29] for more rigorous robustness evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
+TABLE I: SSAT performance under various unlabeled data selection schemes.
+We consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ for SVHN, and $\ell_{\infty}$ perturbations with $\epsilon=0.031$ for CIFAR-10.
+For each configuration, we underline the best robust accuracies.
 Dataset Ratio $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
-[…]
+Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h) Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h)
 SVHN $0\%$ No Sel. $94.4$ $75.6$ $68.3$ $2.62$ $94.4$ $75.6$ $68.3$ $2.62$
@@
 LCS-GMM $95.1$ $81.8$ $72.1$ $3.97$ $95.6$ $81.2$ $71.8$ $4.27$
-LCS-KM $95.2$ $82.7$ $72.9$ $3.90$ $96.4$ $82.4$ $72.6$ $4.18$
+LCS-KM $95.2$ $\underline{82.7}$ $\underline{72.9}$ $3.90$ $96.4$ $\underline{82.4}$ $\underline{72.6}$ $4.18$
 $10\%$ Random $95.3$ $82.0$ $73.0$ $3.28$ $95.6$ $83.6$ $72.4$ $3.28$
@@
 LCS-GMM $96.2$ $83.0$ $74.3$ $3.97$ $95.5$ $84.1$ $74.1$ $4.27$
-LCS-KM $96.1$ $86.3$ $75.2$ $3.90$ $96.6$ $86.6$ $74.8$ $4.18$
+LCS-KM $96.1$ $\underline{86.3}$ $\underline{75.2}$ $3.90$ $96.6$ $\underline{86.6}$ $\underline{74.8}$ $4.18$
 $20\%$ Random $96.2$ $82.3$ $73.2$ $3.28$ $96.3$ $84.6$ $74.0$ $3.28$
@@
 LCS-GMM $96.4$ $84.3$ $75.0$ $3.97$ $95.8$ $85.5$ $74.7$ $4.27$
-LCS-KM $96.2$ $86.6$ $75.1$ $3.90$ $96.7$ $87.2$ $75.3$ $4.18$
+LCS-KM $96.2$ $\underline{86.6}$ $\underline{75.1}$ $3.90$ $96.7$ $\underline{87.2}$ $\underline{75.3}$ $4.18$
 CIFAR-10 $0\%$ No Sel. $84.9$ $55.4$ $49.2$ $10.69$ $84.9$ $55.4$ $49.2$ $10.69$
@@
 LCS-GMM $84.2$ $55.1$ $52.1$ $14.89$ $85.8$ $54.6$ $51.8$ $15.36$
-LCS-KM $85.6$ $56.4$ $52.9$ $14.78$ $85.9$ $56.1$ $52.5$ $15.27$
+LCS-KM $85.6$ $\underline{56.4}$ $\underline{52.9}$ $14.78$ $85.9$ $\underline{56.1}$ $\underline{52.5}$ $15.27$
 $10\%$ Random $85.2$ $56.0$ $52.4$ $14.25$ $85.0$ $56.5$ $52.6$ $14.25$
@@
 LCS-GMM $85.9$ $57.1$ $54.2$ $14.89$ $85.8$ $57.1$ $54.5$ $15.36$
-LCS-KM $87.2$ $58.2$ $55.3$ $14.78$ $86.1$ $58.0$ $55.8$ $15.27$
+LCS-KM $87.2$ $\underline{58.2}$ $\underline{55.3}$ $14.78$ $86.1$ $\underline{58.0}$ $\underline{55.8}$ $15.27$
 $20\%$ Random $87.1$ $57.5$ $54.2$ $14.25$ $85.4$ $57.2$ $54.2$ $14.25$
@@
 LCS-GMM $87.0$ $58.2$ $55.5$ $14.89$ $85.9$ $58.6$ $55.8$ $15.36$
-LCS-KM $88.7$ $60.7$ $57.8$ $14.78$ $85.5$ $60.2$ $57.2$ $15.27$
-5.1. Experimental Settings
-First, we introduce the necessary details to understand our experimental results. All the remaining details for reproducing our experiments are provided in Appendix B.
-Dataset.
-For the image benchmark experiments, the proposed selection schemes are initially applied to external unlabeled datasets, following the implementation protocols outlined by Carmon et al.
-(2019).
-For the SVHN experiments, models are trained on $73$K labeled digit images from the original SVHN dataset, supplemented by $531$K additional unlabeled SVHN images.
-Specifically, the CIFAR-10 dataset comprises $50$K labeled training images and $10$K labeled test images, with models trained using SSAT augmented by an external unlabeled dataset $\mathcal{S}_{u}$ containing $500$K images sampled from the $80$M Tiny Images (80M-TI) dataset.
-To assess the generalizability of our selection schemes, we further evaluate SSAT algorithms using $1$M synthetically generated images created by the Denoising Diffusion Probabilistic Model (DDPM) for both CIFAR-10 and SVHN, following the protocols described by Gowal et al.
-(2021).
-Configuration.
-We primarily conduct experiments on SSAT using WideResNet architectures.
-We train models using TRADES (Zhang et al., 2019) on both labeled and unlabeled data coupled with pseudo labels generated by the intermediate model.
-In Section 6.3, we evaluate the sensitivity of our methods for varying $\ell_{\infty}$ perturbation size, PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations.
-To standardize training, an epoch is defined as processing $50$K data points, regardless of the total dataset size.
-This is achieved by calculating the required number of batches to cover $50$K data points based on the batch size.
-This approach, adopted in prior works (Carmon et al., 2019; Gowal et al., 2021), ensures that training time per epoch remains consistent across datasets of varying sizes.
-SSAT.
-For experiments with external unlabeled data, SSAT models are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs.
-To evaluate the performance of each data selection scheme, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the “best” model.
-This approach aligns with the early stopping practices commonly used in adversarial ML literature (Zhang et al., 2019; Rice et al., 2020).
-In experiments with external data and selection ratios $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN.
-Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs.
-Total training time is reported as the duration required to achieve the best model performance.
-Intermediate Model.
-As a foundational step, the proposed selection schemes require training an intermediate model that utilizes the same architecture as the final model.
-This intermediate model is trained using standard supervised learning for $100$ epochs, with training on the WRN-28-10 architecture taking approximately $54$ minutes and $43$ seconds.
-Notably, this time is excluded from the reported training durations presented in the results table.
-The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for the implementation of any data selection strategy.
-As such, this training step is integral to the overall workflow, irrespective of whether a specific selection method is employed.
-Evaluation Metric.
-To evaluate model robustness, we primarily focus on using multi-step PGD attacks (PGD) (Madry et al., 2017) due to their wide adoption in existing adversarial ML literature.
-We also report the robust accuracy using AutoAttack (AA) (Croce and Hein, 2020b) in our main experiments on image benchmarks for more rigorous evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
-5.2. Main Evaluation on Image Benchmarks
+LCS-KM $88.7$ $\underline{60.7}$ $\underline{57.8}$ $14.78$ $85.5$ $\underline{60.2}$ $\underline{57.2}$ $15.27$
+IV-A Strategic Unlabeled Data Selection
 SVHN.
-Our SVHN evaluation results, summarized in Table 1, demonstrate the efficacy of SSAT algorithms with unlabeled data using various selection schemes and $\alpha\in\{0\%,1\%,10\%,20\%,100\%\}$ trained on a WRN-16-8 architecture.
-Specifically, employing the LCS-KM selection strategy with $10\%$ or $20\%$ of the additional data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which closely approximate the $86.0\%$ robustness achieved using the entire unlabeled dataset.
+Our evaluation results on SVHN are summarized in Table I, which demonstrate the efficacy of the proposed data selection schemes when employed in SSAT algorithms trained using a WRN-16-8 architecture with varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, with comparisons to baseline SSAT performance with $\alpha\in\{0\%,100\%\}$.
+Specifically, employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset.
 In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM.
-The robust accuracies evaluated using AutoAttack (i.e., the “AA” column of Table 1) demonstrate similar patterns compared to that using multi-step PGD Attacks (i.e., the “PGD” column of Table 1).
-Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, reinforcing the robustness and adaptability of the proposed data selection framework.
+The robust accuracies evaluated using AutoAttack (i.e., the AA column of Table I) demonstrate similar patterns compared to those using multi-step PGD Attacks (i.e., the PGD column of Table I).
+Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights.
 Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases.
-We provide detailed analyses and discussions of these ratio hyperparameter configurations in Section 6.2.
-We also note that the use of ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance improvements stem from the proposed data selection strategy rather than label accuracy (see Table 4 in Appendix D.2 for details).
-Additionally, experiments with $1$M data points generated using DDPM for SSAT (Gowal et al., 2021) highlight that selecting only $10\%$$20\%$ of data through LCS-KM achieves comparable performance to utilizing the full unlabeled dataset, with LCS-KM consistently demonstrating superior results.
-Importantly, training SSAT with selected data requires only $75$ epochs to optimize robust accuracy, whereas training with the full dataset demands up to $400$ epochs.
-This significant reduction in computational cost further underscores the practical advantages of the proposed selection methodology, which will be elaborated in Section 5.3.
+We provide detailed sensitivity analyses of these hyperparameters in Section V-C.
+Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline.
+Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime.
+In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs.
+Such a large reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section IV-B.
+Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table IV in Appendix C).
 CIFAR-10.
-For CIFAR-10 experiments, we adopt the WRN-28-10 architecture to train models using self-supervised adversarial training algorithms.
-To assess the generalizability of our observations, we also conduct ablation studies on a ResNet-18 architecture, which are detailed in Table 5 in Appendix D.4).
-The impact of varying proportions, $\alpha$, of pseudo-labeled unlabeled data was analyzed, revealing that selecting $20\%$ of the external data using LCS-KM achieved a robust accuracy of $60.7\%$, comparable to the $61.2\%$ obtained with the entire dataset.
-In contrast, the random selection of the same proportion yielded only $57.8\%$, underscoring the advantages of strategic data selection techniques such as LCS-KM.
-Building on the SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data.
-Using $1$ million images generated by DDPM (Gowal et al., 2021), we discover that utilizing just $20\%$ of the generated data via LCS-KM resulted in performance metrics nearly identical to those achieved with the full dataset.
-These trends are consistent across experiments with $\ell_{2}$ perturbations (see Figure LABEL:fig:two_graphs_ab for evidence).
+We adopt a WRN-28-10 architecture for CIFAR-10 experiments.
+To assess the generalizability of our findings, we also conduct ablations on ResNet-18, which are detailed in Table V in Appendix C.
+Across varying ratios of utilized unlabeled data $\alpha$ and different data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains robust accuracy of $60.7\%$, comparable to the $62.5\%$ obtained with the entire dataset.
+In comparison, random selection with the same ratio parameter achieves only $57.8\%$, underscoring the advantages of our strategic LCS technique with k-means.
+Similar to SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data.
+Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset.
+These trends are also consistent across experiments with $\ell_{2}$ perturbations (see Figure 10 for supporting evidence).
 Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations.
@@
 By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance.
-These results validate the efficacy of our latent clustering-based selection approach in enhancing self-supervised adversarial training (SSAT).
-The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in adversarial training frameworks.
-Table 3 in Appendix C summarizes the comparisons, showcasing that LCS-KM improves model robustness while significantly improving the efficiency of SSAT algorithms with full unlabeled data.
-5.3. Computational Benefits
-As suggested by the runtime comparisons depicted in Table 1, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall computation costs of SSAT while maintaining competitive, robust accuracy.
-This section further visualizes the learning curves of SSAT and discusses the computational benefits associated with our proposed methods.
-Figure LABEL:fig:convergence depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages.
-In our setup, one training epoch corresponds to processing $50$K data points, with $10$ steps for PGD training and $20$ for PGD testing.
-When no additional data is utilized, as shown in Figure LABEL:fig:0per, the highest robust test accuracy is achieved after approximately $75$ epochs.
-As illustrated in Figures LABEL:fig:10percrandom and LABEL:fig:10perc, incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs.
-In contrast, as depicted in Figure LABEL:fig:100perc, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs.
-These findings suggest that including extra unlabeled data in SSAT increases computational demands, as evidenced by the longer training convergence time.
-Strategic selection reduces redundancy in the dataset, allowing the model to achieve optimal performance with fewer training epochs.
-Specifically, we observe that the size of the selected unlabeled dataset plays a critical role in determining the time required for the model to achieve peak accuracy.
-Larger unlabeled datasets usually extend the convergence times of SSAT, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training.
+These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training.
+The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
+(a) Vanilla AT
+(b) Random Selection ($\alpha=10\%$)
+(c) LCS-KM ($\alpha=10\%$)
+(d) Full Utilization ($\alpha=100\%$)
+Figure 1: Illustration of standard and robust accuracy curves of SSAT on CIFAR-10 with different configurations of unlabeled data selection from the external 500500K Tiny Images: (a) No extra data, (b) random selection with $\alpha=10\%$, (c) LCS-KM with $\alpha=10\%$, and (d) utilizing all 500500K unlabeled data.
+TABLE II: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
+Ratio $\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
+$0\%$ No Gen. $100$ $10.7$ $55.4$
+$100\%$ Std. DDPM $3.90$ $400$ $61.0$ $61.8$
+$20\%$ LCS-KM $3.90$ $0.97$ $100$ $19.1$ $60.3$
+$20\%$ LCG-KM $0.43$ $0.77$ $100$ $15.7$ $60.2$
+TABLE III: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
+Ratio $\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
+$0\%$ No Gen. $84.9$ $55.4$ $49.2$
+$100\%$ Std. DDPM $85.7$ $61.8$ $58.4$
+$1\%$ PCG $84.4$ $55.4$ $52.8$
+LCG-GMM $85.3$ $55.8$ $52.6$
+LCG-KM $85.2$ $56.5$ $53.1$
+$10\%$ PCG $84.7$ $56.7$ $54.2$
+LCG-GMM $85.8$ $57.6$ $55.0$
+LCG-KM $84.2$ $57.9$ $55.6$
+$20\%$ PCG $85.9$ $58.3$ $55.9$
+LCG-GMM $84.7$ $58.3$ $56.1$
+LCG-KM $85.7$ $60.3$ $57.4$
+IV-B Computational Benefits of Data Reduction
+As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall training time of SSAT while maintaining competitive robustness.
+We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
+Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages.
+In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing.
+When no additional data is utilized, the highest robust test accuracy is achieved after approximately $75$ epochs (Figure 1a).
+Incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c).
+In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d).
+These findings suggest that including extra unlabeled data in SSAT increases computational demands.
+Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
+Generally speaking, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy.
+Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training.
 However, continued training beyond this point often leads to robust overfitting.
@@
 These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
-5.4. Application to Real-World Medical Data
-We extend our study to a real-world healthcare application with COVID-19-related datasets.
-Specifically, the COVIDGR dataset (Tabik et al., 2020) is utilized as the labeled dataset, and the CoronaHack dataset (Govi, 2020) serves as the source of unlabeled data.
-This extension allowed us to validate the efficacy of the proposed methods in a domain where data scarcity and model robustness are critical considerations.
-Setting.
-The experimental setup consists of multiple stages, leveraging both datasets in conjunction with ResNet-18 as the base model.
-The COVIDGR dataset contains $852$ chest X-ray images evenly distributed across two classes: $426$ COVID-19 positive cases (confirmed via RT-PCR testing) and $426$ negative cases.
-This dataset was split into an $80\%$ training set and a $20\%$ test set for evaluation purposes.
-The CoronaHack dataset, which comprises $5,910$ chest X-ray images, is divided into an $89\%$ training set and an $11\%$ test set.
-This dataset has an imbalanced class distribution, with $73\%$ of the images classified as COVID-19 positive and $27\%$ as negative (non-COVID).
-For our experiments, we strategically select a subset of the training data from the CoronaHack dataset as unlabeled data to assess the effectiveness of our adversarial robustness techniques.
-We applied the LCS-KM selection strategy to identify and incorporate a minimal yet representative subset of unlabeled data from the CoronaHack dataset into the training process.
-This approach enabled us to maintain robust model performance while significantly reducing the volume of external data required.
-The models are trained on SSAT with varying unlabeled data selection schemes using a ResNet-18 architecture.
-We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and vanilla PGD-based adversarial training.
-Result Analysis.
-The results of our experiments confirm the effectiveness of strategic data selection in enhancing robust accuracy and optimizing training efficiency.
-For instance, using LCS-KM to select $10\%$ of the unlabeled data achieved a robust accuracy of $56\%$, compared to $53\%$ when the same proportion was selected randomly, as shown in Figures LABEL:fig:10perrandom and LABEL:fig:10per.
-Additionally, smaller, strategically selected datasets led to faster convergence, underscoring the benefits of early stopping.
-Specifically, the best accuracy of $56\%$ was reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM in Figure LABEL:fig:10per.
-Figure LABEL:fig:20per shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$.
-In contrast, training on the entire unlabeled dataset (i.e., $\alpha=100\%$) required $60$ epochs to achieve the highest robust accuracy of $58\%$ depicted in Figure LABEL:fig:100per.
-These findings demonstrate that smaller but strategically selected subsets of data can achieve robust accuracy comparable to that of the full dataset while significantly reducing computational time and memory requirements.
-The success of our approach in this real-world application highlights its practical relevance, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
-6. Additional Analyses
-In this section, we conduct additional analyses and experiments to assess the generalizability of our proposed method.
-We start by visualizing and comparing the latent-space behavior of PCS, LCS-KM, and LCS-GMM to evaluate their effectiveness in selecting critical data points (Section 6.1).
-Next, we explore the influence of key ratio hyperparameters in our data selection schemes on the robustness performance achieved by SSAT (Section 6.2).
-Additionally, we evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section 6.3).
-Finally, we examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section 6.4).
-Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table 1.
-6.1. Comparative Visualizations
-To facilitate a comprehensive understanding of the proposed data selection schemes, we perform a comparative visualization using the TinyImages dataset in the latent space.
-The latent representations, extracted from the penultimate layer of a pre-trained WRN-28-10 model on CIFAR-10, encode high-dimensional feature embeddings that capture the data’s underlying structure.
-To render these high-dimensional features interpretable, we employ a two-step dimensionality reduction process that balances the retention of global structure and the visualization of local relationships.
-Setting.
-The first step in this two-step pipeline is Principal Component Analysis (PCA).
-By reducing the dimensionality of the latent features to $50$ components, PCA retains the majority of the variance within the data.
-This step ensures that the global structure of the feature space is preserved, maintaining the relative distances and relationships that reflect the overall organization of the data (Gewers et al., 2021).
-PCA also acts as a preprocessing step, making the subsequent visualization process more efficient and effective.
-In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) to project the $50$-dimensional PCA-reduced features into a two-dimensional space.
-Unlike PCA, which emphasizes preserving variance, t-SNE is designed to maintain local relationships among data points (Xia et al., 2021).
-This combination of PCA and t-SNE allows us to visualize the latent feature space in a manner that is both interpretable and faithful to the original high-dimensional structure.
-After deciding on the two-dimensional latent space, we visualize the selected unlabeled data using our proposed schemes: PCS, LCS-KM, and LCS-GMM.
-To ensure methodological rigor and consistency, we strictly adhere to the procedures outlined in Algorithms 1 and 2 for implementing the data selection techniques.
-The ratio hyperparameters $\alpha$ and $\beta$ are set to $0.1$ and $1$, respectively.
-Importantly, these parameters influence the specific distribution of selected points but do not alter the overall interpretability of the visualizations, ensuring that the comparative analysis remains robust by their precise values.
-PCS.
-Figure LABEL:fig:clustering_techniques presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy.
-Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT across various experimental settings.
+(a) Random Selection ($\alpha=10\%$)
+(b) LCS-KM ($\alpha=10\%$)
+(c) LCS-KM ($\alpha=20\%$)
+(d) Full Utilization ($\alpha=100\%$)
+Figure 2: Standard and robust accuracy curves of SSAT with labeled data from COVIDGRand unlabeled data from CoronaHack, with different selection schemes and ratios: (a) random selection with $\alpha=10\%$, (b) LCS-KM with $\alpha=10\%$, (c) LCS-KM with $\alpha=20\%$, and (d) all unlabeled data $\alpha=100\%$.
+IV-C Guided DDPM Fine-Tuning
+We study the performance of our guided DDPM fine-tuning methods as depicted in Algorithm 3 on CIFAR-10.
+Unlike our previous selection schemes, which required pre-generating the whole dataset, here we generate the same amount of unlabeled data matching the ratio $\alpha$.
+The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection.
+We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments.
+Table III presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart.
+For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM.
+Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives.
+These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set can yield performance comparable to a much larger dataset.
+Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup.
+In particular, we choose to present the best-performing methods with the optimal hyperparameters, namely LCS-KM and LCG-KM under the scenario of $\alpha=20\%$, and compare the end-to-end training times.
+We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively.
+The comparison results are summarized in Table II.
+For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process.
+Our DDPM generator’s throughput is roughly $71$ images per second, so the approximate data generation times are 200200K images in $46$ minutes, 11M images in $3.9$ hours, and 100100M images in $16.2$ days.
+As a result, our guided DDPM finetuning (LCG-KM) approach reduces the total runtime from $19.1$ hours to $15.7$ hours when compared to LCS-KM.
+The improvement will be more pronounced when the total amount of data generated by DDPM is much larger.
+These results highlight the trade-off between robustness gains and computational cost.
+While generating millions of synthetic samples can enhance robustness, it is computationally expensive.
+Our guided fine-tuning approach offers a more resource-efficient alternative, reducing SSAT total runtime from $61.0$ hours to $15.7$ hours while preserving strong robustness.
+(a) PCS
+(b) LCS-GMM
+(c) LCS-KM
+Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space.
+Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the ten class clusters, and (c) LCS-KM highlights points selected near decision boundaries across ten classes based on k-means clustering in the latent space.
+(a) CIFAR-10
+(b) SVHN
+(c) CIFAR-10
+(d) SVHN
+Figure 4: Illustration of SSAT performance using our LCS-KM method on image benchmarks with varying boundary and unlabeled data reduction ratio parameters with respect to: (a) and (c) DDPM-generated CIFAR-10 data with $\alpha=20\%$, (b) and (d) 531531K extra SVHN data with $\alpha=10\%$.
+For each figure, we vary the considered ratio parameter while keeping all the remaining hyperparameters fixed.
+V Additional Analyses
+In this section, we present additional analyses to provide further insights into our methods.
+We first evaluate their performance on a real-world medical application dataset (Section V-A), and then visualize the selected data distribution in the latent space to qualitatively understand their ability in identifying critical data points (Section V-B).
+Finally, we provide ablation studies on key hyperparameters of our methods.
+Due to page limit, additional experiments on other hyperparameters, generalizability under different training schemes and perturbations, and the effect of intermediate model quality in low-label regimes, are provided in Appendix D.
+V-A Application to Real-World Medical Data
+To test the generalization of our findings, we extend our study to a real-world healthcare application.
+COVIDGR [30] is used as the labeled dataset, and CoronaHack [31] serves as the source of the unlabeled data.
+Specifically, COVIDGR contains $852$ chest X-ray images, evenly distributed across two classes: $426$ COVID-19 positive (RT-PCR confirmed) and $426$ negative cases.
+It is then split into $80\%$ training and $20\%$ test sets.
+CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set.
+We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios.
+Further details of the experimental setup is discussed in Appendix B.
+The results are demonstrated in Figure 2, which confirms the effectiveness of our selection scheme in enhancing robustness while optimizing efficiency.
+For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b).
+In comparison, using a smaller unlabeled dataset leads to faster convergence, underscoring the benefits of early stopping.
+Specifically, the best performance of $56\%$ is reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM (Figure 2b).
+Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$.
+In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve the peak robust accuracy of $58\%$ (Figure 2d).
+These findings confirm that using a smaller but strategically selected unlabeled data subset can achieve robustness comparable to that of the whole dataset while significantly reducing computational costs.
+These results highlight the practical relevance of our data reduction scheme, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
+V-B Comparative Visualizations of Our Methods
+To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10.
+We first extract their latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model.
+For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships.
+Specifically, we first perform Principal Component Analysis (PCA) [32].
+By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data.
+In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [33] to project the $50$-dimensional PCA-reduced features onto a two-dimensional space.
+Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points.
+After projecting onto the two-dimensional space, we visualize the set of unlabeled data returned by our selection schemes under $\alpha=10\%$.
+Figure 3 presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy.
+Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT.
 In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes.
-Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers or ambiguous samples that offer limited benefit for refining the decision boundary.
-As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points, as shown in Figure LABEL:fig:PCS_selection.
-This amplifies PCS’s shortcomings compared to the more structured sampling of LCS-KM or the probabilistic clustering offered by LCS-GMM.
-Thus, PCS often dedicates training efforts to less informative or highly noisy examples, resulting in weaker robustness improvements.
-LCS.
-LCS-GMM relies on fitting Gaussian distributions to the latent representations.
-Although this probabilistic approach can effectively capture regions where classes overlap, it inherently presupposes that local data clusters follow (approximately) Gaussian distributions, an assumption that may fail in real-world settings characterized by irregular or noisy data.
-As a result, LCS-GMM’s contours can become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Steinley and Brusco, 2011).
-A closer look at the selected boundary samples in Figure LABEL:fig:gmm also reveals that the region these points occupy is broader and less tightly focused on the decision boundary, making it harder to capture truly critical examples for training.
-Due to the concentration, selected data points overlap with each other in the two-dimensional visualization space.
-In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM, further underscoring the latter’s advantage in both coverage and precision.
-Finally, our best-performing LCS-KM method emphasizes leveraging the latent-space clustering structure to identify critical decision-boundary regions in a more balanced manner.
-By partitioning the data representations into multiple k-means clusters, LCS-KM systematically locates the points of lowest classification confidence within each cluster, illustrated in Figure LABEL:fig:kmeans, thereby ensuring that the algorithm focuses on diverse boundary-adjacent regions.
-This cluster-based approach yields a fine-grained analysis of local uncertainties and avoids the risk of overcommitting to isolated, potentially noisy areas in the latent space.
-6.2. Hyperparameter Sensitivity
-The efficacy of our selection strategies critically relies on several key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$.
+Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers that offer limited benefit for refining the decision boundary.
+As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a).
+In comparison, LCS-GMM and LCS-KM yield more structured sampling patterns.
+While LCS-GMM can roughly capture regions where classes overlap, it assumes that each cluster follows approximately Gaussian distributions, an assumption that may fail in real-world settings.
+As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b).
+A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical examples for training.
+In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM.
+Finally, our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner.
+By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c).
+Compared with LCS-GMM, a notable pattern is that the unlabeled data picked by LCS-KM are aligned more linearly.
+This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer.
+We hypothesize that such a local linear property potentially explains why our latent clustering-based approach with k-means always achieves the best performance compared to the other two alternatives, which is worth further investigation.
+V-C Hyperparameter Sensitivity
+The success of our methods relies on a few key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$.
 We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally.
-Due to page limits, experimental results with varying values of $\gamma$ are deferred to Appendix D.1.
-Ratio of Boundary Points $\bm{\beta}$.
-Concentrating exclusively on points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting.
-To mitigate this, we choose a combination of points, with some near the boundary and others further away.
-The proportion of these selections is determined by the value of $\beta$.
-We conduct experiments with different $\beta$ values to assess their impact on robust accuracy.
-Figure LABEL:fig:betacifar illustrates how robust accuracy varies with $\beta$ on the CIFAR-10 dataset using $20\%$ generated data from DDPM, which is approximately four times the size of the original dataset.
+Due to page limits, the detailed settings of these experiments and the ablations on $\gamma$ are deferred to Appendix C-A.
+Ratio of Boundary Data $\bm{\beta}$.
+Concentrating exclusively on data points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting.
+To mitigate this, we choose a combination of points, with some near the boundary and others further away, where the ratio is determined by $\beta$.
+Figure 4a illustrates how robust accuracy varies with $\beta$ on CIFAR-10 using $20\%$ generated data from DDPM.
 In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away.
-Figure LABEL:fig:betasvhn illustrates how robust accuracy varies with $\beta$ on the SVHN dataset using $10\%$ extra data from the same dataset.
-Here, due to the smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting.
-Consequently, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away.
+Figure 4b illustrates how robust accuracy varies with $\beta$ on SVHN using $10\%$ extra data from the same dataset.
+Due to a smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting.
+As a result, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away.
 Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data.
-When more external data is added, a smaller $\beta$, i.e., fewer points near the boundary, may yield better robust accuracy.
+When more external data is added, using a smaller $\beta$ usually leads to better robust accuracy.
 Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.
-Ratio of Selected Unlabeled Data $\bm{\alpha}$.
-Figures LABEL:fig:cofarper and LABEL:fig:svhnper illustrate how the performance varies with selection ratio $\alpha$ for CIFAR-10 and SVHN, respectively.
-We analyze how the quantity of extra data influences the effectiveness of our data-selection method, which is designed to prioritize data points located near decision boundaries.
-When the external data supply is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy.
-However, as the volume of external or generated data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability.
+Data Reduction Ratio $\bm{\alpha}$.
+Figures 4c and 4d describe how accuracies vary with the ratio parameter $\alpha$.
+We analyze how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary.
+When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy.
+However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability.
 Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
-Table 2.
-Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data.
-We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
-Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
-[…]
-$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
-pre-trained CLIP $86.8$ $83.9$ $40.5$
-$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
-pre-trained CLIP $86.8$ $85.3$ $43.2$
-$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
-pre-trained CLIP $86.8$ $85.6$ $47.7$
-100% fully-supervised $90.5$ $86.2$ $58.2$
-pre-trained CLIP $86.8$ $86.5$ $57.4$
-6.3. Further Exploration of Generalizability
-All experiments in this section are conducted on CIFAR-10.
-We compare four different settings: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data (i.e., $\alpha=0\%$), and using the entire dataset (i.e., $\alpha=100\%$).
-Perturbation Size.
-We consider external data from TinyImages as the unlabeled dataset.
-We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table 1.
-The results are depicted in Figures LABEL:fig:4_255 and LABEL:fig:16_255, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data.
-The trends are shown to be consistent across varying perturbation sizes.
+(a) Fine-Tuning Epochs $S$
+(b) Regularization Strength $\lambda$
+Figure 5: Clean and robust accuracy curves of SSAT with LCG-KM using generated data on CIFAR-10 under $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
+Guided DDPM Finetuning.
+We conduct ablation studies on our guided DDPM fine-tuning approach to better illustrate its behavior.
+In particular, we focus on two key hyperparameters, the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$, and study their impact on accuracy and robustness performance when training SSAT algorithms on CIFAR-10 with LCG-KM and $\alpha=10\%$.
+We also investigate the quality of the data samples generated by our fine-tuned DDPM using both quantitative and qualitative metrics.
+Figure 5a illustrates our ablation results on $S$.
+We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process.
+Figure 6 visualizes the images generated by our DDPM models with different fine-tuning epochs.
+We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained for CIFAR-10.
+Training for too many epochs can lead to a degradation in image quality and loss of diversity, thereby adversely influencing the final robust accuracies.
+In addition, Figure 5b shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$.
+Our findings show that choosing $\lambda$ within the range of $[0.4,0.6]$ results in the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
+Figure 6: Visualization of generated data across different fine-tuning epochs
+VI Conclusion and Future Work
+In this work, we proposed various unlabeled data reduction schemes to improve the efficiency of SSAT while preserving robust accuracy.
+In particular, we emphasized that strategically selecting or generating more boundary-adjacent data points can significantly improve the efficiency of semi-supervised adversarial training algorithms, without losing their robustness advantages.
+Among our proposed methods, we found that latent-based approaches using k-means clustering, including LCS-KM and LCG-KM, perform the best, as they reduce computational costs and memory requirements.
+Limitation & Future Work.
+We acknowledge that our work has several limitations.
+While the selection scheme with LCS-KM consistently achieves the best performance, its success requires carefully choosing a few hyperparameters (e.g., to avoid overfitting).
+Future studies can leverage AutoML algorithms to automatically determine these hyperparameters.
+Additionally, it remains unclear whether our best-performing latent-clustering-based methods (i.e., LCS-KM and LCG-KM) can be further optimized to achieve better performance.
+Theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful.
+Furthermore, our guided DDPM fine-tuning approach proposes to adapt a pre-trained model for boundary-seeking data generation.
+An interesting future direction is to investigate whether training-free sampling-based guidance [34] or more advanced generative models [35, 36] can enhance the efficiency and robustness of SSAT.
+Despite these limitations, we believe our work can serve as an important initial step towards the more efficient and scalable development of robust models.
+Availability
+The implementation of the latent-based data reduction methods and all our experiments are available as open-source code at: https://github.com/TrustMLRG/EfficientSSAT.
+Appendix A Related Work
+A-A Adversarially Robsut Learning
+Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [1].
+The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications.
+Earlier-proposed attacks like fast gradient sign method (FGSM) [2] and projected gradient descent (PGD) [7] introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks [37], which search for the minimal perturbation that causes misclassification.
+Recent attacks like Auto-PGD [38] automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack [17] is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
 Adversarial Training.
-We explore the impact of replacing TRADES with vanilla AT as the training algorithm.
-Similarly, we use external data from TinyImages as the unlabeled dataset.
-The results for PGD-based AT are shown in Figure LABEL:fig:-pgd.
-Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure LABEL:fig:pgd-5.
-Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
-$\bm{\ell}_{2}$ Perturbations.
-We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$.
-In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM.
-We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data.
-Our findings are shown in LABEL:fig:externall2 and LABEL:fig:generatedl2, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$.
-These results again confirm the efficiency and effectiveness of LCS-KM when adopted in self-supervised adversarial training algorithms.
-6.4. Low Labeled Data Regime
-We further study how the choice of the intermediate model and the amount of labeled data impact the final performance.
-For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with 10% of data selected from a pool of $500$K unlabeled images sourced from Tiny-ImageNet.
-This additional data is selected using the intermediate model.
-The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied.
-The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low.
-Additionally, we include a pre-trained CLIP model (Radford et al., 2021) as an alternative intermediate model for comparison.
-The results of these experiments are presented in Table 2.
-The intermediate model’s quality significantly impacts the final model’s robust accuracy.
-In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model.
-However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model.
-This trend demonstrates the effect of the intermediate model on the final robust accuracy.
-These observations underscore the importance of having a strong intermediate model for robust performance.
-Additionally, we find that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-label regimes.
-Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix D.3.
-6.5. Further Discussion
-Our study introduces strategic data selection methods, including Latent Clustering-Based Selection with k-means (LCS-KM), Latent Clustering-Based Selection with Gaussian Mixture Models (LCS-GMM), and prediction confidence-based selection, to improve the efficiency of self-supervised adversarial training.
-These approaches tackle key challenges in adversarially robust learning, particularly the significant memory and computational overhead of existing methods.
-By prioritizing boundary-adjacent data, these processes aim to enhance model robustness.
-However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models.
-While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset.
-By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization.
-Nevertheless, deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization.
-Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
-7. Conclusion
-In this work, we illustrated the importance of latent clustering-based data selection in improving the efficiency of SSAT while maintaining robust accuracy.
-Emphasizing boundary data points during selection significantly enhances the model’s overall robustness.
-For SVHN, we also showed improvement in robust accuracy by adding ground-truth labels to a small set of data.
-This suggests manually annotating a few unlabeled data points close to the boundary could be a potential future direction to further enhance model robustness with considerations of memory and computational efficiency.
-Our work opens new avenues for further research, including exploring advanced data selection schemes, making the selection process more interpretable, and investigating the trade-offs between model performance, data selection, and computational costs.
-Appendix A Overall Training Pipeline
-Figure 6 illustrates the overall pipeline, based on the robust self-training framework proposed by Carmon et al.
-(Carmon et al., 2019).
-We first train our intermediate model using the available labeled data and use this intermediate model to select a subsection of our unlabelled or generated data.
-The final robust training is performed on all the labeled data and a strategically selected subset of the unlabelled or generated data.
-Figure 6.
-Overview of the training pipeline.
-An intermediate model is first trained with labeled data and then used to select data from the unlabeled set.
-Finally, we adopt SSAT to learn a robust model on both the labeled and the reduced unlabeled datasets.
+On the defense side, adversarial training and its variants are the most popular methods for achieving robustness.
+For instance, Madry et al.
+proposed to train DNNs on adversarially perturbed inputs using PGD attacks [7], while Zhang et al.
+introduced TRADES [9], which balances the trade-off between robustness and standard accuracy.
+Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications.
+In particular, Rice et al.
+illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training [28], where early stopping is shown to be effective in alleviating such an issue.
+Techniques, such as Free AT [18] and Fast AT [19], offer promising solutions to reduce the computation for adversarial training.
+Nevertheless, they suffer from decreased robustness performance.
+Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently.
+For example, Zhang et al.
+argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [16].
+As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data.
+In a similar vein, Hua et al.
+advocated for a more targeted approach [24], recommending that PGD training be applied primarily to examples situated near the decision boundaries.
+This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks.
+Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
+Self-Supervised Adversarial Training.
+Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research [39].
+In the adversarial context, Schmidt et al.
+[10] highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning.
+They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly.
+Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training [11, 40, 41, 12].
+In particular, Carmon et al.
+proposed robust self-training methods to address the sample complexity issue [11], which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data.
+The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model.
+Concurrently, Najafi et al.
+proposed an optimization framework incorporating labeled and unlabeled data [41], while providing formal guarantees on adversarial robustness.
+Leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature [13, 14, 15].
+For instance, Gowal et al.
+employed various unconditional generative models [13], including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further.
+While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement.
+In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
+A-B Data-Efficient Deep Learning
+Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information.
+Two prominent approaches are dataset distillation and coreset selection.
+Dataset Distillation.
+Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points.
+For instance, Wang et al.
+proposed a foundational approach to dataset distillation [42], proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance.
+Later, Zhao et al.
+introduced the concept of gradient matching [43], where distilled data points are optimized to match the gradients of the original dataset.
+This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data.
+Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [44], facilitating the training of complex models with minimal data while maintaining strong performance.
+Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands.
+Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
+Coreset Selection.
+On the other hand, coreset selection [45, 46] aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset.
+These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning.
+For example, Kaushal et al.
+utilized diverse models to select training data subsets to reduce labeling efforts [20], while Xia et al.
+introduced the concept of “moderate coreset” [21], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions.
+Similarly, Mirzasoleiman et al.
+proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [22].
+In the adversarial context, Dolatabadi et al.
+proposed adversarial coreset selection [47], offering a task-specific solution to enhance the computational efficiency of adversarial training.
+In particular, the method constructs a compact data subset by minimizing the gradient approximation error.
+The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals.
+While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges.
+These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time.
+In such settings, repeatedly selecting data during training may become computationally infeasible.
+In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
 Appendix B Detailed Experimental Setup
 SVHN.
-SVHN is naturally divided into a core training set comprising approximately $73$K images and an extra training set of around $531$K images.
-Initially, the model is trained on the $73$K labeled images.
-To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model (Gowal et al., 2021).
+SVHN is naturally divided into a core training set comprising approximately 7373K images and an extra training set of around 531531K images.
+Initially, the model is trained on the 7373K labeled images.
+To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model [13].
 Configuration.
-For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture (Zagoruyko and Komodakis, 2016) in all our SVHN experiments.
+For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture [48] in all our SVHN experiments.
 We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al.
-(Zhang et al., 2019), with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$.
+[9], with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$.
 Hyperparameters are set the same as in Carmon et al.
-(Carmon et al., 2019) except for the number of epochs.
+[11] except for the number of epochs.
 For SVHN, we use a training batch size of $128$.
-All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with $1$M generated data are run for $400$ epochs.
+All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with 11M generated data are run for $400$ epochs.
 To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping.
-Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
+Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
 Evaluation.
 For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al.
-(Carmon et al., 2019) for better comparison.
+[11] for better comparison.
 We use step size $0.005$, number of attack steps $K=100$.
@@
 We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data.
-However, most images in $80$M-TI do not correspond to CIFAR-10 image categories.
+However, most images in 8080M-TI do not correspond to CIFAR-10 image categories.
 Carmon et al.
-(Carmon et al., 2019) used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model.
-Each class selected an additional $50$K images from 80M-TI using the model’s predicted scores to create a $500$K pseudo-labeled dataset, which we used in our experiments.
-We train the intermediate model with $50$K labeled data and use this model to select data from the $500$K pseudo-labeled data or a synthetic dataset.
+[11] used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model.
+Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments.
+We train the intermediate model with 5050K labeled data and use this model to select data from the 500500K pseudo-labeled data or a synthetic dataset.
 We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset.
-We also conducted further experiments utilizing $1$M synthetical data generated by the DDPM model (Gowal et al., 2021).
+We also conducted further experiments utilizing 11M synthetical data generated by the DDPM model [13].
+For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [49] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
 Configuration.
-For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) (Zagoruyko and Komodakis, 2016) architecture to train models.
+For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models.
 Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$.
 For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$.
-In section 6.3 when $\epsilon=4/255$, we use $0.0035$ as the PGD step size.
+In section D-A when $\epsilon=4/255$, we use $0.0035$ as the PGD step size.
 For $\epsilon=16/255$, we use $0.015$ as the PGD step size.
 Hyperparameters used are the same as in Carmon et al.
-(Carmon et al., 2019) except for the number of epochs.
+[11] except for the number of epochs.
 Here, we use a training batch size of $256$.
-The experiments using all the $500$K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs.
+The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs.
 For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process.
-Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
+Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
 Evaluation.
 The PGD attack evaluation is conducted similarly to that of Carmon et al.
-(Carmon et al., 2019) for fair comparisons.
+[11] for fair comparisons.
 We use step size $\alpha=0.01$, number of attack steps $K=40$.
@@
 Medical Application.
-Below, we detail the setup employed for our experiments on the medical datasets outlined in Section 5.4.
-For this experiment, we trained models using a ResNet-18 (He et al., 2016) architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$.
+Below, we detail the setup employed for our experiments on the medical datasets outlined in Section V-A.
+For this experiment, we trained models using a ResNet-18 [50] architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$.
 We set the attack step size to $0.02$, and the number of PGD steps to $10$.
-The final model is trained with SSAT using the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$.
+The final model is trained with SSAT using the Adam optimizer [51] with a learning rate of $0.001$.
 We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
-Table 3.
-Comparisons of SSAT methods with different selection schemes based on average efficiency score and robust accuracy.
-Here, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the “No Selection” method, we use $\alpha=100\%$, taking the average of the external and generated data results for each setting.
-[…]
-SVHN CIFAR-10
-Random $1.00$ $83.1\pm 1.04$ $56.8\pm 0.59$
-PCS $0.89$ $84.0\pm 0.92$ $57.4\pm 0.53$
-LCS-GMM $0.87$ $84.2\pm 0.89$ $57.8\pm 0.67$
-LCS-KM $0.88$ $86.6\pm 0.33$ $59.3\pm 1.15$
-No Selection $0.33$ $86.2\pm 0.15$ $62.2\pm 0.35$
-Appendix C Detailed Comparisons of Different Data Selection Schemes
-This section presents a comprehensive comparison of all the selection methods, including the case where no selection is applied.
-It highlights the tradeoff between performance and efficiency, offering a detailed analysis of the conditions under which each method is most suitable.
-Table 3 summarizes the performance of SSAT with different data selection strategies in terms of efficiency and robust accuracy.
-Each method strikes a balance between computational cost and accuracy improvements.
-Random selection serves as a baseline with minimal computation but lacks precision, while our strategic selection methods prioritize boundary points, which achieve progressively improved performance but with moderate increases in computational demands.
-Using $100\%$ of the data without selection yields the highest robust accuracy but at a significant computational cost.
-More specifically, we calculate a normalized efficiency score for each method defined as $\text{Eff Score}_{ij}=\frac{T_{\text{min},j}}{T_{ij}}$, where $T_{\text{min},j}$ is the minimum training time for dataset $j$, and $T_{ij}$ is the actual training time for method $i$.
-The overall efficiency score for method $i$ is the average across all datasets computed as $\frac{1}{n}\sum_{j}\text{Eff Score}_{ij}$, where $n$ denotes the number of datasets and sources considered.
-Effectiveness is measured by the average robust accuracy across CIFAR-10 and SVHN.
-In particular, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the no-selection method, we use $\alpha=100\%$.
-In all cases, we take the average of the results from both the external and generated data.
-These metrics are introduced to provide a fair comparison between methods.
-Appendix D Additional Experiments
-D.1. Other Hyperparameter Analysis Details and Results
-In Section 6.2, we present a detailed analysis of how accuracies are influenced by different hyperparameter values.
-In these experiments, we varied one hyperparameter at a time while keeping the others constant.
-Figure LABEL:fig:betacifar illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the $1$M data samples generated by the DDPM model, using a per-batch ratio of $0.3$.
-Figure LABEL:fig:betasvhn shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
-In Figure LABEL:fig:ratiocifar, we present how robust accuracy for the CIFAR-10 dataset is affected by varying the ratio of per-batch ratio $\gamma$ when $20\%$ extra data from the $1$M generated DDPM samples is used, with $\beta$ fixed at $0.4$.
-Similarly, Figure LABEL:fig:ratiosvhn displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$.
-In addition, Figures LABEL:fig:cofarper and LABEL:fig:svhnper explore how robust accuracy varies with the amount of data used.
-For CIFAR-10 in Figure LABEL:fig:cofarper, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$.
-For the SVHN dataset in Figure LABEL:fig:svhnper, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$.
-We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures LABEL:fig:cofarper and LABEL:fig:svhnper, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
-Table 4.
-Comparison results of self-supervised adversarial training algorithms with varying data selection schemes on SVHN.
-Instead of pseudo labels, the selected unlabeled data are assigned with the ground-truth SVHN labels provided in the dataset.
-[…]
+Appendix C Additional Experiments
+(a) CIFAR-10
+(b) SVHN
+Figure 7: Illustration of clean and robust accuracy curves of SSAT algorithms with LCS-KM by varying per-batch ratio $\gamma$, which defines how much labeled and unlabeled data are included in each mini-batch, on: (a) CIFAR-10 with $20\%$ extra unlabeled data selected from 11M generated images using DDPM, and (b) SVHN with $10\%$ selected from the external 531531K SVHN images.
+TABLE IV: Comparison results of semi-supervised adversarial training algorithms with varying data selection schemes on SVHN.
+Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
+$\textbf{Ratio }\bm{\alpha}$ Method Clean Accuracy ($\%$) Robust Accuracy ($\%$) Number of Epochs
 $1\%$ LCS-KM $95.6$ $82.9$ $75$
@@
 $100\%$ No Selection $97.5$ $86.4$ $200$
-D.2. SVHN with Ground-Truth Labels
-We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data.
-Table 4 shows the comparison results.
-With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
-D.3. Pre-Trained CLIP as Intermediate Model
-In all our previous experiments, we began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$.
-This intermediate model was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset.
-However, when dealing with generated or external data that already comes with labels, training an intermediate model specifically for selection undermines the selection process’s intended benefits, primarily to reduce computational and time complexity.
-In addition, in application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
-To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings (Radford et al., 2021).
+C-A Hyperparameter Sensitivity Analysis Details
+In Section V-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values.
+In these experiments, we varied one hyperparameter at a time while keeping the others constant.
+Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$.
+Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
+In Figure 7a, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$.
+Similarly, Figure 7b displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$.
+In addition, Figures 4c and 4d explore how robust accuracy varies with the amount of data used.
+For CIFAR-10 experiments in Figure 4c, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$.
+For SVHN in Figure 4d, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$.
+We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures 4c and 4d, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
+(a) CIFAR-10
+(b) SVHN
+Figure 8: Comparisons of SSAT methods with varying amounts of unlabeled data selected using LCS-KM with the CLIP model on (a) CIFAR-10 and (b) SVHN.
+We report standard and robust accuracies on the labeled test dataset.
+C-B Pre-Trained CLIP as Intermediate Model
+In our experiments, we always began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$.
+$f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset.
+However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the data selection process’s intended benefits, primarily to reduce computational and time complexity.
+In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
+To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings [52].
 CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training.
@@
 Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies.
-Figure LABEL:fig:comparison_clip shows our result when using CLIP instead of the previous intermediate model for data selection.
-D.4. Resnet-18 Model Architecture
-Our main experiments focus on the WidResNet architecture.
+Figure 8 shows our result when using CLIP instead of the previous intermediate model for data selection.
+C-C SVHN with Ground-Truth Labels
+We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data.
+Table IV shows the comparison results.
+With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
+C-D Resnet-18 Model Architecture
+Our main experiments focus on the WideResNet architecture.
 We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model.
-Similar to our previous experiments, we train the ResNet-18 model on $50$K labeled samples from CIFAR-10 and a chosen subset of data from $500$K unlabeled images of Tiny-ImageNet.
-Our evaluation results are shown in Table 5, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
-Table 5.
-Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations.
-The model training dataset consists of $50$K labeled CIFAR-10 images and $500$K unlabeled images drawn from Tiny-ImageNet.
-[…]
+Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet.
+Our evaluation results are shown in Table V, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
+TABLE V: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations.
+The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
+$\textbf{Ratio }\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
 $1\%$ Random $79.8$ $50.6$
@@
 $100\%$ No Selection $83.2$ $54.7$
+Appendix D Extended Analyses
+In this section, we further evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section D-A), and examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section D-B).
+Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table I.
+(a) TRADES ($\epsilon=4/255$, PGD-10)
+(b) TRADES ($\epsilon=16/255$, PGD-10)
+(c) AT ($\epsilon=8/255$, PGD-5)
+(d) AT ($\epsilon=8/255$, PGD-10)
+Figure 9: Illustration of SSAT performance on CIFAR-10 using LCS-KM under $\ell_{\infty}$ perturbations under various configurations: TRADES using $10$ PGD steps with (a) $\epsilon=4/255$ and (b) $\epsilon=16/255$; AT with $\epsilon=8/255$ using (c) 55 PGD steps and (d) $10$ PGD steps.
+TABLE VI: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data.
+We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
+Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
+Training Method Clean Acc ($\%$) Clean Acc ($\%$) Robust Acc ($\%$)
+$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
+pre-trained CLIP $86.8$ $83.9$ $40.5$
+$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
+pre-trained CLIP $86.8$ $85.3$ $43.2$
+$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
+pre-trained CLIP $86.8$ $85.6$ $47.7$
+100% fully-supervised $90.5$ $86.2$ $58.2$
+pre-trained CLIP $86.8$ $86.5$ $57.4$
+D-A Further Exploration of Generalizability
+We compare four different settings of SSAT on CIFAR-10: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data ($\alpha=0\%$), and using the entire dataset ($\alpha=100\%$).
+Perturbation Size.
+We consider external data from TinyImages as the unlabeled dataset.
+We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table I.
+The results are depicted in Figures 9a and 9b, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data.
+The trends are shown to be consistent across varying perturbation sizes.
+Adversarial Training.
+We explore the impact of replacing TRADES with vanilla AT as the training algorithm.
+Similarly, we use external data from TinyImages as the unlabeled dataset.
+The results for PGD-based AT are shown in Figure 9d.
+Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure 9c.
+Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
+(a) External data
+(b) Generated data
+Figure 10: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$: (a) external TinyImages and (b) DDPM-generated data.
+$\bm{\ell}_{2}$ Perturbations.
+We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$.
+In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM.
+We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data.
+Our findings are shown in 10a and 10b, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$.
+These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
+D-B Low Labeled Data Regime
+We further study how the choice of the intermediate model and the amount of labeled data impact the final performance.
+For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet.
+This additional data is selected using the intermediate model.
+The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied.
+The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low.
+In addition, we include a pre-trained CLIP model [52] as an alternative intermediate model for comparison.
+The results of these experiments are presented in Table VI.
+The intermediate model’s quality significantly impacts the final model’s robust accuracy.
+In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model.
+However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model.
+This trend demonstrates the effect of the intermediate model on the final robust accuracy.
+These observations underscore the importance of having a strong intermediate model for robust performance.
+Besides, we observe that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-labeled-data regimes.
+Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
+D-C Further Discussion
+Our work introduces several data reduction strategies that tackle the key challenges in adversarially robust learning, particularly the significant memory and computational overhead of state-of-the-art SSAT methods.
+By prioritizing boundary-adjacent data, these processes aim to enhance model robustness.
+However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models.
+While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset.
+By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization.
+Nevertheless, a deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization.
+Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
+Reducing the overall time required for adversarial training is critical for improving the efficiency and scalability of robust learning.
+Broadly, there are two complementary strategies to achieve this: (i) decreasing the computational cost of each training epoch, and (ii) reducing the total number of epochs needed for convergence.
+The first strategy aims to make each epoch faster without sacrificing robustness.
+Several prior works [18, 19, 53, 54] focus on this direction by introducing more efficient attack generation and optimization schemes.
+These approaches demonstrate that careful algorithmic design can substantially reduce per-epoch runtime, making adversarial training far more practical in large-scale settings.
+However, as shown in previous studies [15, 13], the total number of training epochs grows sharply with the size of the dataset.
+Consequently, even if the cost per epoch is reduced, increasing the amount of training data can still lead to a dramatic rise in overall training time.
+For example, adversarial training with approximately 11M generated samples typically requires around $400$ epochs to converge, whereas training with 100100M samples may demand as many as $2000$ epochs.
+Developing strategies that mitigate this dependence so that large-scale training does not necessitate a proportional increase in epochs would provide substantial computational savings and further improve the practicality of adversarially robust learning.

### v3

@@
 Abstract
-Achieving high model robustness under adversarial settings is widely recognized as demanding considerable training samples.
-Recent works propose semi-supervised adversarial training (SSAT) methods with external unlabeled or synthetically generated data, which are the current state-of-the-art.
+Learning robust models under adversarial settings is widely recognized as requiring a considerably large number of training samples.
+Recent work proposes semi-supervised adversarial training (SSAT), which utilizes external unlabeled or synthetically generated data and is currently the state of the art.
 However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage.
-In this paper, we propose unlabeled data reduction strategies to improve the efficiency of SSAT.
-Specifically, we design novel latent clustering-based techniques to select or generate a small critical subset of data samples near the model’s decision boundary.
-While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting.
-Comprehensive experiments on benchmark datasets demonstrate that our methods can significantly reduce SSAT’s data requirement and computation costs while preserving its strong robustness advantages.
-In particular, our latent-space selection scheme based on k-means clustering and our guided DDPM fine-tuning approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data and approximately $4\times$ less total runtime.
+In this paper, we propose data reduction strategies to improve the efficiency of SSAT by optimizing the amount of additional data incorporated.
+Specifically, we design novel latent clustering-based techniques to select or generate a small, critical subset of data samples near the model’s decision boundary.
+While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points, thereby avoiding overfitting.
+Comprehensive experiments across image benchmarks demonstrate that our methods can effectively reduce SSAT’s data requirements and computational costs while preserving its strong robustness advantages.
+In particular, our latent-space selection scheme based on k-means clustering and our guided diffusion-based approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data.
+When compared to full SSAT trained to convergence, our methods reduce total runtime by approximately $3\times$ to $4\times$ due to strategic prioritization of unlabeled data.
 I Introduction
-Deep neural networks (DNNs) are highly vulnerable to adversarially perturbed inputs [1, 2].
-This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness [2, 3, 4, 5].
-Among them, adversarial-training-based algorithms are most popular [7, 8, 9].
+Deep neural networks (DNNs) are shown to be extremely vulnerable to adversarially perturbed inputs [40, 15].
+Such a phenomenon has raised serious concerns about the reliability of deploying DNNs in security- or safety-critical applications and has driven numerous research into designing methods to enhance model robustness [15, 31, 3, 2].
+Among them, adversarial training and its variants are most popular [27, 43, 50].
 As illustrated by Schmidt et al.
@@
 To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [5, 49] or generated synthetic data [17, 35, 44].
-Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store the extra data and a much longer training time for algorithms like adversarial training to converge (see Appendix A for detailed discussions on related work).
+Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically require vast amounts of additional data, suggesting the need for much larger hardware to store the extra data and a longer training time for algorithms like adversarial training to converge.
 Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy.
@@
 [51], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary.
-Specifically, we introduce three data selection strategies, including PCS, a simple approach that prioritizes unlabeled data points with low prediction confidence, and two advanced latent clustering-based selection schemes: LCS-KM and LCS-GMM, where we identify the most uncertain data in the model’s latent embedding space based on the corresponding clustering technique.
+Specifically, we introduce three selection strategies, including PCS, a simple approach that prioritizes data points with low prediction confidence, and two advanced latent clustering-based schemes: LCS-KM and LCS-GMM, where we identify the most uncertain region in the model’s latent embedding space using clustering techniques.
 The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset.
-In addition, we propose novel DDPM fine-tuning approaches with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models.
-Our data reduction schemes balance the ratio of boundary critical points and the remaining non-boundary points to avoid potential overfitting to the boundary distribution.
-By applying the aforementioned data reduction schemes, the high unlabeled data and computation requirements of SSAT can be significantly alleviated while largely preserving its robustness performance.
+In addition, we propose a novel generative method by fine-tuning a denoising diffusion probabilistic model (DDPM) with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models.
+All our techniques are designed to be lightweight while balancing the ratio of boundary-critical points and the remaining non-boundary points, thereby reducing overall computational costs and avoiding potential overfitting to the boundary distribution.
+By applying our data reduction techniques, the high requirements for unlabeled data in SSAT can be largely alleviated.
 Contributions.
-We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section II).
-To realize this goal, we propose to optimize the performance of SSAT by refining the model’s decision boundary in the regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points (Section III-A).
-Specifically, we propose various data selection schemes to identify boundary-adjacent points (Section III-B), including a simple method based on prediction confidence and two advanced selection techniques by clustering the unlabeled or generated data in the latent embedding space.
-Additionally, we introduce three guided DDPM fine-tuning approaches (Section III-C), aiming to generate the subset of desirable data directly, thus avoiding the computation overhead of pre-generating the whole extra dataset.
-Comprehensive experiments on image benchmarks demonstrate that our data selection schemes, especially for the best-performing LCS-KM approach, significantly improve the data and computational efficiency of baseline SSAT methods, while achieving comparable robust accuracies (Section IV-A).
-For instance, utilizing $10\%$ extra unlabeled data identified by LCS-KM in SSAT can achieve similar robustness performance on SVHN compared to using the entire dataset, while reducing the total runtime to approximately $25\%$.
-By directly generating the exact amount of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can reduce the computation overhead while maintaining robustness (Section IV-C).
-Moreover, we demonstrate the generalizability of our data reduction schemes to a real-world COVID-related medical application (Section V-A) and visualize the distributions of selected data points to further illustrate why our LCS-KM method always performs the best (Section V-B).
-Finally, we provide general guidelines on how to choose the key hyperparameters for maximizing the SSAT performance (Section V-C), and conclude with more discussions on related literature, limitations of our study, and future research directions (Section VI).
-Our work reveals the importance of prioritizing boundary-adjacent points for improving SSAT’s efficiency and highlights the superiority of latent-based k-means clustering, potentially assisting the development of more efficient and scalable robust learning algorithms for real-world applications.
-II Inefficiency of Existing SSAT Methods
-II-A Semi-Supervised Adversarial Training
+We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT algorithms (Section III).
+To achieve this goal, we propose optimizing SSAT performance by refining the model’s decision boundary in regions of high uncertainty and strategically prioritizing boundary-adjacent data points (Section IV-A).
+Specifically, we introduce different selection schemes to identify boundary-adjacent points (Section IV-B), including a straightforward method based on prediction confidence and two advanced selection techniques that utilize clustering of unlabeled or generated data in the model’s latent embedding space.
+In addition, we propose three guided DDPM fine-tuning approaches (Section IV-C) to directly generate a subset of desirable data, thereby avoiding the computational overhead of pre-generating the entire dataset.
+Comprehensive experiments on image benchmarks demonstrate that our data selection strategies, particularly for the best-performing LCS-KM approach, significantly improve data efficiency compared to baseline SSAT methods, while achieving comparable test robust accuracies and reducing overall computation costs (Section V-A).
+For instance, selecting $10\%$ extra DDPM-generated data based on LCS-KM in the SSAT algorithm can achieve robustness performance on SVHN similar to that of using the entire unlabeled dataset.
+Compared to full SSAT trained to convergence, this reduces the total runtime to approximately $25\%$ due to faster convergence enabled by the strategically selected data.
+By directly generating the exact number of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can further reduce computational overhead while maintaining robustness (Section V-C).
+Moreover, we demonstrate the generalizability of our data reduction schemes to a COVID-related medical application task (Section VI-A) and visualize the distributions of selected data points to illustrate why our LCS-KM method consistently performs the best (Section VI-B).
+We conclude with the limitations of our study and future directions (Section VII).
+Our work highlights the importance of prioritizing boundary-adjacent points and latent-based k-means clustering in SSAT, potentially assisting in the development of more efficient, robust learning algorithms for real-world applications.
+II Related Work
+II-A Adversarially Robust Learning
+Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [40].
+Earlier attacks, such as fast gradient sign method (FGSM) [15] and projected gradient descent (PGD) [27], proposed gradient-based constrained optimization to find adversarial examples.
+Carlini and Wagner introduced CW attacks to search for minimal but unconstrained perturbations that cause misclassifications [4].
+Recent Auto-PGD attacks [10] automate gradient steps to more efficiently produce worst-case perturbations, while AutoAttack [8] is an ensemble of diverse, parameter-free attacks, making them most reliable for evaluating the robustness of DNNs.
+Adversarial Training.
+To attain robustness, adversarial training (AT) and its variants are most popular.
+For instance, Madry et al.
+proposed to train DNNs on adversarial inputs using PGD attacks [27], while Zhang et al.
+introduced TRADES [50], which balances the trade-off between robustness and standard accuracy.
+Nevertheless, adversarial training algorithms are often criticized for overfitting to adversarially perturbed samples produced during training and for their high computational costs, which hinder their deployment in real-world applications.
+In particular, Rice et al.
+empirically demonstrated that robust overfitting is a prevalent phenomenon in adversarial training [33], where early stopping is suggested as an effective approach for mitigating this issue.
+Techniques, such as Free AT [37] and Fast AT [45], offer promising solutions to reduce the computational burden of adversarial training.
+However, they inevitably suffer from decreased robustness performance.
+Moreover, a line of research has pointed out that treating all data points equally in adversarial training is not optimal, as different examples can impact model robustness differently.
+For example, Zhang et al.
+argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [51].
+Consequently, they should be given less importance during training to prevent the model from over-optimization, which could negatively affect its generalization on unseen test data.
+In a similar vein, Hua et al.
+advocated for a more targeted approach [22], recommending PGD training applied primarily to examples situated near the model’s decision boundaries.
+The above techniques enhance the model’s adversarial robustness, as perturbations on boundary-adjacent data are more informative in guiding the optimization process of adversarial training.
+Our work builds on these ideas by strategically prioritizing boundary points from the unlabeled dataset to address the unique challenges associated with SSAT algorithms.
+SSAT.
+Utilizing additional unlabeled data to improve model performance is an active field of research [6].
+In the adversarial context, Schmidt et al.
+[34] highlighted a key observation: robust generalization often necessitates a significantly larger training set size to attain a similar level of performance compared to standard learning.
+To address the sample complexity challenge, a line of work has proposed semi-supervised adversarial training (SSAT), which utilizes extra unlabeled data, often acquired from a similar but slightly different distribution [5, 1, 29, 49].
+In particular, Carmon et al.
+introduced the robust self-training framework [5], which first trains an intermediate model using available labeled data and then uses the model to generate pseudo-labels for unlabeled data.
+The pseudo-labeled samples, combined with the original labeled dataset, are subsequently used to train a final robust model.
+Concurrently, Najafi et al.
+proposed an optimization framework that incorporates both labeled and unlabeled data [29], providing formal guarantees on adversarial robustness.
+Leveraging synthetic data produced by state-of-the-art generative models to build robust models has been explored in more recent adversarial ML literature [17, 35, 44].
+For instance, Gowal et al.
+employed various unconditional generative models, including DDPM [21], to produce a large synthetic dataset for incorporation in adversarial training [17].
+While enhancing robustness, all the aforementioned SSAT methods require a considerable amount of additional unlabeled data.
+In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data, aiming to achieve both efficiency and effectiveness in semi-supervised adversarial training.
+II-B Data-Efficient Deep Learning
+Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information.
+Two prominent approaches are dataset distillation and coreset selection.
+Dataset Distillation.
+Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions.
+For instance, Wang et al.
+proposed a foundational approach to dataset distillation [42], condensing large datasets into smaller, synthetic subsets while preserving model performance.
+Later, Zhao et al.
+introduced the concept of gradient matching [52], where distilled data points are optimized to match the gradients of the original dataset.
+This technique enhances distillation performance, allowing models to learn more effectively from reduced data.
+In addition, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [39], facilitating the training of complex models with minimal data while maintaining strong performance.
+Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands.
+Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
+Coreset Selection.
+Coreset selection [36, 24] aims to identify a small, representative subset of data that retains the model performance when trained on it, often relying on diversity-based metrics or optimization techniques to ensure that the selected subset is sufficiently informative.
+For example, Kaushal et al.
+utilized diverse models to select training data subsets to reduce labeling efforts [23], while Xia et al.
+introduced the concept of moderate coreset [47], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions.
+Similarly, Mirzasoleiman et al.
+proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [28].
+Moreover, Dolatabadi et al.
+adapted the coreset selection technique to adversarial contexts, offering a task-specific solution to enhance the efficiency of adversarial training [12].
+In particular, they proposed adversarial coreset selection, which identifies a compact data subset by minimizing the gradient approximation error, and then utilizes it for adversarial training, with selection taking place at regular intervals.
+Compared with vanilla AT, the growing reliance on large volumes of generated or external unlabeled data for SSAT presents unique challenges.
+Due to the significantly expanded training dataset of SSAT, iteratively refining data selection during the entire training process is computationally infeasible.
+Our work complements the aforementioned methods for efficient SSAT, focusing on developing improved data reduction schemes that do not compromise robustness or scalability.
+III Inefficiency of Existing SSAT Methods
+III-A Semi-Supervised Adversarial Training
 We work with adversarially robust classification tasks and focus on semi-supervised adversarial training (SSAT), which was initially introduced to the field by Carmon et al.
@@
 For simplicity, we term $\gamma$ as the per-batch ratio.
-When $\gamma$ is set as $1$, Equation II-A reduces to the optimization objective of vanilla adversarial training (AT) [7].
-In Equation II-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss.
+When $\gamma$ is set as $1$, Equation III-A reduces to the optimization objective of vanilla adversarial training (AT) [27].
+In Equation III-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss.
 Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
@@
 Aligned with prior literature [27], we adopt multi-step PGD to approximately solve for the optimization problem in Equation 3.
-II-B Limitations of Prior Work
+III-B Limitation of Prior Work
 Data Inefficiency.
@@
 [17] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [27] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data.
-Training SSAT algorithms on the huge unlabeled dataset demands significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and carbon footprint.
-In addition, handling a larger dataset in ML often means increased GPU processing load and memory usage.
+Training SSAT algorithms on the vast, unlabeled dataset requires significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and a larger carbon footprint.
+In addition, handling a larger dataset in machine learning often means increased GPU processing load and memory usage.
 High Computation Cost.
-Apart from the data inefficiency issue, SSAT also requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of total training epochs compared with vanilla AT (see Figure 1 for supporting evidence).
-Intuitively, learning meaningful and robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more difficult.
-As our experiments illustrate, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples.
+Apart from the data inefficiency, SSAT requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times training epochs compared with vanilla AT (see Figure 1 for supporting evidence).
+Intuitively, learning meaningful, robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more challenging.
+As will be illustrated in our experiments, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples.
 Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster.
-We note that vanilla AT has already been criticized for its high computation [18, 19], due to the iterative PGD steps involved in dealing with the inner maximization problem (Equation 3).
-SSAT methods incur even higher computations, since they converge more slowly and usually adopt larger model architectures (see Table I for concrete runtime comparisons).
+We note that vanilla AT has already been criticized for its high computation costs [37, 45], due to the iterative PGD steps involved in solving the inner maximization problem (Equation 3).
+SSAT algorithms incur even higher computations, as they converge more slowly and often employ larger model architectures (see Table I for concrete runtime comparisons).
 Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
-II-C Problem Formulation
-Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for achieving high model robustness.
+III-C Problem Formulation
+Recognizing the data and computational inefficiencies of SSAT, we propose investigating whether the large size of unlabeled data is necessary for achieving high model robustness.
 To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
@@
 $\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big),\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (4)
-where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation II-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint.
+where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation III-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint.
 Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
-$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$ (5)
+$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right],$ (5)
 where $\ell_{0/1}$ denotes the 0-1 loss function.
 Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly.
-As we will illustrate in Section III-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
+As we will illustrate in Section IV-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
 Guided Diffusion.
@@
 where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model.
-As will be detailed in Section III-C, we adopt a pretrained DDPM [23] as the backbone while strategically fine-tuning it with novel training losses for guided data generations.
+As will be detailed in Section IV-C, we adopt a pretrained DDPM [21] as the backbone while strategically fine-tuning it with novel training losses for guided data generations.
 Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
-III Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
+IV Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
 So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement.
-In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section III-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections III-B and III-C).
-III-A Prioritize Boundary-Adjacent Unlabeled Data
+In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section IV-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections IV-B and IV-C).
+IV-A Prioritize Boundary-Adjacent Unlabeled Data
 Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity.
-Inspired by prior literature [16, 24] that emphasizes the unbalanced data importance for vanilla adversarial training (see Appendix A for more discussions), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT.
-In particular, we propose to identify or generate a small set of vulnerable yet valuable unlabeled data points, which are close to the model’s decision boundary.
+Inspired by prior literature [51, 22] that emphasizes the unbalanced data importance for adversarial training, we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT.
+In particular, we propose identifying or generating a small set of vulnerable yet valuable unlabeled data points that are close to the model’s decision boundary.
 Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify.
-Thus, improving their classification can yield more robustness enhancement to adversarial inputs.
-We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy function for locating difficult-to-classify data points.
-Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points.
-By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.
+Thus, improving their classification can yield enhanced robustness against adversarial inputs.
+We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to provide a meaningful proxy for locating difficult-to-classify data points.
+Since $f_{\hat{\theta}}$ is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide useful information for identifying these critical points.
+By focusing more on these critical boundary points, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.
 While a plausible approach is to select a set of unlabeled data based on how much their label predictions change under adversarial perturbations returned by PGD attacks, it is computationally expensive, undermining the efficiency gains we aim for.
@@
 Therefore, we seek better alternatives to effectively identify boundary data points without incurring high computational overhead.
-Algorithm 1 Prediction Confidence-Based Selection
-1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
-2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
-3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
-4:  Initialize $\mathcal{A}_{u}$ as empty set
-5:  $\mathrm{Conf}(\bm{x})\leftarrow$ get confidence of $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
-6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
-7:  $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with lowest $\mathrm{Conf}(\bm{x})$
-8:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
-9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
-10:  Output: selected subset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
-III-B Strategic Selection via Latent Clustering
+IV-B Strategic Selection via Latent Clustering
 PCS.
-To identify a small, critical subset of boundary-adjacent data points while accounting for computational efficiency and scalability, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point.
-The pseudocode for such a selection scheme is depicted in Algorithm 1.
-Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized.
+To identify a core subset of boundary-adjacent points while accounting for computational efficiency, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point.
+Algorithm 1 in Appendix A presents the pseudocode of this selection scheme.
+Initially, all the unlabeled data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized.
 The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection.
 Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes.
-The reason for involving such a trade-off parameter will be further discussed in Section V-C.
+The reason for involving such a trade-off parameter will be further discussed in Section VI-C.
 The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data.
@@
 In particular, we explore two classical clustering techniques for their unique benefits in identifying boundary points in the LCS framework.
-Algorithm 2 presents the pseudocode of the two proposed LCS methods.
-In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to capture local geometric structures around decision boundaries effectively.
+Algorithm 2 depicts the pseudocode of the methods.
+In particular, LCS with k-means (LCS-KM) generates latent representations of unlabeled data and clusters them based on Euclidean distances to the centroids, prioritizing data points that are equidistant from multiple centroids to effectively capture local geometric structures around decision boundaries.
 On the other hand, LCS with Gaussian mixture models (LCS-GMM) fits the latent representations to Gaussian mixture models, using posterior probabilities across multiple fitted Gaussians to identify points that are likely near decision boundaries.
 Compared to PCS, both LCS-KM and LCS-GMM leverage latent space clustering to provide a more accurate characterization of boundary vulnerabilities.
-Algorithm 2 Latent Clustering-Based Selection
-1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
-2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
-3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
-4:  Initialize $\mathcal{A}_{u}$ as empty set
-5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
-6:  if using LCS-KM then
-7:   $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}\leftarrow$ k-means clustering on $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
-8:   $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in\mathcal{S}_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
-9:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta d$
-10:  else if using LCS-GMM then
-11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
-12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
-13:   $\Delta p\leftarrow$ compute $\Delta p=|p_{1}-p_{2}|$ for each $\bm{z}$, where $p_{1}$ and $p_{2}$ are the top two highest probabilities
-14:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta\gamma$
-15:  end if
-16:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
-17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
-18:  Output: selected dataset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
 LCS-KM.
@@
 Finally, the top $\alpha|\mathcal{S}_{u}|$ unlabeled data with the smallest $\Delta p$ values are used to form the selected subset.
-As will be illustrated in Section IV, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=10\%$ or $\alpha=20\%$).
-More detailed discussions of the difference between these data selection schemes are provided in Section V-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
-Algorithm 3 SSAT with Guided DDPM Fine-Tuning
-1:  Input: labeled $\mathcal{S}_{l}$; pre-trained DDPM $\theta_{\mathrm{pre}}$; fine-tuning epochs $S$; trade-off parameter $\lambda$; other parameters $\epsilon$, $\alpha|\mathcal{S}_{u}|$, $\beta$, $\gamma$, $k$, $\eta$
-2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
-3:  Initialize $\mathcal{G}_{u}$ as empty set
-4:  $\mathcal{G}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points generated by pre-trained $\theta_{\mathrm{pre}}$
-5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
-6:  for $s=1,2,\ldots,S$ do
-7:   Sample batches of size $b$: $\bm{x}_{0}\sim\mathcal{S}_{l}$, $t\sim\mathcal{U}(T)$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I})$
-8:   $\bm{x}_{t}\leftarrow\sqrt{\bar{\alpha}_{t_{i}}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$
-9:   $L_{\mathrm{DDPM}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|_{2}^{2}$
-10:   if using PCG then
-11:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{PC}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
-12:   else if using LCG-KM then
-13:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{KM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
-14:   else if using LCG-GMM then
-15:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{GMM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
-16:   end if
-17:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\big(L_{\mathrm{DDPM}}+\lambda\cdot L_{\mathrm{reg}}\big)$
-18:  end for
-19:  $\mathcal{G}_{u}\leftarrow$ add $\beta\cdot\alpha|\mathcal{S}_{u}|$ points generated by fine-tuned $\theta$
-20:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)$
-21:  Output: generated dataset $\mathcal{G}_{u}$, final model $\theta_{\mathrm{final}}$
-III-C Guided Diffusion via Latent Clustering
-While our selection-based strategy can significantly reduce the unlabeled data set size and lower the overall training time, it has a critical drawback–when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement.
-As shown in Table II, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU.
-Note that the best current SSAT method documented in the RobustBench leaderboard [17] utilizes a few hundred million synthetically generated images, suggesting the total generation time can be a significant computational overhead.
+As will be illustrated in Section V, the robustness performance of SSAT can be largely maintained when using LCS-KM and LCS-GMM to select a small subset of unlabeled data (e.g., $\alpha=20\%$).
+More detailed discussions of the difference between the proposed data selection schemes are provided in Section VI-B, where we visualize the selected unlabeled data in a two-dimensional latent space.
+IV-C Guided Diffusion via Latent Clustering
+While our selection strategy can effectively reduce the unlabeled data set size and lower the training time, it has a critical drawback: when SSAT employs a generative approach, all our previous methods require generating the full synthetic unlabeled dataset as a prerequisite before selection, which is inefficient in terms of both generation time and storage requirement.
+As shown in Table III, generating 1M CIFAR-10-like synthetic images via DDPM requires around $4$ hours on a single A100 GPU.
+Note that the best current SSAT method, as documented in the RobustBench leaderboard [8], utilizes a few hundred million synthetically generated images, suggesting that the total generation time can be a significant computational overhead.
 Therefore, a natural question arises: whether pre-generating synthetic data and then applying strategic selection schemes can be further improved.
-In this section, we propose an alternative generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively.
-In particular, we fine-tune a pretrained DDPM using novel guidance losses for more efficient unlabeled data generation based on the insights from previous sections.
-In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the computation overhead.
+In this section, we introduce an alternative data generative pipeline, detailed in Algorithm 3, which answers the above question affirmatively.
+In particular, we fine-tune a pre-trained DDPM using novel guidance losses to achieve more efficient unlabeled data generation, building on the similar insights presented in the previous sections.
+In this way, we directly generate a small, critical set of boundary-adjacent data points, thereby avoiding redundant data generation and further reducing the overall computation overhead.
 DDPM.
-Diffusion model resembles a milestone in the history of generative modeling, which is the current state-of-the-art for high-quality image generation and synthesis.
+Diffusion models represent a milestone in the history of generative modeling, achieving state-of-the-art performance for high-quality image generation and synthesis.
 A denoising diffusion probabilistic model (DDPM) is a deep neural network trained to learn how to step-by-step reverse the forward process of adding Gaussian noise to an image [21].
@@
 A larger value of $\ell_{\mathrm{PC}}(\bm{x})$ suggests that $\bm{x}$ is more likely to be closer to the model’s decision boundary.
-As we’ve explained in Section III-B, the prediction confidence of a neural network classifier can be an unreliable estimator.
+As we’ve explained in Section IV-B, the prediction confidence of a neural network classifier can be an unreliable estimator.
 Therefore, we propose alternative guidance losses using more robust latent clustering-based approaches.
@@
 For LCG-KM, we first obtain the latent embeddings of all the labeled data with respect to the intermediate model, namely $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{l}\}$, and then apply k-means clustering algorithm to partition the embeddings into $k$ clusters.
-Note that this step is different from LCS-KM, because we now do not pre-generate the unlabeled dataset.
-Let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters.
-We define the guidance loss as:
+Note that this step differs from LCS-KM, as we no longer pre-generate the unlabeled dataset.
+Formally, let $\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{k}\}$ denote the centroids corresponding to the $k$ returned clusters, then we define the guidance loss for LCG-KM as:
 […]
@@
 By penalizing the model weights through the regularization term, our method gradually pushes $\theta$ towards higher probabilities of generating boundary-adjacent data using stochastic gradient descent (SGD).
-Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to bias the diffusion model to generate the desirable subset of synthetic images.
-IV Experiments
-In this section, we test the performance of our data reduction strategies proposed in Section III compared with baseline SSAT algorithms.
-In particular, we first conduct experiments to evaluate our selection-based methods in terms of robustness and efficiency advantages when integrated into the SSAT pipeline (Section IV-A).
-Then, we highlight the computational advantages of our approaches, particularly their capabilities of accelerating the convergence and reducing the total runtime of SSAT (Section IV-B).
-Finally, we study the effectiveness of our DDPM fine-tuning schemes in further improving the efficiency of SSAT algorithms under a generative setup (Section IV-C).
-To begin with, we lay out the main experimental setup.
-All our experiments are conducted using a single A100 GPU.
-All the remaining details, including all the hyperparameters, for reproducing our experiments are provided in Appendix B.
+Our later experiments show that when starting from a pre-trained DDPM (i.e., on CIFAR-10), a few fine-tuning epochs are sufficient to guide the diffusion model to generate the desirable subset of synthetic images.
+V Experiments
+In this section, we evaluate the performance of our data selection methods (Section V-A), discuss their computational advantages compared to full SSAT (Section V-B), and study the effectiveness of our guided DDPM fine-tuning schemes (Section V-C).
+All our experiments are conducted on a single A100 GPU.
+Below, we briefly introduce the main experimental setup, where full details are deferred to Appendix B.
 Dataset.
-We test two image benchmarks in our evaluation, SVHN [26] and CIFAR-10 [27], both widely used in prior adversarial ML literature.
-We consider two scenarios in terms of the source of the utilized unlabeled dataset $\mathcal{S}_{u}$ for testing our selection-based strategies: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model.
+We consider two benchmark image datasets, SVHN [30] and CIFAR-10 [26], and two scenarios in terms of the source of the unlabeled dataset $\mathcal{S}_{u}$ for evaluating our methods: (i) external–using an external dataset, and (ii) generative–using a synthetic dataset generated by some pre-trained generative model.
 For external scenarios, we follow the implementation protocols outlined by Carmon et al.
@@
 For CIFAR-10, the original dataset comprises 5050K labeled training images and 1010K labeled test images, where models trained using SSAT are augmented by an external unlabeled dataset $\mathcal{S}_{u}$ with 500500K images sampled from the 8080M Tiny Images (80M-TI) dataset.
-For generative setup, we adopt 11M synthetic images generated by DDPM for both datasets, following the protocols documented by Gowal et al.
+For the generative setup, we use 11M synthetic images generated by DDPM for both datasets, following the protocols in Gowal et al.
 [17].
 Configuration.
-Following prior literature, we mainly consider $L_{\infty}$-norm bounded perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10.
-We adopt WideResNet (WRN) architectures and train models using TRADES [9] on both labeled and unlabeled data, coupled with pseudo labels returned by the intermediate model.
+We consider $\ell_{\infty}$ perturbations with $\epsilon=0.015$ on SVHN and $\epsilon=0.031$ on CIFAR-10.
+We adopt WideResNet (WRN) architectures and train models using TRADES [50] on both labeled and unlabeled data, incorporating pseudo labels generated by the intermediate model.
 In Appendix D, we evaluate the generalizability of our methods for varying $\ell_{\infty}$ perturbation size, vanilla PGD-based adversarial training with varying attack steps, and $\ell_{2}$ perturbations.
@@
 Prior works also adopted this approach [5, 17], which ensures that training time per epoch remains consistent across datasets of varying sizes.
-SSAT.
-For experiments with external unlabeled data, SSAT algorithms are trained for a total of $200$ epochs, whereas experiments with generated data extend to $400$ epochs.
-To compare the performance of our methods in a fair way, models are saved every $25$ epoch of SSAT, and the model achieving the highest robust accuracy is selected as the best model.
-This approach aligns with the early stopping practices commonly used in adversarial ML literature [9, 28].
-In experiments with external data and ratio parameter $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN.
-Conversely, when no data selection scheme is applied, peak performance is delayed to approximately $200$ epochs.
-Total training time is reported as the duration required to achieve the best model performance.
-Intermediate Model.
-Our data selection and guided diffusion schemes require training an intermediate model $f_{\hat{\theta}}$, which we adopt the same architecture as the final model.
-This intermediate model is trained using standard supervised learning for $100$ epochs on a WRN-28-10 architecture, which takes around $55$ minutes.
-Notably, this time is excluded from the reported total runtime.
-The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our proposed strategies.
-As such, this training step is integral to the overall workflow, irrespective of whether a specific selection or fine-tuning method is employed.
-Evaluation Metric.
-To test the robustness of the final model returned by SSAT, we mainly use multi-step PGD attacks (PGD) [7], due to their broad adoption in adversarial ML.
-We also report robust accuracies against AutoAttack (AA) [29] for more rigorous robustness evaluation and the model’s clean accuracy on normal inputs (Clean) as a reference.
 TABLE I: SSAT performance under various unlabeled data selection schemes.
@@
 For each configuration, we underline the best robust accuracies.
-Dataset Ratio $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
+Dataset $\bm{\alpha}$ Method External Unlabeled DDPM-Generated
 Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h) Clean ($\%)$ PGD ($\%)$ AA ($\%)$ Time (h)
@@
 LCS-KM $88.7$ $\underline{60.7}$ $\underline{57.8}$ $14.78$ $85.5$ $\underline{60.2}$ $\underline{57.2}$ $15.27$
-IV-A Strategic Unlabeled Data Selection
-SVHN.
-Our evaluation results on SVHN are summarized in Table I, which demonstrate the efficacy of the proposed data selection schemes when employed in SSAT algorithms trained using a WRN-16-8 architecture with varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, with comparisons to baseline SSAT performance with $\alpha\in\{0\%,100\%\}$.
-Specifically, employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset.
-In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM.
-The robust accuracies evaluated using AutoAttack (i.e., the AA column of Table I) demonstrate similar patterns compared to those using multi-step PGD Attacks (i.e., the PGD column of Table I).
-Additionally, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights.
-Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases.
-We provide detailed sensitivity analyses of these hyperparameters in Section V-C.
-Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline.
-Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime.
-In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs.
-Such a large reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section IV-B.
-Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table IV in Appendix C).
-CIFAR-10.
-We adopt a WRN-28-10 architecture for CIFAR-10 experiments.
-To assess the generalizability of our findings, we also conduct ablations on ResNet-18, which are detailed in Table V in Appendix C.
-Across varying ratios of utilized unlabeled data $\alpha$ and different data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains robust accuracy of $60.7\%$, comparable to the $62.5\%$ obtained with the entire dataset.
-In comparison, random selection with the same ratio parameter achieves only $57.8\%$, underscoring the advantages of our strategic LCS technique with k-means.
-Similar to SVHN experiments, we extend our analysis to CIFAR-10 with synthetically generated data.
-Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset.
-These trends are also consistent across experiments with $\ell_{2}$ perturbations (see Figure 10 for supporting evidence).
-Our findings emphasize the robustness and adaptability of the proposed data selection schemes, especially for LCS-KM, even under varying experimental configurations.
-In terms of computational efficiency, CIFAR-10 experiments highlight the significant advantages of data selection.
-When leveraging selected external unlabeled data, optimal robust performance was achieved after $100$ epochs of training.
-By contrast, using the full unlabeled or generated datasets without selection required extended training durations of $200$ and $400$ epochs, respectively, to reach comparable performance.
-These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training.
-The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also substantially reduces computational costs and accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
+SSAT.
+With external data, we train SSAT algorithms for a total of $200$ epochs using stepwise learning rate scheduling, whereas training is extended to $400$ epochs under generative setups.
+To ensure fair comparisons, models are saved every $25$ epoch, and the model achieving the highest robust accuracy is selected as the best model.
+This approach aligns with the early stopping practices widely used in prior literature [50, 33].
+When the ratio of selected unlabeled data $\alpha\in\{10\%,20\%\}$, we consistently observe peak robust accuracy around $100$ epochs for CIFAR-10 and $75$ epochs for SVHN.
+Conversely, when no data selection is applied, peak robustness is delayed to approximately $200$ epochs.
+The total training time is reported as the duration required to achieve optimal performance.
+Intermediate Model.
+Our data selection and guided diffusion approaches require training an intermediate model $f_{\hat{\theta}}$, which adopts the same architecture as the final model.
+The intermediate model fulfills two critical roles: it not only facilitates the pseudo-labeling of unlabeled data but also constitutes a necessary component for implementing any of our data reduction strategies.
+To be more specific, we train the intermediate model $f_{\hat{\theta}}$ using standard supervised learning for $100$ epochs on WRN-28-10, which takes approximately $55$ minutes.
+We exclude this time from the reported total runtime, as this step is integral to the overall pipeline of any tested SSAT algorithm.
+Evaluation Metric.
+To evaluate the robustness of the final model returned by SSAT, we mainly employ multi-step PGD attacks (PGD) [27], due to their broad adoption in adversarial ML.
+We also report robust accuracies against AutoAttack (AA) [9] to provide a more rigorous robustness evaluation, as well as the model’s clean accuracy on normal inputs (Clean).
+V-A Strategic Unlabeled Data Selection
+Table I documents our evaluation results on SVHN and CIFAR-10, showing the efficacy of our data reduction schemes when incorporated into semi-supervised adversarial training under varying data ratios $\alpha\in\{1\%,10\%,20\%\}$, compared with baseline TRADES ($\alpha=0\%$) and full SSAT ($\alpha=100\%$).
 (a) Vanilla AT
@@
 Figure 1: Illustration of standard and robust accuracy curves of SSAT on CIFAR-10 with different configurations of unlabeled data selection from the external 500500K Tiny Images: (a) No extra data, (b) random selection with $\alpha=10\%$, (c) LCS-KM with $\alpha=10\%$, and (d) utilizing all 500500K unlabeled data.
-TABLE II: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
-Ratio $\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
+SVHN.
+Specifically for SVHN, we adopt a WRN-16-8 architecture.
+Employing our LCS-KM strategy with $10\%$ or $20\%$ of extra unlabeled data achieves PGD robust accuracies of $86.3\%$ and $86.6\%$, respectively, which matches and even slightly outperforms the $86.0\%$ robustness achieved using the entire unlabeled dataset.
+In contrast, random selection with $20\%$ of the data yields a robust accuracy of only $82.3\%$, underscoring the superior efficacy of informed selection strategies such as LCS-KM.
+The robust accuracies evaluated using AutoAttack demonstrate similar patterns compared to those using multi-step PGD Attacks.
+In addition, alternative methods, including PCS and LCS-GMM, consistently outperform random selection, confirming the effectiveness of our insights.
+Notably, the robustness of models trained using SSAT peaks when the boundary ratio parameter $\beta$ is inversely proportional to the selection ratio $\alpha$, suggesting that smaller $\beta$ values are favorable as the data inclusion proportion increases.
+We provide detailed sensitivity analyses of these hyperparameters in Section VI-C.
+Our evaluations with 11M DDPM-generated data highlight that when using $10\%$ to $20\%$ of unlabeled data selected via LCS-KM, SSAT can achieve comparable robustness to the SSAT baseline.
+Since all SSAT algorithms listed under the column “DDPM-Generated” in Table I require the same amount of pre-generation time, we exclude it from the total runtime.
+In particular, training SSAT with reduced data requires only $75$ epochs to achieve the peak robust accuracy, while training with the full dataset requires up to $400$ epochs.
+Such a reduction in computation time underscores the practical advantages of the proposed data selection strategies, which are further elaborated in Section V-B.
+Besides, we observe in our experiments that using ground-truth labels reveals negligible differences compared to pseudo-labeled data, indicating that the primary performance gains stem from the selection strategy itself rather than label accuracy (see Table V in Appendix C).
+CIFAR-10.
+We adopt a WRN-28-10 architecture for CIFAR-10 experiments.
+Across varying ratios $\alpha$ and data selection schemes, Table I reveals that selecting $20\%$ of the external data using LCS-KM attains PGD robust accuracy of $60.7\%$, close to the $62.5\%$ obtained with the entire dataset.
+In contrast, random selection with the same ratio achieves only $57.5\%$, underscoring the advantages of our strategic LCS technique.
+Similar to SVHN, we extend our analysis to CIFAR-10 with synthetically generated data.
+Using 11M DDPM-generated CIFAR-10-like images, we discover that using just $20\%$ of the generated data via LCS-KM results in performance nearly identical to that achieved with the full dataset.
+The above observations are also consistent across other settings, such as $\ell_{2}$ perturbations and a ResNet-18 model architecture (see Figure 9 and Table VI in the appendix for supporting evidence).
+The above findings emphasize the adaptability of our data selection schemes, particularly for LCS-KM, under varying experimental conditions.
+When leveraging a small subset of selected external unlabeled data, optimal robust performance was achieved at $100$ epochs of training.
+In contrast, using the full external or generated datasets without selection requires extended training of $200$ and $400$ epochs, respectively, to reach peak performance.
+These results validate the efficacy of our latent clustering-based selection approach in enhancing semi-supervised adversarial training.
+The proposed method not only maintains the robust accuracy of the final model, $\theta_{\mathrm{final}}$, but also effectively accelerates the training process, underscoring its practical utility and scalability in SSAT frameworks.
+V-B Computational Benefit of Data Reduction
+As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can effectively reduce the overall training time of SSAT while maintaining competitive robustness.
+We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
+Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages.
+In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing.
+When no additional unlabeled data is utilized, the highest test robust accuracy is achieved after approximately $75$ epochs (Figure 1a).
+Incorporating $10\%$ external unlabeled data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c).
+In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d).
+These observations suggest that including extra unlabeled data in SSAT increases computational demands.
+Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
+TABLE II: Performance of full SSAT ($\alpha=100\%$) with early stopping at epoch $100$ on CIFAR-10 under $\ell_{\infty}$ perturbations with $\epsilon=0.031$.
+Data Source Schedule Clean ($\%)$ PGD ($\%)$ AA ($\%)$
+External (500K) Cosine $84.1$ $54.0$ $50.4$
+Stepwise $88.0$ $60.6$ $57.6$
+DDPM (1M) Cosine $78.6$ $52.1$ $48.6$
+Stepwise $83.0$ $58.5$ $55.9$
+TABLE III: Runtime comparison of SSAT with synthetically generated unlabeled data across various settings on CIFAR-10.
+$\bm{\alpha}$ Method Fine-Tuning (h) Generation (h) Selection (h) #Epochs Total Time (h) PGD (%)
 $0\%$ No Gen. $100$ $10.7$ $55.4$
@@
 $20\%$ LCG-KM $0.43$ $0.77$ $100$ $15.7$ $60.2$
-TABLE III: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
-Ratio $\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
+TABLE IV: Performance of our DDPM fine-tuning strategies on CIFAR-10 under varying $\alpha$. The best robust accuracies are underlined.
+$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) AA ($\%$)
 $0\%$ No Gen. $84.9$ $55.4$ $49.2$
@@
 LCG-KM $85.7$ $60.3$ $57.4$
-IV-B Computational Benefits of Data Reduction
-As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall training time of SSAT while maintaining competitive robustness.
-We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
-Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages.
-In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing.
-When no additional data is utilized, the highest robust test accuracy is achieved after approximately $75$ epochs (Figure 1a).
-Incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c).
-In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d).
-These findings suggest that including extra unlabeled data in SSAT increases computational demands.
-Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
-Generally speaking, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy.
+Generally, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy.
 Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training.
-However, continued training beyond this point often leads to robust overfitting.
-Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations.
-Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance.
-To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time.
-These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
+Early stopping offers a computational advantage in scenarios with limited data by preventing overfitting with a reduced number of training epochs.
+Nevertheless, early stopping proves less effective in full SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance.
+Table II summarizes the model performance on CIFAR-10 with respect to full SSAT when it is early stopped at the same training epoch as our methods.
+We tested two learning rate schedulers: cosine annealing and stepwise decay.
+The results show that the training process of full SSAT algorithms has not converged at the early-stopped training epoch.
+Importantly, even when comparing at the same number of training epochs, our LCS-KM method with $20\%$ DDPM-generated data achieves $60.2\%$ PGD robust accuracy at epoch $100$, outperforming early-stopped full SSAT by $1.7\%$.
+This demonstrates that our performance gain stems not only from reduced epochs but also from accelerated convergence enabled by strategically selecting more informative samples.
+Due to space limits, full results and detailed discussions are provided in Appendix D-C.
+These findings highlight the improved trade-offs of our methods in terms of computational costs, data efficiency, and robustness, offering a better solution to address the challenges of SSAT with extensive datasets.
 (a) Random Selection ($\alpha=10\%$)
@@
 Figure 2: Standard and robust accuracy curves of SSAT with labeled data from COVIDGRand unlabeled data from CoronaHack, with different selection schemes and ratios: (a) random selection with $\alpha=10\%$, (b) LCS-KM with $\alpha=10\%$, (c) LCS-KM with $\alpha=20\%$, and (d) all unlabeled data $\alpha=100\%$.
-IV-C Guided DDPM Fine-Tuning
-We study the performance of our guided DDPM fine-tuning methods as depicted in Algorithm 3 on CIFAR-10.
-Unlike our previous selection schemes, which required pre-generating the whole dataset, here we generate the same amount of unlabeled data matching the ratio $\alpha$.
+V-C Guided DDPM Fine-Tuning
+We evaluate the performance of our guided DDPM fine-tuning methods on CIFAR-10.
+Unlike our selection-based schemes, which require pre-generating the entire dataset, here we generate the same amount of unlabeled data, matching the ratio $\alpha$.
 The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection.
 We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments.
-Table III presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart.
+Table IV presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart.
 For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM.
 Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives.
-These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set can yield performance comparable to a much larger dataset.
+These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set of samples can yield performance comparable to a much larger dataset.
 Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup.
@@
 We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively.
-The comparison results are summarized in Table II.
+The comparison results are summarized in Table III.
 For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process.
@@
 Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space.
-Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the ten class clusters, and (c) LCS-KM highlights points selected near decision boundaries across ten classes based on k-means clustering in the latent space.
-(a) CIFAR-10
-(b) SVHN
-(c) CIFAR-10
-(d) SVHN
-Figure 4: Illustration of SSAT performance using our LCS-KM method on image benchmarks with varying boundary and unlabeled data reduction ratio parameters with respect to: (a) and (c) DDPM-generated CIFAR-10 data with $\alpha=20\%$, (b) and (d) 531531K extra SVHN data with $\alpha=10\%$.
-For each figure, we vary the considered ratio parameter while keeping all the remaining hyperparameters fixed.
-V Additional Analyses
-In this section, we present additional analyses to provide further insights into our methods.
-We first evaluate their performance on a real-world medical application dataset (Section V-A), and then visualize the selected data distribution in the latent space to qualitatively understand their ability in identifying critical data points (Section V-B).
-Finally, we provide ablation studies on key hyperparameters of our methods.
-Due to page limit, additional experiments on other hyperparameters, generalizability under different training schemes and perturbations, and the effect of intermediate model quality in low-label regimes, are provided in Appendix D.
-V-A Application to Real-World Medical Data
-To test the generalization of our findings, we extend our study to a real-world healthcare application.
+Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the $10$ class clusters, and (c) LCS-KM highlights points selected near decision boundaries across $10$ classes based on k-means clustering in the latent space.
+VI Further Analyses
+This section presents further analyses, including evaluations on a real-world medical dataset (Section VI-A), visualizations of selected unlabeled data (Section VI-B), and ablation studies on key hyperparameters (Section VI-C).
+Due to the page limit, additional experiments under different training schemes and perturbations, as well as the effect of the intermediate model in low-label regimes, are provided in Appendix D.
+VI-A Application to Real-World Medical Data
+We extend our study to a real-world healthcare application.
 COVIDGR [41] is used as the labeled dataset, and CoronaHack [16] serves as the source of the unlabeled data.
@@
 CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set.
-We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios.
-Further details of the experimental setup is discussed in Appendix B.
-The results are demonstrated in Figure 2, which confirms the effectiveness of our selection scheme in enhancing robustness while optimizing efficiency.
+We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios (see Appendix 3 for more details).
+Figure 2 demonstrates the results, which again confirm the effectiveness of our selection scheme.
 For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b).
@@
 Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$.
-In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve the peak robust accuracy of $58\%$ (Figure 2d).
-These findings confirm that using a smaller but strategically selected unlabeled data subset can achieve robustness comparable to that of the whole dataset while significantly reducing computational costs.
-These results highlight the practical relevance of our data reduction scheme, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
-V-B Comparative Visualizations of Our Methods
+In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve a peak robust accuracy of $58\%$ (Figure 2d).
+These findings confirm that using a smaller, but strategically selected, unlabeled data subset can achieve robustness comparable to that of the entire dataset.
+These results highlight the practical relevance of our data reduction scheme, particularly for applications where labeled data is limited yet robust model performance is essential.
+(a) CIFAR-10 ($\alpha=20\%$)
+(b) SVHN ($\alpha=10\%$)
+(c) Fine-Tuning Epochs $S$
+(d) Regularization Strength $\lambda$
+Figure 4: Figures (a) and (b): Illustration of SSAT performance using LCS-KM by varying boundary ratio parameters $\beta$ with respect to DDPM-generated CIFAR-10 data with $\alpha=20\%$ and 531531K extra SVHN data with $\alpha=10\%$.
+Figures (c) and (d): Clean and robust accuracy curves of SSAT with LCG-KM on CIFAR-10 with $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
+VI-B Comparative Visualization
 To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10.
-We first extract their latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model.
+We first extract the latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model.
 For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships.
-Specifically, we first perform Principal Component Analysis (PCA) [32].
+In particular, we first perform Principal Component Analysis (PCA) [14].
 By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data.
-In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [33] to project the $50$-dimensional PCA-reduced features onto a two-dimensional space.
+Then, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [46] to project the $50$-dimensional PCA features onto a $2$-dimensional space.
 Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points.
-After projecting onto the two-dimensional space, we visualize the set of unlabeled data returned by our selection schemes under $\alpha=10\%$.
-Figure 3 presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy.
-Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT.
-In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes.
-Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers that offer limited benefit for refining the decision boundary.
+Finally, we visualize the set of unlabeled data selected by our methods under $\alpha=10\%$.
+Figure 3 shows the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point.
+The highlighted black dots indicate the data points selected by the respective selection strategy.
+Specifically, PCS adopts a purely confidence-based selection strategy, choosing the points with the globally lowest confidence without considering their distribution within or across classes.
+Although this approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers.
 As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a).
@@
 As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b).
-A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical examples for training.
-In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM.
-Finally, our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner.
+A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical samples for SSAT.
+Our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner.
 By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c).
@@
 This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer.
-We hypothesize that such a local linear property potentially explains why our latent clustering-based approach with k-means always achieves the best performance compared to the other two alternatives, which is worth further investigation.
-V-C Hyperparameter Sensitivity
-The success of our methods relies on a few key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$.
-We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally.
-Due to page limits, the detailed settings of these experiments and the ablations on $\gamma$ are deferred to Appendix C-A.
+We hypothesize that this local linear property may explain why our latent clustering-based approach with k-means consistently achieves the best performance compared to the other two alternatives.
+VI-C Hyperparameter Sensitivity
+We conduct ablations to study the sensitivity of the hyperparameters involved in our methods The detailed setup and the ablations on the unlabeled data ratio $\alpha$ and the per-batch ratio of extra to original data $\gamma$ are deferred to Appendix C-A.
 Ratio of Boundary Data $\bm{\beta}$.
@@
 Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data.
-When more external data is added, using a smaller $\beta$ usually leads to better robust accuracy.
-Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.
-Data Reduction Ratio $\bm{\alpha}$.
-Figures 4c and 4d describe how accuracies vary with the ratio parameter $\alpha$.
-We analyze how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary.
-When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy.
-However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability.
-Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
-(a) Fine-Tuning Epochs $S$
-(b) Regularization Strength $\lambda$
-Figure 5: Clean and robust accuracy curves of SSAT with LCG-KM using generated data on CIFAR-10 under $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
-Guided DDPM Finetuning.
-We conduct ablation studies on our guided DDPM fine-tuning approach to better illustrate its behavior.
-In particular, we focus on two key hyperparameters, the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$, and study their impact on accuracy and robustness performance when training SSAT algorithms on CIFAR-10 with LCG-KM and $\alpha=10\%$.
-We also investigate the quality of the data samples generated by our fine-tuned DDPM using both quantitative and qualitative metrics.
-Figure 5a illustrates our ablation results on $S$.
+When more external data is added, using a smaller $\beta$ typically yields better robust accuracy.
+Conversely, choosing a larger $\beta$ is favorable when there is less extra unlabeled data.
+Guided DDPM Fine-Tuning ($S$ & $\lambda$).
+We conduct ablations on our best-performing LCG-KM approach by varying the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$ on CIFAR-10 under $\alpha=10\%$.
+Figure 4c illustrates our ablation results on $S$.
 We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process.
-Figure 6 visualizes the images generated by our DDPM models with different fine-tuning epochs.
-We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained for CIFAR-10.
-Training for too many epochs can lead to a degradation in image quality and loss of diversity, thereby adversely influencing the final robust accuracies.
-In addition, Figure 5b shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$.
-Our findings show that choosing $\lambda$ within the range of $[0.4,0.6]$ results in the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
-Figure 6: Visualization of generated data across different fine-tuning epochs
-VI Conclusion and Future Work
-In this work, we proposed various unlabeled data reduction schemes to improve the efficiency of SSAT while preserving robust accuracy.
-In particular, we emphasized that strategically selecting or generating more boundary-adjacent data points can significantly improve the efficiency of semi-supervised adversarial training algorithms, without losing their robustness advantages.
-Among our proposed methods, we found that latent-based approaches using k-means clustering, including LCS-KM and LCG-KM, perform the best, as they reduce computational costs and memory requirements.
-Limitation & Future Work.
-We acknowledge that our work has several limitations.
-While the selection scheme with LCS-KM consistently achieves the best performance, its success requires carefully choosing a few hyperparameters (e.g., to avoid overfitting).
-Future studies can leverage AutoML algorithms to automatically determine these hyperparameters.
-Additionally, it remains unclear whether our best-performing latent-clustering-based methods (i.e., LCS-KM and LCG-KM) can be further optimized to achieve better performance.
-Theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful.
-Furthermore, our guided DDPM fine-tuning approach proposes to adapt a pre-trained model for boundary-seeking data generation.
-An interesting future direction is to investigate whether training-free sampling-based guidance [34] or more advanced generative models [35, 36] can enhance the efficiency and robustness of SSAT.
-Despite these limitations, we believe our work can serve as an important initial step towards the more efficient and scalable development of robust models.
-Availability
-The implementation of the latent-based data reduction methods and all our experiments are available as open-source code at: https://github.com/TrustMLRG/EfficientSSAT.
-Appendix A Related Work
-A-A Adversarially Robsut Learning
-Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [1].
-The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications.
-Earlier-proposed attacks like fast gradient sign method (FGSM) [2] and projected gradient descent (PGD) [7] introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks [37], which search for the minimal perturbation that causes misclassification.
-Recent attacks like Auto-PGD [38] automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack [17] is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
-Adversarial Training.
-On the defense side, adversarial training and its variants are the most popular methods for achieving robustness.
-For instance, Madry et al.
-proposed to train DNNs on adversarially perturbed inputs using PGD attacks [7], while Zhang et al.
-introduced TRADES [9], which balances the trade-off between robustness and standard accuracy.
-Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications.
-In particular, Rice et al.
-illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training [28], where early stopping is shown to be effective in alleviating such an issue.
-Techniques, such as Free AT [18] and Fast AT [19], offer promising solutions to reduce the computation for adversarial training.
-Nevertheless, they suffer from decreased robustness performance.
-Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently.
-For example, Zhang et al.
-argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [16].
-As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data.
-In a similar vein, Hua et al.
-advocated for a more targeted approach [24], recommending that PGD training be applied primarily to examples situated near the decision boundaries.
-This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks.
-Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
-Self-Supervised Adversarial Training.
-Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research [39].
-In the adversarial context, Schmidt et al.
-[10] highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning.
-They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly.
-Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training [11, 40, 41, 12].
-In particular, Carmon et al.
-proposed robust self-training methods to address the sample complexity issue [11], which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data.
-The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model.
-Concurrently, Najafi et al.
-proposed an optimization framework incorporating labeled and unlabeled data [41], while providing formal guarantees on adversarial robustness.
-Leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature [13, 14, 15].
-For instance, Gowal et al.
-employed various unconditional generative models [13], including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further.
-While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement.
-In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
-A-B Data-Efficient Deep Learning
-Within the broader literature of data-efficient deep learning, various techniques have been proposed to reduce the amount of training data while retaining critical information.
-Two prominent approaches are dataset distillation and coreset selection.
-Dataset Distillation.
-Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points.
-For instance, Wang et al.
-proposed a foundational approach to dataset distillation [42], proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance.
-Later, Zhao et al.
-introduced the concept of gradient matching [43], where distilled data points are optimized to match the gradients of the original dataset.
-This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data.
-Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [44], facilitating the training of complex models with minimal data while maintaining strong performance.
-Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands.
-Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
-Coreset Selection.
-On the other hand, coreset selection [45, 46] aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset.
-These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning.
-For example, Kaushal et al.
-utilized diverse models to select training data subsets to reduce labeling efforts [20], while Xia et al.
-introduced the concept of “moderate coreset” [21], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions.
-Similarly, Mirzasoleiman et al.
-proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [22].
-In the adversarial context, Dolatabadi et al.
-proposed adversarial coreset selection [47], offering a task-specific solution to enhance the computational efficiency of adversarial training.
-In particular, the method constructs a compact data subset by minimizing the gradient approximation error.
-The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals.
-While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges.
-These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time.
-In such settings, repeatedly selecting data during training may become computationally infeasible.
-In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
+We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained on CIFAR-10.
+Training for too many epochs can lead to a degradation in image quality and loss of diversity (see Figure 5 in the appendix for visualizations), thereby adversely influencing the final robust accuracies.
+Figure 4d shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$.
+Our findings indicate that selecting $\lambda$ within the range of $[0.4,0.6]$ yields the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
+VII Conclusion and Future Work
+In this work, we developed various data reduction schemes to improve the efficiency of SSAT.
+We illustrated that strategically selecting or generating more boundary-adjacent data points can significantly improve the data efficiency of SSAT while preserving its robustness advantages.
+In particular, our latent-based k-means clustering approach performs the best, reducing both computational costs and memory requirements.
+Limitation & Future Direction.
+While LCS-KM consistently achieves the best performance, its success requires careful hyperparameter tuning; future work can leverage AutoML algorithms to automate this process.
+In addition, theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful.
+Finally, extending our guided DDPM fine-tuning with training-free sampling-based guidance [11] or advanced generative models [38, 13] is a promising direction.
+Further discussions are provided in Appendix D-D.
+Despite these challenges, we believe our work can serve as an important initial step towards more efficient and scalable development of robust models.
+Appendix A Algorithm Pseudocode
+Algorithms 1-3 present the pseudocode of our data reduction schemes, corresponding to prediction confidence-based selection (PCS), latent clustering-based selection (LCS), and guided DDPM-based fine-tuning, as introduced in Section IV.
 Appendix B Detailed Experimental Setup
@@
 Initially, the model is trained on the 7373K labeled images.
-To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model [13].
+To evaluate the generalizability of our selection schemes, we apply them to a synthetic dataset generated using a DDPM model [17].
 Configuration.
@@
 For SVHN, we use a training batch size of $128$.
-All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with 11M generated data are run for $400$ epochs.
+All of our experiments that utilize all of the extra 531K data are run for $200$ epochs, and the experiments with 11M generated data are run for $400$ epochs.
 To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping.
-Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
+Following TRADES [50], we employ an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$, along with a stepwise learning rate scheduler with a decaying factor of $10$ at epochs $75$ and $90$.
+Algorithm 1 Prediction Confidence-Based Selection (PCS)
+1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
+2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
+3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
+4:  Initialize $\mathcal{A}_{u}$ as empty set
+5:  $\mathrm{Conf}(\bm{x})\leftarrow$ get confidence of $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
+6:  Sort $\mathcal{S}_{u}$ by ascending order of $\mathrm{Conf}(\bm{x})$
+7:  $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with lowest $\mathrm{Conf}(\bm{x})$
+8:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
+9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
+10:  Output: selected subset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
+Algorithm 2 Latent Clustering-Based Selection (LCS)
+1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
+2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
+3:  $\hat{y}\leftarrow$ predict pseudo label $f_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
+4:  Initialize $\mathcal{A}_{u}$ as empty set
+5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
+6:  if using LCS-KM then
+7:   $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{k}\}\leftarrow$ k-means clustering on $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
+8:   $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in\mathcal{S}_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
+9:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta d$
+10:  else if using LCS-GMM then
+11:   Fit a GMM with $k$ components to $\{\bm{z}\}_{\bm{x}\in\mathcal{S}_{u}}$
+12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
+13:   $\Delta p\leftarrow$ compute $\Delta p=|p_{1}-p_{2}|$ for each $\bm{z}$, where $p_{1}$ and $p_{2}$ are the top two highest probabilities
+14:   $\mathcal{A}_{u}\leftarrow$ add top $\beta\cdot\alpha|\mathcal{S}_{u}|$ points with smallest $\Delta\gamma$
+15:  end if
+16:  $\mathcal{A}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points randomly from $\mathcal{S}_{u}\setminus\mathcal{A}_{u}$
+17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
+18:  Output: selected dataset $\mathcal{A}_{u}$, final model ${\theta}_{\mathrm{final}}$
+Algorithm 3 Guided DDPM Fine-Tuning
+1:  Input: labeled $\mathcal{S}_{l}$; pre-trained DDPM $\theta_{\mathrm{pre}}$; fine-tuning epochs $S$; trade-off parameter $\lambda$; other parameters $\epsilon$, $\alpha|\mathcal{S}_{u}|$, $\beta$, $\gamma$, $k$, $\eta$
+2:  $\hat{\theta}\leftarrow$ train a standard classification model on $\mathcal{S}_{l}$
+3:  Initialize $\mathcal{G}_{u}$ as empty set
+4:  $\mathcal{G}_{u}\leftarrow$ add $(1-\beta)\cdot\alpha|\mathcal{S}_{u}|$ points generated by pre-trained $\theta_{\mathrm{pre}}$
+5:  Initialize $\theta$ as $\theta_{\mathrm{pre}}$
+6:  for $s=1,2,\ldots,S$ do
+7:   Sample batches of size $b$: $\bm{x}_{0}\sim\mathcal{S}_{l}$, $t\sim\mathcal{U}(T)$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\mathbf{I})$
+8:   $\bm{x}_{t}\leftarrow\sqrt{\bar{\alpha}_{t_{i}}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$
+9:   $L_{\mathrm{DDPM}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\|_{2}^{2}$
+10:   if using PCG then
+11:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{PC}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+12:   else if using LCG-KM then
+13:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{KM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+14:   else if using LCG-GMM then
+15:    $L_{\mathrm{reg}}\leftarrow\frac{1}{b}\sum_{i\in[b]}\|\ell_{\mathrm{GMM}}(g_{\theta}(\bm{x}_{t},t))\|_{2}^{2}$
+16:   end if
+17:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\big(L_{\mathrm{DDPM}}+\lambda\cdot L_{\mathrm{reg}}\big)$
+18:  end for
+19:  $\mathcal{G}_{u}\leftarrow$ add $\beta\cdot\alpha|\mathcal{S}_{u}|$ points generated by fine-tuned $\theta$
+20:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)$
+21:  Output: generated dataset $\mathcal{G}_{u}$, final model $\theta_{\mathrm{final}}$
 Evaluation.
-For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al.
+For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to those of Carmon et al.
 [5] for better comparison.
-We use step size $0.005$, number of attack steps $K=100$.
-We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training.
-We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
+We use a step size $0.005$, and a number of attack steps of $K=100$.
+We evaluate models at $\epsilon=0.015$, which is the same value used during training.
+We also use Auto Attack with the same $\epsilon$ value as in the PGD attack.
 CIFAR-10.
@@
 Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments.
-We train the intermediate model with 5050K labeled data and use this model to select data from the 500500K pseudo-labeled data or a synthetic dataset.
+We train the intermediate model using 5050K labeled data and utilize this model to select data from either the 500500K pseudo-labeled data or a synthetic dataset.
 We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset.
-We also conducted further experiments utilizing 11M synthetical data generated by the DDPM model [13].
+We also conducted further experiments using 11M synthetic data generated by the DDPM model [17].
 For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [20] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
@@
 For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models.
-Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$.
+Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$.
 For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$.
-In section D-A when $\epsilon=4/255$, we use $0.0035$ as the PGD step size.
+In section D-A, when $\epsilon=4/255$, we use $0.0035$ as the PGD step size.
 For $\epsilon=16/255$, we use $0.015$ as the PGD step size.
@@
 The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs.
-For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process.
-Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
+For our experiments with limited data, we get the best results at $100$ epochs, where we early stop the training process.
+We employ an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$, along with a stepwise scheduler with a decaying factor of $10$ at epochs $75$ and $90$, respectively.
 Evaluation.
@@
 [5] for fair comparisons.
-We use step size $\alpha=0.01$, number of attack steps $K=40$.
+We use a step size $\alpha=0.01$, and a number of attack steps $K=40$.
 We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training.
 For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$.
-We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
+We also use Auto Attack with the same $\epsilon$ value as in the PGD attack.
 Medical Application.
-Below, we detail the setup employed for our experiments on the medical datasets outlined in Section V-A.
-For this experiment, we trained models using a ResNet-18 [50] architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$.
+For experiments on the medical dataset outlined in Section VI-A, we train models using a ResNet-18 architecture [19], and consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$.
 We set the attack step size to $0.02$, and the number of PGD steps to $10$.
-The final model is trained with SSAT using the Adam optimizer [51] with a learning rate of $0.001$.
+The final model is trained using SSAT with the Adam optimizer [25] and a learning rate of $0.001$.
 We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
+Figure 5: Visualization of generated images with various fine-tuning epochs $S$.
 Appendix C Additional Experiments
@@
 (b) SVHN
-Figure 7: Illustration of clean and robust accuracy curves of SSAT algorithms with LCS-KM by varying per-batch ratio $\gamma$, which defines how much labeled and unlabeled data are included in each mini-batch, on: (a) CIFAR-10 with $20\%$ extra unlabeled data selected from 11M generated images using DDPM, and (b) SVHN with $10\%$ selected from the external 531531K SVHN images.
-TABLE IV: Comparison results of semi-supervised adversarial training algorithms with varying data selection schemes on SVHN.
-Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
-$\textbf{Ratio }\bm{\alpha}$ Method Clean Accuracy ($\%$) Robust Accuracy ($\%$) Number of Epochs
-$1\%$ LCS-KM $95.6$ $82.9$ $75$
-$10\%$ Random $96.3$ $82.7$ $75$
-$10\%$ LCS-KM $96.7$ $86.3$ $75$
-$20\%$ LCS-KM $96.4$ $86.5$ $75$
-$100\%$ No Selection $97.5$ $86.4$ $200$
-C-A Hyperparameter Sensitivity Analysis Details
-In Section V-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values.
-In these experiments, we varied one hyperparameter at a time while keeping the others constant.
+(c) CIFAR-10
+(d) SVHN
+Figure 6: Figures (a) and (b): Illustration of SSAT performance using LCS-KM by varying unlabeled data ratios $\alpha$ for CIFAR-10 with DDPM generation ($\beta=0.4$) and SVHN with external data ($\beta=0.6$).
+Figures (c) and (d): Clean and robust accuracy curves of SSAT with LCS-KM by varying per-batch ratio $\gamma$ on CIFAR-10 with 11M generated images using DDPM ($\alpha=20\%$, $\beta=0.4$) and on SVHN with external 531531K images ($\alpha=10\%$, $\beta=0.6$).
+C-A Other Ablations on Hyperparameters
+In Section VI-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values.
+We vary one hyperparameter at a time while keeping the others constant.
 Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$.
 Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
-In Figure 7a, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$.
-Similarly, Figure 7b displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$.
-In addition, Figures 4c and 4d explore how robust accuracy varies with the amount of data used.
-For CIFAR-10 experiments in Figure 4c, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$.
-For SVHN in Figure 4d, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$.
-We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures 4c and 4d, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
+For our LCG ablations on $S$ and $\lambda$ (Figures 4c and 4d), we set $\alpha=20\%$, $\beta=0.4$ and $\gamma=0.3$ for CIFAR-10, and choose $\alpha=10\%$, $\beta=0.6$ and $\gamma=0.5$ for SVHN, matching the same set of optimal hyperparameters as identified for our selection-based strategies.
+Figure 5 visualizes the images generated by our guided DDPM fine-tuning techniques, including PCG, LCG-GMM, and LCG-KM, on CIFAR-10 with different numbers of fine-tuning epochs $S\in\{0,5,15,25,35,45\}$.
+As $S$ is increased, the generated CIFAR-10 images become more difficult to recognize, reflecting a higher degree of uncertainty.
+Figures 6a and 6b demonstrate how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary.
+For CIFAR-10, the per-batch ratio is set to $0.3$, and $\beta$ is fixed at $0.4$.
+For SVHN, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$.
+We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that as the amount of included unlabeled data increases, the optimal accuracy is reached at a considerably later epoch.
+When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy.
+However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained in the model’s generalization ability.
+Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
+In Figure 6c, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$.
+Similarly, Figure 6d displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$.
 (a) CIFAR-10
@@
 $f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset.
-However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the data selection process’s intended benefits, primarily to reduce computational and time complexity.
-In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
-To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings [52].
+However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the intended benefits of the data selection process, primarily to reduce computational and time complexity.
+In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
+To address these issues, we investigate the feasibility of utilizing an alternative pre-trained model in LCS for data selection, specifically by leveraging the CLIP model to extract image embeddings [32].
 CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training.
 By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency.
-This substitution allows us to select a relevant subset without having to train an intermediate model from scratch.
+This substitution enables us to select a relevant subset without requiring the training of an intermediate model from scratch.
 Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies.
-Figure 8 shows our result when using CLIP instead of the previous intermediate model for data selection.
-C-C SVHN with Ground-Truth Labels
-We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data.
-Table IV shows the comparison results.
-With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
-C-D Resnet-18 Model Architecture
-Our main experiments focus on the WideResNet architecture.
-We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model.
-Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet.
-Our evaluation results are shown in Table V, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
-TABLE V: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations.
+Figure 7 shows the result when using CLIP instead of the previous intermediate model for unlabeled data selection in SSAT.
+TABLE V: Comparison results of SSAT algorithms with varying data selection schemes on SVHN.
+Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
+$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) #Epochs
+$1\%$ LCS-KM $95.6$ $82.9$ $75$
+$10\%$ Random $96.3$ $82.7$ $75$
+$10\%$ LCS-KM $96.7$ $86.3$ $75$
+$20\%$ LCS-KM $96.4$ $86.5$ $75$
+$100\%$ No Selection $97.5$ $86.4$ $200$
+TABLE VI: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations.
 The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
-$\textbf{Ratio }\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
+$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
 $1\%$ Random $79.8$ $50.6$
@@
 $100\%$ No Selection $83.2$ $54.7$
-Appendix D Extended Analyses
-In this section, we further evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section D-A), and examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section D-B).
-Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table I.
 (a) TRADES ($\epsilon=4/255$, PGD-10)
@@
 Figure 8: Illustration of SSAT performance on CIFAR-10 using LCS-KM under $\ell_{\infty}$ perturbations under various configurations: TRADES using $10$ PGD steps with (a) $\epsilon=4/255$ and (b) $\epsilon=16/255$; AT with $\epsilon=8/255$ using (c) 55 PGD steps and (d) $10$ PGD steps.
-TABLE VI: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data.
+C-C SVHN with Ground-Truth Labels
+We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data.
+Table V shows the comparison results.
+With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
+C-D ResNet-18 Model Architecture
+Our main experiments focus on the WideResNet architecture.
+We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model.
+Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet.
+Our evaluation results are shown in Table VI, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
+Appendix D Extended Analyses
+In this section, we further evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section D-A), and examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section D-B).
+Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table I.
+TABLE VII: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data.
 We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
 Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
-Training Method Clean Acc ($\%$) Clean Acc ($\%$) Robust Acc ($\%$)
+Training Method Clean Accuracy ($\%$) Clean Accuracy ($\%$) PGD Robust Accuracy ($\%$)
 $1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
@@
 We consider external data from TinyImages as the unlabeled dataset.
-We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table I.
-The results are depicted in Figures 9a and 9b, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data.
-The trends are shown to be consistent across varying perturbation sizes.
+We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm, similar to the experiments in Table I.
+The results are depicted in Figures 8a and 8b, suggesting that our LCS-KM method is highly effective compared to random selection and no selection, approaching the robust accuracies achieved by fully utilizing unlabeled data.
+The trends are consistent across varying perturbation sizes.
 Adversarial Training.
@@
 (b) Generated data
-Figure 10: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$: (a) external TinyImages and (b) DDPM-generated data.
+Figure 9: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$ with: (a) 500500K external TinyImages and (b) 11M DDPM-generated data.
 $\bm{\ell}_{2}$ Perturbations.
@@
 We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data.
-Our findings are shown in 10a and 10b, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$.
+Our findings are presented in Figures 9a and 9b, which reveal that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$.
 These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
 D-B Low Labeled Data Regime
-We further study how the choice of the intermediate model and the amount of labeled data impact the final performance.
+We further investigate how the choice of intermediate model and the amount of labeled data affect the final performance.
 For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet.
@@
 The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied.
-The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low.
+The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, provided the labeled data ratio is low.
 In addition, we include a pre-trained CLIP model [32] as an alternative intermediate model for comparison.
-The results of these experiments are presented in Table VI.
-The intermediate model’s quality significantly impacts the final model’s robust accuracy.
+Table VII summarizes our results.
+The quality of the intermediate model significantly impacts the robust accuracy of the final model.
 In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model.
-However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model.
+However, as the amount of labeled data for the intermediate model increases (e.g., $20\%$ or $100\%$), the performance of the final model becomes similar to or better than that of its counterparts based on the CLIP intermediate model.
 This trend demonstrates the effect of the intermediate model on the final robust accuracy.
-These observations underscore the importance of having a strong intermediate model for robust performance.
-Besides, we observe that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-labeled-data regimes.
-Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
-D-C Further Discussion
-Our work introduces several data reduction strategies that tackle the key challenges in adversarially robust learning, particularly the significant memory and computational overhead of state-of-the-art SSAT methods.
-By prioritizing boundary-adjacent data, these processes aim to enhance model robustness.
-However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models.
-While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset.
-By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization.
-Nevertheless, a deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization.
-Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
-Reducing the overall time required for adversarial training is critical for improving the efficiency and scalability of robust learning.
-Broadly, there are two complementary strategies to achieve this: (i) decreasing the computational cost of each training epoch, and (ii) reducing the total number of epochs needed for convergence.
-The first strategy aims to make each epoch faster without sacrificing robustness.
-Several prior works [18, 19, 53, 54] focus on this direction by introducing more efficient attack generation and optimization schemes.
-These approaches demonstrate that careful algorithmic design can substantially reduce per-epoch runtime, making adversarial training far more practical in large-scale settings.
-However, as shown in previous studies [15, 13], the total number of training epochs grows sharply with the size of the dataset.
-Consequently, even if the cost per epoch is reduced, increasing the amount of training data can still lead to a dramatic rise in overall training time.
-For example, adversarial training with approximately 11M generated samples typically requires around $400$ epochs to converge, whereas training with 100100M samples may demand as many as $2000$ epochs.
-Developing strategies that mitigate this dependence so that large-scale training does not necessitate a proportional increase in epochs would provide substantial computational savings and further improve the practicality of adversarially robust learning.
+These observations highlight the importance of having a robust intermediate model for achieving strong performance.
+Additionally, we observe that pre-trained models, such as CLIP, can be a good alternative to training an intermediate model, especially in low-labeled-data regimes.
+Detailed results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
+TABLE VIII: Full version of Table IV.
+We evaluate full SSAT ($\alpha=100\%$) on CIFAR-10 under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ under various learning rate schedulers with/without early stopping.
+Setting Schedule Epoch Clean ($\%)$ PGD ($\%)$ AA ($\%)$
+External Cosine $100$ $84.1$ $54.0$ $50.4$
+$200$ $89.3$ $62.2$ $59.8$
+Stepwise $100$ $88.0$ $60.6$ $57.6$
+$200$ $89.7$ $62.5$ $58.6$
+DDPM Cosine $100$ $78.6$ $52.1$ $48.6$
+$400$ $86.2$ $62.4$ $59.9$
+Stepwise $100$ $83.0$ $58.5$ $55.9$
+$400$ $85.7$ $61.8$ $58.4$
+D-C Performance of Full SSAT Stopped at Earlier Epochs
+To investigate whether early stopping can serve as an effective strategy for full SSAT rather than training for the complete number of epochs, we evaluate SSAT models on CIFAR-10 using $100\%$ of the unlabeled data with two popular learning rate scheduling schemes, both with an initial learning rate of $0.1$: (Cosine) cosine annealing and (Stepwise) stepwise decay with a decaying factor of $10$ at training epochs $75$ and $90$.
+We compare the performance of SSAT stopped at epoch $100$, the same as our data reduction methods, against SSAT models trained for the full duration: $200$ epochs for 500500K external TinyImages and $400$ epochs for 11M DDPM-generated data.
+Table VIII presents the complete results.
+Below, we summarize the key observations regarding the interplay between learning rate scheduling and the training duration of SSAT:
+Cosine vs.
+Stepwise.
+The two learning rate schedulers exhibit distinct training dynamics during the early epochs.
+At epoch $100$, models trained with cosine annealing show significantly lower performance (e.g., $50.4\%$ AA robust accuracy for external data and $48.6\%$ for DDPM-generated data) compared to stepwise decay ($57.6\%$ and $55.9\%$, respectively).
+However, when trained to completion, cosine annealing achieves more competitive final robustness performance than stepwise scheduling, approximately $1\%-1.5\%$ improvement in robust accuracy when evaluated using AutoAttack.
+Effectiveness of Early Stopping.
+For the stepwise scheduler, early stopping at epoch $100$ can achieve robust accuracy not far from the robustness of the final model returned by full SSAT.
+For external data, the robustness gap between epochs $100$ and $200$ is only $1\%$ with a doubled training time.
+Similarly, for DDPM data, early stopping at epoch $100$ yields $55.9\%$ AA robust accuracy compared to $58.4\%$ at epoch $400$, a difference of $2.5\%$ with $4\times$ more training.
+In comparison, early stopping is not effective for full SSAT when using the cosine annealing scheduler.
+Prolonged training of full SSAT algorithms from epoch $100$ gradually increases the clean and robust accuracy performance until reaching peak performance, regardless of the type of learning rate scheduler implemented.
+These observations are aligned with the full SSAT curves illustrated in Figure 1d, suggesting that learning is not complete for SSAT with $\alpha=100\%$ if it is simply stopped at an early epoch.
+Comparison with LCS-KM.
+Even with the stepwise scheduler and early stopping employed, full SSAT early stopped at epoch $100$ achieves AA robust accuracy of $57.6\%$ (external) and $55.9\%$ (DDPM-generated), slightly lower than our LCS-KM method with $20\%$ unlabeled data that achieves $57.8\%$ and $57.2\%$ for the respective settings (Table I).
+Under cosine scheduling, the robustness improvement of using LCS-KM is much more pronounced compared to full SSAT with early stopping.
+These results demonstrate the advantages of our data reduction methods, offering a viable solution to improve the data efficiency of SSAT algorithms and accelerate the learning process while largely preserving their robustness advantages.
+D-D Further Discussions
+Our work introduced several strategies to improve the data efficiency of SSAT algorithms by prioritizing unlabeled data adjacent to the model’s decision boundaries.
+However, the success of our methods hinges on the hypothesis that the intermediate model’s latent representations are reliable for pseudo-labeling and prioritizing unlabeled data, thereby improving robust generalization, which may falter in scenarios where the model is undertrained or poorly calibrated.
+In addition, while our methods effectively reduce the volume of unlabeled data, their final performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points.
+Developing automatic tools to automate the process of hyperparameter selection, as well as conducting in-depth theoretical studies to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization, along with their adaptability to potential variations, would be meaningful future directions.
+Moreover, reducing the overall runtime required for adversarial training without sacrificing robustness is critical for achieving efficient and scalable robust learning.
+In the prior literature, two complementary strategies have been proposed: (i) decreasing the computation of each training epoch, such as free and fast adversarial training [37, 45], and other more efficient optimizers for generating perturbations in an adversarial training framework [7, 53], and (ii) reducing the total number of epochs needed for convergence, such as implementing early stopping [33].
+While the former techniques can reduce the per-epoch runtime, the overall training process is less stable compared to vanilla adversarial training, where PGD with sufficient iterations is used to solve the inner maximization problem.
+Our work falls into the second category, where we adopt early stopping to accelerate the convergence of SSAT algorithms.
+Nevertheless, developing effective strategies based on our data reduction schemes and more efficient inner maximization solvers represents a promising future direction for further improving the practicality of SSAT algorithms.